### v1

Abstract
In this work, we propose a noise-centric formulation of the distributionally robust Kalman filter (DRKF) for discrete-time linear stochastic systems with uncertain noise statistics. By placing Wasserstein ambiguity sets directly on the process and measurement noise distributions, the proposed DRKF preserves the analytical structure of the classical Kalman filter while providing a priori spectral bounds on all feasible covariances. In the time-invariant setting, we derive a steady-state DRKF from a single stationary semidefinite program, yielding a constant-gain estimator with the same per-step computational complexity as the standard Kalman filter. We establish conditions guaranteeing the existence, uniqueness, and convergence of this steady-state solution, and we prove its asymptotic minimax optimality with respect to the worst-case mean-square error. Numerical experiments validate the theory and demonstrate that the proposed DRKF improves estimation accuracy under unknown or uncertain noise models while offering computational advantages over existing robust and distributionally robust filters.
1 Introduction
State estimation is a central problem in control, perception, signal processing, and robotics, where the goal is to infer system states from noisy and partial measurements. The Kalman filter (KF) [1] is optimal for linear–Gaussian systems when the noise statistics are known. In practice, however, these statistics are rarely available a priori; they are estimated from limited data, subject to modeling error, and often vary over time. Even mild misspecification can degrade estimation accuracy, while severe mismatch may lead to filter divergence.
To address this sensitivity, several robust estimation frameworks have been proposed. The $H_{\infty}$ filter [2] bounds the disturbance-to-error energy gain and guarantees finite error under worst-case noise, but it is often conservative in stochastic settings due to its pessimistic nature. Risk-sensitive filters [3, 4] penalize large errors exponentially, but their performance depends on a risk parameter that lacks a clear statistical interpretation and is difficult to tune, frequently resulting in excessive conservatism or instability.
Recent advances in distributionally robust optimization (DRO) have inspired the development of distributionally robust state estimators (DRSE). Although distributional robustness has already evolved into a rich research area in control design (e.g., [5, 6, 7, 8, 9, 10, 11]), its exploration in state estimation remains relatively limited. DRSE methods generally assume that the true noise distributions lie within an ambiguity set centered at a nominal distribution and seek estimators that minimize the worst-case expected error over this set. Within this framework, a variety of ambiguity sets have been considered, including $\phi$-divergence balls, moment-based ambiguity sets [12, 13, 14, 15], and Wasserstein balls [16, 17, 18, 19, 20]. Importantly, DRSE formulations differ not only in the type of ambiguity set employed, but also in the distributions to which the ambiguity is applied. For example, [16] imposes a Wasserstein ambiguity set on the joint state–measurement distribution; [17] places separate ambiguity sets on the prior state and measurement-noise covariances; and [19] introduces a bicausal optimal-transport ambiguity set that explicitly encodes temporal causality.
Despite this progress, steady-state formulations of DRSE remain underdeveloped. Analyzing the infinite-horizon regime is crucial for long-term operation and computational tractability: time-varying DRKFs typically require solving an optimization problem at every time step or a large optimization problem whose size grows with the horizon. The frequency-domain minimax formulation of [20] characterizes the optimal linear time-invariant estimator, but the resulting filter is non-rational and does not admit a finite-dimensional state-space realization.
Our previous work [21] developed a steady-state DRKF that placed ambiguity sets on the prior state and measurement-noise distributions. In this paper, we instead adopt a noise-centric formulation, placing Wasserstein ambiguity sets directly on the initial state, process noise, and measurement noise distributions. Unlike existing approaches (e.g., [20]), the resulting steady-state DRKF is obtained from a single stationary semidefinite program (SDP) that directly returns a constant-gain solution. This noise-centric formulation ensures a form of dynamic consistency—all admissible priors are generated by propagating the system dynamics under feasible noise sequences—and enables several theoretical and computational advantages.
Our main contributions are summarized as follows.
  • Noise-centric ambiguity modeling. Wasserstein ambiguity sets are defined directly on the initial state, process noise, and measurement noise distributions. This design $(i)$ enforces a form of dynamic consistency by ensuring that all admissible priors are reachable through the system dynamics under admissible noise realizations; $(ii)$ provides a priori spectral bounds on the DRKF covariance recursion, unlike ambiguity sets using joint state–measurement laws, which may include unreachable priors and typically lack such spectral guarantees; and $(iii)$ yields a tractable time-varying DRKF whose stage-wise gains are precomputed via small convex SDPs and applied online with Kalman-level per-step complexity.
  • Steady-state DRKF via a single SDP. A constant-gain DRKF is obtained by solving one stationary SDP offline. Its online computational cost matches that of the classical KF. The resulting steady-state covariances solve a time-invariant convex program, and the gains of the time-varying DRKF converge to this stationary gain.
  • Theoretical guarantees. $(i)$ Spectral boundedness and KF sandwiching: Wasserstein balls impose explicit eigenvalue bounds on feasible covariances, ensuring that the DRKF’s prior and posterior covariances lie between those of two appropriately chosen KFs. This property prevents arbitrarily inflated or singular covariances under worst-case deviations in the noise statistics. $(ii)$ Explicit convergence radii: we derive closed-form sufficient conditions on the ambiguity radii guaranteeing contraction of the DR Riccati recursion and convergence of the time-varying DRKF to its steady-state counterpart. $(iii)$ Stability and optimality: under standard controllability and observability assumptions, the estimation error dynamics are Schur stable, and the steady-state DRKF achieves asymptotic minimax mean-square error (MSE) optimality within the specified ambiguity sets.
The remainder of this paper is organized as follows. Section˜2 reviews preliminaries on DRSE and Wasserstein ambiguity sets. Section˜3 presents the time-varying and steady-state DRKF algorithms and establishes their key properties. Section˜4 develops convergence, stability, and optimality results. Section˜5 demonstrates the performance and utility of the proposed framework through numerical experiments.
2 Problem Setup
We begin by introducing notation. Let $\mathbb{S}^{n}$ denote the set of symmetric matrices in $\real{n\times n}$, with $\mathbb{S}^{n}_{+}$ and $\mathbb{S}^{n}_{++}$ denoting the positive semidefinite and positive definite cones. For $A,B\in\mathbb{S}^{n}$, the relations $A\succeq B$ and $A\succ B$ indicate that $A-B\in\mathbb{S}^{n}_{+}$ and $A-B\in\mathbb{S}^{n}_{++}$, respectively. For $A\in\real{m\times n}$, $\|A\|_{2}$ and $\|A\|_{F}$ denote the spectral and Frobenius norms. For $A\in\mathbb{S}^{n}$, $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denote its smallest and largest eigenvalues, and $A^{+}$ denotes its Moore–Penrose pseudoinverse.
Let $\mathcal{P}(\mathcal{W})$ be the set of Borel probability measures supported on a set $\mathcal{W}\subseteq\real{n}$, and let $\mathcal{P}_{2}(\mathcal{W})$ denote the subset of measures with finite second-order moments. Given two probability measures $\mathbb{P}$ and $\mathbb{Q}$, their product measure is denoted by $\mathbb{P}\otimes\mathbb{Q}$. For a sequence $\{w_{k}\}_{k\geq 0}$ with $w_{k}\in\real{n}$, the stacked block vector of length $N$ at time $t$ is defined as $\bm{w}_{t}^{N}\coloneqq[\,w_{tN+N-1}^{\top},\dots,w_{tN}^{\top}\,]^{\top}$.
2.1 Distributionally Robust State Estimation Problem
Consider the discrete-time linear stochastic system [^1] Throughout, we assume that the system matrices $A$ and $C$ are known, and we focus exclusively on uncertainty in the noise statistics.
$\begin{split}x_{t+1}&=A_{t}x_{t}+w_{t},\\ y_{t}&=C_{t}x_{t}+v_{t},\end{split}$ (1)
where $x_{t}\in\real{n_{x}}$ and $y_{t}\in\real{n_{y}}$ denote the system state and the output. The process noise $w_{t}\in\real{n_{x}}$ and the measurement noise $v_{t}\in\real{n_{y}}$ follow distributions $\mathbb{Q}_{w,t}\in\mathcal{P}(\real{n_{x}})$ and $\mathbb{Q}_{v,t}\in\mathcal{P}(\real{n_{y}})$, while the initial state $x_{0}$ follows $\mathbb{Q}_{x,0}\in\mathcal{P}(\real{n_{x}})$. We assume that $x_{0}$, $\{w_{t}\}$, and $\{v_{t}\}$ are mutually independent, and that each of $\{w_{t}\}$ and $\{v_{t}\}$ is temporally independent; that is, $w_{t}$ and $w_{t^{\prime}}$ (and likewise $v_{t}$ and $v_{t^{\prime}}$) are independent for all $t\neq t^{\prime}$. [^2] Such independence assumptions are standard in the control and estimation literature; see, e.g., [22, Sec. 2.2] and [2, Sec. 5.1].
At each time $t\geq 0$, let $\mathcal{Y}_{t}\coloneqq\{y_{0},\dots,y_{t}\}$ denote the measurement history. The goal is to estimate $x_{t}$ given $\mathcal{Y}_{t}$. Since outputs arrive sequentially, a natural approach is to compute the minimum mean-square error (MMSE) estimator recursively:
$\min_{\psi_{t}\in\mathcal{F}_{t}}J_{t}(\psi_{t},\mathbb{Q}_{e,t})\coloneqq\mathbb{E}\left[\|x_{t}-\psi_{t}(\mathcal{Y}_{t})\|^{2}\mid\mathcal{Y}_{t-1}\right],$ (2)
where $\mathcal{F}_{t}=\{\,\psi_{t}:(\real{n_{y}})^{t+1}\rightarrow\real{n_{x}}:\psi_{t}\text{ measurable}\,\}$ denotes the set of admissible estimators. The expectation is taken with respect to the conditional joint distribution of $(x_{t},y_{t})$ given the past observations $\mathcal{Y}_{t-1}$. Under (1), this conditional distribution is fully determined, for $t>0$, by the posterior distribution $\mathbb{P}_{x,t-1}$ of $x_{t-1}$ given $\mathcal{Y}_{t-1}$ and the joint noise distribution $\mathbb{Q}_{e,t}\coloneqq\mathbb{Q}_{w,t-1}\otimes\mathbb{Q}_{v,t}$ of $(w_{t-1},v_{t})$. At $t=0$, no posterior from a previous step exists. By slight abuse of notation, we adopt the convention $\mathcal{Y}_{-1}=\emptyset$ and set $\mathbb{P}_{x,0}^{-}\coloneqq\mathbb{Q}_{x,0}$ and $\mathbb{Q}_{e,0}\coloneqq\mathbb{P}_{x,0}^{-}\otimes\mathbb{Q}_{v,0}$, so that the definition in (2) also covers $t=0$, where the expectation reduces to an unconditional expectation with respect to the joint law of $(x_{0},v_{0})$. [^3] We use the superscript “-” for prior quantities; e.g., $\bar{x}^{-}_{t}$, $\Sigma_{x,t}^{-}$, and $\mathbb{P}_{x,t}^{-}$ denote the prior mean, covariance, and distribution of $x_{t}$ conditioned on $\mathcal{Y}_{t-1}$.
When $x_{0}\sim\mathcal{N}(\hat{x}_{0}^{-},\Sigma_{x,t}^{-})$ and $\{w_{t}\},\{v_{t}\}$ are independent Gaussian sequences with known covariances, the optimal causal MMSE estimator reduces to the classical KF. In practice, noise statistics are estimated from limited data and thus subject to ambiguity. Typically only nominal distributions $\hat{\mathbb{Q}}_{w,t}$, $\hat{\mathbb{Q}}_{v,t}$, and $\hat{\mathbb{Q}}_{x,0}$ are available from identification methods [23, 24, 25, 26]. [^4] A hat, $\hat{\cdot}$, denotes nominal quantities; e.g., $\hat{x}_{0}^{-}$ and $\hat{\mathbb{Q}}_{x,0}$ represent the nominal initial state mean and distribution, respectively. Even modest deviations between nominal and true distributions can degrade performance or cause divergence, especially when disturbances accumulate over time (e.g., marginally stable or unstable systems).
To address this, we formulate a stage-wise minimax problem in which the estimator competes against an adversary that selects distributions from ambiguity sets $\mathbb{D}_{j,t}\subset\mathcal{P}(\real{n_{j}})$ for $j\in\{w,v,x_{0}\}$. Each ambiguity set is centered at its corresponding nominal distribution and parameterized by a robustness radius. The DRSE problem at time $t$ is then given by
$\min_{\psi_{t}\in\mathcal{F}_{t}}\max_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t}),$ (3)
where $\mathbb{D}_{e,0}\coloneqq\mathbb{D}_{x,0}\times\mathbb{D}_{v,0}$ and $\mathbb{D}_{e,t}\coloneqq\mathbb{D}_{w,t-1}\times\mathbb{D}_{v,t}$ for $t>0$. The corresponding joint distributions are given by $\mathbb{P}_{e,0}\coloneqq\mathbb{P}_{x,0}^{-}\otimes\mathbb{P}_{v,0}$ and $\mathbb{P}_{e,t}\coloneqq\mathbb{P}_{w,t-1}\otimes\mathbb{P}_{v,t}$. This product form reflects the assumed independence between process and measurement noise at each stage.
Remark 1.
In contrast to [17] and [21], which define the ambiguity set around the joint prior-measurement distribution, our formulation places ambiguity directly on the noise distributions. This enforces a form of dynamic consistency: priors arise only through the system dynamics (1) and admissible noise realizations. By contrast, DR-MMSE formulations may allow priors that are not dynamically reachable. Our formulation also enables the adversary to couple estimation performance across time, capturing long-term accumulation of disturbances via the state recursion, even though the ambiguity sets themselves are stage-wise. For systems with slow or unstable dynamics, even small biases in $w_{t}$ can lead to unbounded state growth, making this sequential robustness essential.
While DRSE formulations have been explored in several contexts [17, 18, 16], their application to control-oriented state estimation remains limited. Moreover, understanding their steady-state behavior as $t\to\infty$ is critical for assessing long-term robustness and computational tractability.
2.2 Wasserstein Ambiguity Sets
Wasserstein ambiguity sets are widely used in the DRO literature due to their strong analytical and computational properties, including convex reformulations (e.g., [27, 28]). Unlike $\phi$-divergence balls, which may include statistically irrelevant distributions, Wasserstein balls account for distances between support elements and thus capture meaningful perturbations. This property prevents pathological distributions and leads to more reliable solutions [29].
The type-2 Wasserstein distance between two probability distributions $\mathbb{P},\mathbb{Q}\in\mathcal{P}(\mathcal{W})$ is defined as
$W_{2}(\mathbb{P},\mathbb{Q})\coloneqq\inf_{\tau\in\mathcal{T}(\mathbb{P},\mathbb{Q})}\Bigg\{\!\left(\int_{\mathcal{W}\times\mathcal{W}}\|x-y\|^{2}\,d\tau(x,y)\right)^{\!\frac{1}{2}}\Bigg\},$
where $\mathcal{T}(\mathbb{P},\mathbb{Q})$ denotes the set of joint probability measures on $\mathcal{W}\times\mathcal{W}$ with marginals $\mathbb{P}$ and $\mathbb{Q}$. The optimization variable $\tau$ represents the transportation plan that redistributes probability mass from $\mathbb{P}$ to $\mathbb{Q}$. [^5] We use the Euclidean norm to measure the transportation cost $\|x-y\|$.
We define the ambiguity sets as Wasserstein balls centered at nominal distributions:
$\mathbb{D}_{j,t}\coloneqq\left\{\mathbb{P}_{j,t}\in\mathcal{P}(\real{n_{j}})\,|\,W_{2}(\mathbb{P}_{j,t},\hat{\mathbb{Q}}_{j,t})\leq\theta_{j}\right\}$
for $j\in\{w,v,x_{0}\}$, where $\theta_{j}$ denotes the Wasserstein ball radius, which specifies the level of robustness. [^6] Although we model the process and measurement noise uncertainties using separate ambiguity sets, the formulation can be extended to correlated noise by placing a single Wasserstein ball around their joint distribution. Extending the theory to such joint ambiguity sets is conceptually straightforward, but it leads to higher-dimensional optimization problems for filter design.
In general, computing $W_{2}$ explicitly is difficult. A useful surrogate is the Gelbrich distance. Specifically, the Gelbrich distance between $\mathbb{P}\in\mathcal{P}(\mathbb{R}^{n})$ and $\mathbb{Q}\in\mathcal{P}(\mathbb{R}^{n})$ with mean vectors $\mu_{1},\mu_{2}\in\real{n}$ and covariance matrices $\Sigma_{1},\Sigma_{2}\in\mathbb{S}^{n}_{+}$ is defined as
$G(\mathbb{P},\mathbb{Q})\coloneqq\sqrt{\|\mu_{1}-\mu_{2}\|_{2}^{2}+\mathcal{B}^{2}(\Sigma_{1},\Sigma_{2})},$
where
$\mathcal{B}(\Sigma_{1},\Sigma_{2})\coloneqq\sqrt{\operatorname{Tr}[\Sigma_{1}+\Sigma_{2}-2\big(\Sigma_{2}^{\frac{1}{2}}\Sigma_{1}\Sigma_{2}^{\frac{1}{2}}\big)^{\frac{1}{2}}]}$
is the Bures–Wasserstein distance.
The Gelbrich distance always provides a lower bound on the type-2 Wasserstein distance, i.e., $G(\mathbb{P},\mathbb{Q})\leq W_{2}(\mathbb{P},\mathbb{Q})$. Moreover, equality holds when both distributions are Gaussian, or more generally, elliptical with the same density generator [30].
3 Distributionally Robust Kalman Filters
We now develop tractable DR state estimators consistent with the noise-centric ambiguity model introduced in Section˜2. We first present a finite-horizon formulation whose recursion mirrors that of the KF under the least-favorable noise covariances. We then specialize to the time-invariant case and derive a steady-state DRKF that uses a single stationary SDP, resulting in a constant-gain filter with Kalman-level online complexity.
3.1 Finite-Horizon Case
We begin with assumptions on the nominal distributions that guarantee tractability of the DRSE problem and preserve the analytical structure of the classical KF, specifically the Gaussian and affine relationships that enable closed-form recursive updates.
Assumption 1.
The nominal distributions of the initial state $x_{0}$, process noise $w_{t}$, and measurement noise $v_{t}$ are Gaussian for all $t$, i.e., $\hat{\mathbb{Q}}_{x,0}=\mathcal{N}(\hat{x}_{0}^{-},\hat{\Sigma}_{x,0}^{-}),\hat{\mathbb{Q}}_{w,t}=\mathcal{N}(\hat{w}_{t},\hat{\Sigma}_{w,t})$, and $\hat{\mathbb{Q}}_{v,t}=\mathcal{N}(\hat{v}_{t},\hat{\Sigma}_{v,t})$ with mean vectors $\hat{x}_{0}^{-}\in\real{n_{x}},\hat{w}_{t}\in\real{n_{x}}$, and $\hat{v}_{t}\in\real{n_{y}}$ and covariance matrices $\hat{\Sigma}_{x,0}^{-}\in\mathbb{S}^{n_{x}}_{+},\hat{\Sigma}_{w,t}\in\mathbb{S}^{n_{x}}_{+},\hat{\Sigma}_{v,t}\in\mathbb{S}^{n_{y}}_{++}$, respectively.
Under this assumption, the DRSE problem (3) admits the following convex reformulations.
Lemma 1.
Suppose Assumption˜1 holds. Then, the DRSE problem (3) satisfies the following properties.
$(i)$ At the initial stage, the following convex optimization problem has the same optimal value as (3), and its optimal solutions are optimal for (3):
$\begin{split}\max_{\Sigma_{x,0}^{-},\Sigma_{v,0}}&\operatorname{Tr}[\Sigma_{x,0}^{-}-\Sigma_{x,0}^{-}C_{0}^{\top}(C_{0}\Sigma_{x,0}^{-}C_{0}^{\top}+\Sigma_{v,0})^{-1}C_{0}\Sigma_{x,0}^{-}]\\ \mbox{s.t.}\;&\mathrm{Tr}[\Sigma_{x,0}^{-}+\hat{\Sigma}_{x,0}^{-}-2((\hat{\Sigma}_{x,0}^{-})^{\frac{1}{2}}\Sigma_{x,0}^{-}(\hat{\Sigma}_{x,0}^{-})^{\frac{1}{2}})^{\frac{1}{2}}]\leq\theta_{x_{0}}^{2}\\ &\mathrm{Tr}[\Sigma_{v,0}+\hat{\Sigma}_{v,0}-2\big(\hat{\Sigma}_{v,0}^{\frac{1}{2}}\Sigma_{v,0}\hat{\Sigma}_{v,0}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{v}^{2}\\ &\Sigma_{x,0}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\Sigma_{v,0}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}\\ &\Sigma_{x,0}^{-}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,0}\in\mathbb{S}^{n_{y}}_{+}.\end{split}$ (4)
$(ii)$ For any $t>0$, fix the state distribution $\mathbb{P}_{x,t-1}=\mathcal{N}(\bar{x}_{t-1},\Sigma_{x,t-1})$. Then the following convex optimization problem has the same optimal value as (3), and its optimal solutions are optimal for (3):
$\begin{split}\max_{\Sigma_{w,t-1},\Sigma_{v,t}}\;&\operatorname{Tr}[\Sigma_{x,t}^{-}-\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t})^{-1}C_{t}\Sigma_{x,t}^{-}]\\ \mbox{s.t.}\;&\operatorname{Tr}[\Sigma_{w,t-1}+\hat{\Sigma}_{w,t-1}-2\big(\hat{\Sigma}_{w,t-1}^{\frac{1}{2}}\Sigma_{w,t-1}\hat{\Sigma}_{w,t-1}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,t}+\hat{\Sigma}_{v,t}-2\big(\hat{\Sigma}_{v,t}^{\frac{1}{2}}\Sigma_{v,t}\hat{\Sigma}_{v,t}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{v}^{2}\\ &\Sigma_{x,t}^{-}=A_{t}\Sigma_{x,t-1}A_{t}^{\top}+\Sigma_{w,t-1}\\ &\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}\\ &\Sigma_{w,t-1}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,t}\in\mathbb{S}^{n_{y}}_{+}.\end{split}$ (5)
$(iii)$ If $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ and $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ solve (4) and (5), respectively, then the maximum in (3) is attained by the Gaussian distributions $\mathbb{P}_{x,0}^{-,*}=\mathcal{N}(\hat{x}_{0}^{-},\Sigma_{x,0}^{-,*}),\mathbb{P}_{w,t}^{*}=\mathcal{N}(\hat{w}_{t},\Sigma_{w,t}^{*})$ and $\mathbb{P}_{v,t}^{*}=\mathcal{N}(\hat{v}_{t},\Sigma_{v,t}^{*})$ for each $t\geq 0$. Furthermore, at any time $t$, the optimal state estimator achieving the minimum in (3) is affine and given by
$\psi_{t}^{*}(\mathcal{Y}_{t})=\bar{x}^{-}_{t}+\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t}^{*})^{-1}(y_{t}-C_{t}\bar{x}^{-}_{t}-\hat{v}_{t}),$ (6)
with $\bar{x}^{-}_{t}=A_{t-1}\bar{x}_{t-1}+\hat{w}_{t-1},\Sigma_{x,t}^{-}=A_{t}\Sigma_{x,t-1}A_{t}^{\top}+\Sigma_{w,t}^{*}$, and $\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*},\bar{x}^{-}_{0}=\hat{x}_{0}^{-}$.
The problems (4) and (5) define the stage-wise minimax interaction: the estimator selects $\psi_{t}$, while the adversary selects noise covariances within Wasserstein balls. Specifically, problem (4) yields the least-favorable prior and measurement covariances $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ at $t=0$, whereas (5) yields the least-favorable process and measurement covariances $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ for $t\geq 1$. The corresponding optimal values represent the least-favorable posterior MSE at each stage.
Lemma˜1 extends [17, Theorem 3.1] to the noise-centric setting by replacing ambiguity on the prior state (for $t>0$) with ambiguity on the process noise. Importantly, although the Wasserstein constraints are expressed in terms of first- and second-order moments, this does not restrict the adversary in a meaningful way under Assumption˜1. For Gaussian nominal distributions, the least-favorable distributions are Gaussian, and their worst-case behavior is fully characterized by their means and covariances. Furthermore, even without Gaussian nominal assumptions, (6) remains minimax optimal within the class of affine estimators [17].
Remark 2.
By adapting the structural argument used in [17, Lemma A.3], DRSE problem (3) admits at least one maximizer whose covariance matrices satisfy $\Sigma_{x,0}^{-,*}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\Sigma_{v,0}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}$, for $t=0$, and $\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}$ for all $t>0$ (see Appendix A). Because such a maximizer always exists, the explicit eigenvalue lower-bound constraints included in (4) and (5) are redundant: they do not alter the optimal value but simply restrict attention to maximizers that satisfy these bounds. In particular, the versions of (4) and (5) without these explicit bounds are equivalent to the original DRSE problem (3).
Remark 3.
In the inner maximization problems (4) and (5), the least-favorable distributions retain the nominal noise means. Any shift of a noise mean would introduce a systematic bias in the estimation error while also increasing the Wasserstein distance, making such a shift never advantageous for the adversary. Because of this trade-off, the ambiguity budget is used entirely to perturb the covariances, which is reflected in Lemma˜1 and Theorem˜1.
Both problems (4) and (5) are solvable for $\hat{\Sigma}_{v,t}\in\mathbb{S}^{n_{y}}_{++}$ since their objectives are continuous over compact feasible sets. Applying a standard Schur complement argument (e.g., [31, Appendix 5.5]), these problems reduce to SDPs that are efficiently solvable using off-the-shelf solvers.
Corollary 1.
Suppose Assumption˜1 holds. Then, for any time $t>0$, the optimization problem (5) is equivalent to the following tractable convex SDP:
$\begin{split}\max_{\begin{subarray}{c}\Sigma_{x,t}^{-},\Sigma_{x,t},\\ \Sigma_{w,t-1},\Sigma_{v,t},\\ Y,Z\end{subarray}}\;&\operatorname{Tr}[\Sigma_{x,t}]\\ \mbox{s.t.}\;&\begin{bmatrix}\Sigma_{x,t}^{-}-\Sigma_{x,t}&\Sigma_{x,t}^{-}C_{t}^{\top}\\ C_{t}\Sigma_{x,t}^{-}&C_{t}\Sigma_{x,t}^{-}C^{\top}+\Sigma_{v,t}\end{bmatrix}\succeq 0\\ &\begin{bmatrix}\hat{\Sigma}_{w,t-1}&Y\\ Y^{\top}&\Sigma_{w,t-1}\end{bmatrix}\succeq 0,\;\begin{bmatrix}\hat{\Sigma}_{v,t}&Z\\ Z^{\top}&\Sigma_{v,t}\end{bmatrix}\succeq 0\\ &\operatorname{Tr}[\Sigma_{w,t-1}+\hat{\Sigma}_{w,t-1}-2Y]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,t}+\hat{\Sigma}_{v,t}-2Z]\leq\theta_{v}^{2}\\ &\Sigma_{x,t}^{-}=A_{t}\Sigma_{x,t-1}A_{t}^{\top}+\Sigma_{w,t-1}\\ &\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}\\ &\Sigma_{x,t}^{-},\Sigma_{x,t},\Sigma_{w,t-1}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,t}\in\mathbb{S}^{n_{y}}_{+}\\ &Y\in\real{n_{x}\times n_{x}},\;Z\in\real{n_{y}\times n_{y}}.\end{split}$ (7)
Similar results hold for $t=0$. Solving (7) offline for all $t$ yields the least-favorable covariances $(\Sigma_{w,t}^{*},\Sigma_{v,t}^{*})$. The next theorem shows that these covariances can be used directly in a Kalman-style recursion, resulting in a DR state estimator that is robust to distributional uncertainty.
Theorem 1 (DR Kalman Filter).
Under Assumption˜1, the minimax DR state estimate $\psi^{*}_{t}(\mathcal{Y}_{t})$ at any time $t$ coincides with the conditional mean $\bar{x}_{t}$ under the least-favorable distributions. Moreover, the least-favorable prior and posterior remain Gaussian: $\mathbb{P}_{x,t}^{-,*}=\mathcal{N}(\bar{x}^{-}_{t},\Sigma_{x,t}^{-})$, $\mathbb{P}_{x,t}^{*}=\mathcal{N}(\bar{x}_{t},\Sigma_{x,t})$. These distributions are recursively computed for $t=0,1,\dots$ as follows:
  • (Measurement Update) Compute the DR Kalman gain
    $K_{t}=\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}{\Sigma_{x,t}^{-}}C_{t}^{\top}+\Sigma_{v,t}^{*})^{-1},$ (8)
    where $\Sigma_{v,t}^{*}$ is the maximizer of (5) for $t>0$ and (4) for $t=0$. Then, update $\bar{x}_{t}$ and $\Sigma_{x,t}$ based on the measurement $y_{t}$ as follows:
    $\bar{x}_{t}=K_{t}(y_{t}-C_{t}\bar{x}^{-}_{t}-\hat{v}_{t})+\bar{x}^{-}_{t}$ (9)
    $\Sigma_{x,t}=(I-K_{t}C_{t})\Sigma_{x,t}^{-}.$ (10)
    The initial values are $\bar{x}^{-}_{0}=\hat{x}_{0}^{-},\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*}$, where $\Sigma_{x,0}^{-,*}$ is the maximizer of (4) .
  • (State Prediction) Predict $\bar{x}^{-}_{t+1}$ and $\Sigma_{x,t+1}^{-}$ as
    $\begin{split}\bar{x}^{-}_{t+1}&=A_{t}\bar{x}_{t}+\hat{w}_{t}\end{split}$ (11)
    $\Sigma_{x,t+1}^{-}$ $=A_{t}\Sigma_{x,t}A_{t}^{\top}+\Sigma_{w,t}^{*},$ (12)
    where $\Sigma_{w,t}^{*}$ is the maximizer of (5) for $t>0$.
Notably, the measurement update and state prediction equations (9)–(12) mirror those of the classical KF when the system is driven by the least-favorable distributions $\mathbb{P}_{x,0}^{-,*},\mathbb{P}_{w,t}^{*}$, and $\mathbb{P}_{v,t}^{*}$.
3.2 Infinite-Horizon Case
While the DRKF in Theorem˜1 provides a tractable solution to the finite-horizon DRSE problem, our focus is on its long-term behavior. To study this regime, we assume that both the system dynamics and the nominal uncertainty statistics are time invariant.
Assumption 2.
The system (1) is time invariant, i.e., $A_{t}\equiv A,C_{t}\equiv C$. Moreover, the nominal uncertainty distributions are stationary, so that $\hat{w}_{t}\equiv\hat{w},\hat{v}_{t}\equiv\hat{v}$ and $\hat{\Sigma}_{v,t}\equiv\hat{\Sigma}_{v},\hat{\Sigma}_{w,t}\equiv\hat{\Sigma}_{w}$ for all $t\geq 0$, where $\hat{\Sigma}_{w}\in\mathbb{S}^{n_{x}}_{++}$ and $\hat{\Sigma}_{v}\in\mathbb{S}^{n_{y}}_{++}$.
Under Assumption˜2, the stage-wise SDPs (5)–(7) reduce to a time-invariant recursion for the covariance matrices. The resulting sequences of least-favorable noise covariances $\{\Sigma_{w,t}^{*},\Sigma_{v,t}^{*}\}$ and state covariances $\{\Sigma_{x,t}^{-},\Sigma_{x,t}\}$ may converge to limiting values $\{\Sigma_{w,\infty},\Sigma_{v,\infty}\}$ and $\{\Sigma_{x,\infty}^{-},\Sigma_{x,\infty}\}$ as $t\to\infty$. Whether such convergence holds depends on additional conditions, which will be established later in Section˜4.2. Motivated by this possibility, we seek a steady-state DR estimator with constant gain.
In the infinite-horizon setting, the DRSE problem reduces to finding these steady-state covariances and the corresponding steady-state DR Kalman gain. Since the dynamics and nominal statistics are fixed, the stage-wise SDP (5) simplifies to the following time-invariant convex program:
$\begin{split}\max_{\begin{subarray}{c}\Sigma_{w,\infty},\Sigma_{v,\infty}\\ \Sigma_{x,\infty}^{-},\Sigma_{x,\infty}\end{subarray}}\;&\operatorname{Tr}[\Sigma_{x,\infty}]\\ \mbox{s.t.}\;&\Sigma_{x,\infty}=\Sigma_{x,\infty}^{-}-\Sigma_{x,\infty}^{-}C^{\top}(C\Sigma_{x,\infty}^{-}C^{\top}+\Sigma_{v,\infty})^{-1}C\Sigma_{x,\infty}^{-}\\ &\Sigma_{x,\infty}^{-}=A\Sigma_{x,\infty}A^{\top}+\Sigma_{w,\infty}\\ &\operatorname{Tr}[\Sigma_{w,\infty}+\hat{\Sigma}_{w}-2\big(\hat{\Sigma}_{w}^{\frac{1}{2}}\Sigma_{w,\infty}\hat{\Sigma}_{w}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,\infty}+\hat{\Sigma}_{v}-2\big(\hat{\Sigma}_{v}^{\frac{1}{2}}\Sigma_{v,\infty}\hat{\Sigma}_{v}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{v}^{2}\\ &\Sigma_{w,\infty}\succeq\lambda_{\min}(\hat{\Sigma}_{w})I_{n_{x}},\Sigma_{v,\infty}\succeq\lambda_{\min}(\hat{\Sigma}_{v})I_{n_{y}}\\ &\Sigma_{x,\infty}^{-},\Sigma_{x,\infty},\Sigma_{w,\infty}\in\mathbb{S}^{n_{x}}_{+},\Sigma_{v,\infty}\in\mathbb{S}^{n_{y}}_{+}.\end{split}$
As in (7), the optimization problem above can be reformulated as an equivalent steady-state SDP:
$\begin{split}\max_{\begin{subarray}{c}\Sigma_{x,\infty}^{-},\Sigma_{x,\infty},\\ \Sigma_{w,\infty},\Sigma_{v,\infty},\\ Y,Z\end{subarray}}\;&\operatorname{Tr}[\Sigma_{x,\infty}]\\ \mbox{s.t.}\;&\begin{bmatrix}\Sigma_{x,\infty}^{-}-\Sigma_{x,\infty}&\Sigma_{x,\infty}^{-}C^{\top}\\ C\Sigma_{x,\infty}^{-}&C\Sigma_{x,\infty}^{-}C^{\top}+\Sigma_{v,\infty}\end{bmatrix}\succeq 0\\ &\begin{bmatrix}\hat{\Sigma}_{w}&Y\\ Y^{\top}&\Sigma_{w,\infty}\end{bmatrix}\succeq 0,\quad\begin{bmatrix}\hat{\Sigma}_{v}&Z\\ Z^{\top}&\Sigma_{v,\infty}\end{bmatrix}\succeq 0\\ &\Sigma_{x,\infty}^{-}=A\Sigma_{x,\infty}A^{\top}+{\Sigma}_{w,\infty}\\ &\operatorname{Tr}[\Sigma_{w,\infty}+\hat{\Sigma}_{w}-2Y]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,\infty}+\hat{\Sigma}_{v}-2Z]\leq\theta_{v}^{2}\\ &\Sigma_{w,\infty}\succeq\lambda_{\min}(\hat{\Sigma}_{w})I_{n_{x}},\Sigma_{v,\infty}\succeq\lambda_{\min}(\hat{\Sigma}_{v})I_{n_{y}}\\ &\Sigma_{x,\infty}^{-},\Sigma_{x,\infty},\Sigma_{w,\infty}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,\infty}\in\mathbb{S}^{n_{y}}_{+}\\ &Y\in\real{n_{x}\times n_{x}},\;Z\in\real{n_{y}\times n_{y}}.\end{split}$ (13)
The optimal solution $\{\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*},\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*}\}$ defines the least-favorable stationary covariances, from which the corresponding steady-state DR Kalman gain is
$K_{\infty}^{*}:=\Sigma_{x,\infty}^{-,*}C^{\top}(C\Sigma_{x,\infty}^{-,*}C^{\top}+\Sigma_{v,\infty}^{*})^{-1}.$ (14)
The resulting time-invariant state estimator is
[…]
with prior
$\bar{x}^{-}_{t+1}=A\bar{x}_{t}+\hat{w}$. (16)
Compared with the finite-horizon DRKF, the steady-state filter (15) requires solving only a single stationary SDP offline. This makes it well-suited for real-time implementation. More importantly, it optimizes directly for worst-case asymptotic performance, ensuring robustness to persistent disturbances and long-term distributional errors. These properties are particularly valuable in safety-critical systems with slow or unstable dynamics.
3.3 Algorithms
We now summarize the proposed DRKF methods for both the time-varying and steady-state settings. These algorithms follow directly from the tractable reformulations in Section˜3 and the recursive structure established in Theorem˜1. Although both versions share the same online recursion as the classical KF, they differ in how the covariance matrices, and hence the Kalman gains, are computed to ensure robustness against distributional ambiguity.
3.3.1 Time-Varying DRKF
As shown in Algorithm˜1, the time-varying DRKF is implemented by first solving the stage-wise SDP problems (4) and (5) offline for each time step. If the nominal covariance matrices are known a priori, the least-favorable noise covariances and the DR gains $\{K_{t}\}$ can be precomputed offline for a given horizon $T$. At each stage, the gain $K_{t}$ is computed using the least-favorable prior state covariance and the least-favorable measurement noise covariance. The online stage then consists only of the measurement update and state prediction steps (9) and (11). Upon receiving a new measurement $y_{t}$, the update and prediction steps are carried out exactly as in the classical KF, except that they use the DR gains computed offline and the nominal noise means.
Algorithm 1 Time-Varying DRKF
1:System matrices $\{A_{t},C_{t}\}_{t=0}^{T-1}$, nominal means and covariances $(\hat{w}_{t},\hat{v}_{t},\hat{\Sigma}_{w,t},\hat{\Sigma}_{v,t})$, prior state mean $\hat{x}_{0}^{-}$ and covariance $\hat{\Sigma}_{x,0}^{-}$, ambiguity radii $(\theta_{w},\theta_{v},\theta_{x_{0}})$, horizon length $T$
2:Solve (4) to obtain $(\Sigma_{x,0}^{-*},\Sigma_{v,0}^{*})$ $\triangleright$ Offline stage
3:Compute DR gain $K_{0}$ via (8)
4:for $t=1,\dots,T-1$ do
5:  Solve (7) to obtain $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*},\Sigma_{x,t}^{-},\Sigma_{x,t})$
6:  Compute DR gain $K_{t}$ via (8)
7:end for
8:Set $\bar{x}^{-}_{0}=\hat{x}_{0}^{-}$ $\triangleright$ Online stage
9:for $t=0,\dots,T-1$ do
10:  Observe $y_{t}$
11:  Update the state estimate $\bar{x}_{t}$ via (9) using $K_{t}$
12:  Predict the prior mean $\bar{x}^{-}_{t+1}$ via (11)
13:end for
3.3.2 Steady-State DRKF
Under the time-invariant assumptions of Section˜3.2, the DRKF recursion may converge to a stationary solution. In this setting, the procedure in Algorithm˜1 simplifies considerably. Rather than solving an SDP at every stage, one solves the single stationary SDP (13) offline to obtain the least-favorable noise covariances, the steady-state prior and posterior covariances, and the resulting steady-state DR Kalman gain $K_{\infty}^{*}$. The online recursion is identical in structure to the classical steady-state KF, making the approach particularly suitable for long-horizon or real-time applications. The complete procedure is given in Algorithm˜2.
Algorithm 2 Steady-State DRKF
1:System matrices $(A,C)$, nominal means and covariances $(\hat{w},\hat{v},\hat{\Sigma}_{w},\hat{\Sigma}_{v})$, prior state mean $\hat{x}_{0}^{-}$, ambiguity radii $(\theta_{w},\theta_{v})$
[…]
3:Solve (13) to obtain $(\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*},\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*})$
4:Compute the steady-state DR gain via (14)
5:Set $\bar{x}^{-}_{0}=\hat{x}_{0}^{-}$ $\triangleright$ Online stage
6:for $t=0,1,2,\dots$ do
7:  Observe $y_{t}$
8:  Update the state estimate $\bar{x}_{t}$ via (15) using $K_{\infty}^{*}$
9:  Predict the prior mean $\bar{x}^{-}_{t+1}$ via (16)
10:end for
In the steady state regime, ambiguity in the initial state distribution becomes irrelevant: the Wasserstein radius $\theta_{x_{0}}$ affects only the transient phase and does not influence the asymptotic solution. As discussed in Section˜4.2 and Remark˜5, this dependence disappears after finitely many steps. Thus, the steady-state DRKF depends only on $(\theta_{w},\theta_{v})$ and the nominal noise covariances.
3.4 Discussion on Computational Complexity
The main computational burden in both DRKF variants arises in the offline stage, where SDPs must be solved. The online recursion has the same per-step complexity as the classical KF.
[…]
For comparison, the stage-wise DRKF of [16] and the DR-MMSE estimator of [17] also require solving a convex program at each stage. The former uses the joint state-measurement covariance of dimension $(n_{x}+n_{y})$, leading to similar polynomial scaling but with larger LMI blocks and constants. Like our time-varying DRKF, these methods scale linearly in $T$; however, neither admits a steady-state variant that reduces the offline computation to a single SDP.
Temporally coupled ambiguity sets, as in [20], can model long-range dependence across the noise sequence but incur substantially higher computational cost: all $T$ stages are merged into a single horizon-level SDP whose number of variables and LMI block sizes both grow linearly with $T$. The resulting interior-point complexity, $\tilde{\mathcal{O}}(n_{x}^{6}T^{6})$, quickly becomes prohibitive even for moderate horizons. By contrast, our noise-centric formulation solves a fixed-size SDP at each stage, so the offline cost grows only linearly in $T$.
In the infinite-horizon setting, [20] characterizes the steady-state filter via a frequency-domain max-min problem over Toeplitz operators, producing a generally non-rational filter that requires spectral factorization and approximation for implementation. No finite-dimensional formulation is available. By contrast, our steady-state DRKF is obtained from a single stationary SDP of size $(n_{x}+n_{y})$ with offline complexity $\tilde{\mathcal{O}}\left((n_{x}+n_{y})^{3.5}\right)$, directly yielding a constant gain $K_{\infty}^{*}$ and an online recursion matching the classical steady-state KF. Although this stage-wise ambiguity does not model temporal noise correlations, it preserves the separability of the Riccati recursion and keeps the DR filter tractable for real-time use.
4 Theoretical Analyses
Having established the DRKF formulations in both the finite- and infinite-horizon settings, we now analyze their fundamental theoretical properties. Our goal is to show that DRKF retains key structural guarantees of the classical KF, such as stability and boundedness, while providing provable robustness under distributional uncertainty. In particular, we demonstrate that the DRKF covariance recursion remains spectrally bounded and that its steady-state solution inherits standard convergence properties of the classical filter.
4.1 Spectral Boundedness
We begin by quantifying how Wasserstein ambiguity sets constrain the least-favorable covariance matrices, and how these constraints propagate through the DRKF recursion in Theorem˜1. This analysis shows that the DRKF operates within a uniformly bounded spectral envelope, a property that underpins its stability and distinguishes it from several existing robust formulations.
Our starting point is a well-known characterization of the Bures–Wasserstein distance.
Lemma 2.
[33, Theorem 1] Let $A,B\in\mathbb{S}^{n}_{+}$. The Bures–Wasserstein distance satisfies
$\mathcal{B}(A,B)\;=\;\min_{U\in O(n)}\|A^{1/2}-B^{1/2}U\|_{F},$
where $O(n)$ denotes the orthogonal group. The minimizer exists and is given by the unitary factor in the polar decomposition of $B^{1/2}A^{1/2}$.
This characterization enables sharp spectral bounds for matrices within a Bures–Wasserstein ball.
Proposition 1.
Fix $\hat{X}\in\mathbb{S}^{n}_{+}$. For any $X\in\mathbb{S}^{n}_{+}$ satisfying $\mathcal{B}(X,\hat{X})\leq\theta$, the minimum and maximum eigenvalues of $X$ are bounded as
$\underline{\lambda}(\hat{X},\theta)\;\leq\;\lambda_{\min}(X)\;\leq\;\lambda_{\max}(X)\;\leq\;\overline{\lambda}(\hat{X},\theta),$
where
$\underline{\lambda}(\hat{X},\theta)$ $\coloneqq\left(\max\left\{0,\sqrt{\lambda_{\min}(\hat{X})}-\theta\right\}\right)^{2},$
$\overline{\lambda}(\hat{X},\theta)$ $\coloneqq\left(\sqrt{\lambda_{\max}(\hat{X})}+\theta\right)^{2}.$
The proof is provided in Appendix B.1. This result shows that Bures–Wasserstein balls impose both lower and upper eigenvalue bounds on any feasible covariance matrix. Equivalently, the Bures–Wasserstein ball defines an eigenvalue tube around the nominal covariance. This geometric interpretation is essential: it guarantees that our DRKF cannot produce arbitrarily inflated or degenerate covariances, even under worst-case perturbations of the noise distributions allowed by the ambiguity sets. In contrast, KL-based ambiguity sets can admit pathological least-favorable distributions [29].
We now translate these spectral bounds to the least-favorable noise covariances that enter the DRKF recursion.
Corollary 2.
Let $\hat{\Sigma}_{v,t}\in\mathbb{S}^{n_{y}}_{++}$ and $\hat{\Sigma}_{w,t}\in\mathbb{S}^{n_{x}}_{+}$ denote the nominal noise covariances. For ambiguity radii $\theta_{v},\theta_{w},\theta_{x_{0}}$, define
$\begin{split}\underline{\lambda}_{v,t}&:=\lambda_{\min}(\hat{\Sigma}_{v,t}),\qquad\overline{\lambda}_{v,t}:=\overline{\lambda}(\hat{\Sigma}_{v,t},\theta_{v}),\\ \underline{\lambda}_{w,t}&:=\lambda_{\min}(\hat{\Sigma}_{w,t}),\qquad\overline{\lambda}_{w,t}:=\overline{\lambda}(\hat{\Sigma}_{w,t},\theta_{w})\\ \underline{\lambda}_{x,0}&:=\lambda_{\min}(\hat{\Sigma}_{x,0}^{-}),\qquad\overline{\lambda}_{x,0}:=\overline{\lambda}(\hat{\Sigma}_{x,0}^{-},\theta_{x_{0}}).\end{split}$ (17)
Then, for all $t>0$, the least-favorable covariances $\Sigma_{v,t}^{*},\Sigma_{w,t}^{*}$ and $\Sigma_{x,0}^{-,*}$ obtained from (4)–(5) satisfy
$\underline{\lambda}_{v,t}I_{n_{y}}\preceq\Sigma_{v,t}^{*}\preceq\overline{\lambda}_{v,t}I_{n_{y}},$
$\underline{\lambda}_{w,t}I_{n_{x}}\preceq\Sigma_{w,t}^{*}\preceq\overline{\lambda}_{w,t}I_{n_{x}},$
$\underline{\lambda}_{x,0}I_{n_{x}}\preceq\Sigma_{x,0}^{-,*}\preceq\overline{\lambda}_{x,0}I_{n_{x}}.$
The proof is provided in Appendix B.2. In the time-invariant setting, the constants $\underline{\lambda}_{v},\overline{\lambda}_{v},\underline{\lambda}_{w},\overline{\lambda}_{w}$ are fixed. Thus, Corollary˜2 ensures that the DRKF operates within a compact spectral envelope determined solely by the nominal statistics and the Wasserstein radii.
Theorem 2.
Consider two classical KFs with noise covariances chosen according to the lower and upper bounds in (17). Let $(\underline{\Sigma}_{x,t}^{-},\underline{\Sigma}_{x,t})$ and $(\overline{\Sigma}_{x,t}^{-},\overline{\Sigma}_{x,t})$ denote their respective prior and posterior error covariances, with initial conditions $\underline{\Sigma}_{x,0}^{-}=\underline{\lambda}_{x,0}I$ and $\overline{\Sigma}_{x,0}^{-}=\overline{\lambda}_{x,0}I$. Then, for all $t\geq 0$, the DRKF covariances $(\Sigma_{x,t}^{-},\Sigma_{x,t})$ satisfy
$\underline{\Sigma}_{x,t}^{-}\preceq\Sigma_{x,t}^{-}\preceq\overline{\Sigma}_{x,t}^{-},\qquad\underline{\Sigma}_{x,t}\preceq\Sigma_{x,t}\preceq\overline{\Sigma}_{x,t}.$
The proof is provided in Appendix B.3. Given the initial prior state covariance, the lower and upper bounds in Theorem˜2 can be precomputed by running two standard KF recursions before solving the optimization problems (4) (for $t=0$) and (5) (for $t>0$).
This spectral boundedness is a direct consequence of our ambiguity set design, which models uncertainty separately for the process and measurement noise distributions. In contrast, previous DRKFs based on ambiguity sets centered on the joint prior-measurement distributions, as in [21], do not directly admit such precomputed spectral bounds. This highlights an additional advantage of our noise-centric construction, complementing the benefit noted in Remark˜1.
4.2 Convergence
Having established the key theoretical properties of the least-favorable covariance matrices, we now derive conditions under which the DRKF admits a unique steady-state solution, and the sequence of time-varying gains converges to its steady-state value. Specifically, recall that the steady-state DR Kalman gain $K_{\infty}$ is defined in (14). Our goal is to show that
$\lim_{t\to\infty}K_{t}=K_{\infty}.$
Under the standing assumptions, the steady-state SDP (13) has at least one optimal solution, since its objective is continuous and the feasible set is nonempty and compact due to the Bures-Wasserstein spectral bounds. Similarly, the time-varying SDP problems (4) and (5) admit optimal solutions, and therefore the sequence of DR Kalman gains $\{K_{t}\}$ is well defined.
The prior covariance recursion (12) can be equivalently expressed as
$\Sigma_{x,t+1}^{-}=A\bigl((\Sigma_{x,t}^{-})^{-1}+C^{\top}(\Sigma_{v,t}^{*})^{-1}C\bigr)^{-1}A^{\top}+\Sigma_{w,t}^{*},$ (18)
with $\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*}$. Since $\Sigma_{x,0}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}}$ and $\Sigma_{x,t}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{w})I_{n_{x}}$ for $t>0$ at the optimum, $\Sigma_{x,t}^{-}$ is positive definite and the recursion is well defined. This motivates the following definition of the DR Riccati mapping:
$r_{\mathbb{D}}(\Sigma_{x,t}^{-}):=A\bigl((\Sigma_{x,t}^{-})^{-1}+C^{\top}(\Sigma_{v,t}^{*})^{-1}C\bigr)^{-1}A^{\top}+\Sigma_{w,t}^{*}.$ (19)
As in the classical KF, controllability and observability are needed to guarantee a unique steady state.
Assumption 3.
The pair $(A,\hat{\Sigma}_{w}^{1/2})$ is controllable and the pair $(A,C)$ is observable.
To isolate the effect of distributional robustness in the process channel, we parameterize the perturbation through $\Delta_{t}$.
Assumption 4.
There exists $\Delta_{t}\in\mathbb{S}^{n_{x}}_{+}$ such that $A\Delta_{t}A^{\top}=\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}$.
Under Assumption˜4, the recursion (18) can be rewritten as
$r_{\mathbb{D}}(\Sigma_{x,t}^{-})=A\bigl((\Sigma_{x,t}^{-})^{-1}+C^{\top}\hat{\Sigma}_{v}^{-1}C-\Phi_{t}\bigr)^{-1}A^{\top}+\hat{\Sigma}_{w},$ (20)
where
$\Phi_{t}\coloneqq(\Sigma_{x,t}^{-})^{-1}+C^{\top}\hat{\Sigma}_{v}^{-1}C-\left(I+X_{t}\Delta_{t}\right)^{-1}X_{t},$ (21)
and
$X_{t}:=(\Sigma_{x,t}^{-})^{-1}+C^{\top}(\Sigma_{v,t}^{*})^{-1}C.$
Assumption˜4 is a structural compatibility condition ensuring that the perturbation in the process-noise covariance lies in the range of the dynamics matrix $A$. While the DR Riccati iteration often converges numerically even when this condition is mildly violated, our formal convergence and uniqueness guarantees rely on it.
Remark 4.
Assumption˜4 holds whenever $\mathrm{range}(\Sigma_{w,t}^{*}-\hat{\Sigma}_{w})\subseteq\mathrm{range}(A)$. In this case, a canonical choice of $\Delta_{t}$ is the minimum-Frobenius-norm solution
$\Delta_{t}:=A^{+}(\Sigma_{w,t}^{*}-\hat{\Sigma}_{w})(A^{+})^{\top}.$ (22)
This matrix $\Delta_{t}$ is positive semidefinite, satisfies $A\Delta_{t}A^{\top}=\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}$, and obeys the bound
$\|\Delta_{t}\|_{2}\leq\|A^{+}\|_{2}^{2}\|\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}\|_{2}.$
In practice, verifying $\mathrm{range}(\Sigma_{w,t}^{*}-\hat{\Sigma}_{w})\subseteq\mathrm{range}(A)$ requires knowledge of the time-varying covariance $\Sigma_{w,t}^{*}$. A convenient sufficient condition that can be checked a priori uses the spectral bound $\overline{\lambda}_{w}$ from Corollary˜2: if $\mathrm{range}(\overline{\lambda}_{w}I-\hat{\Sigma}_{w})\subseteq\mathrm{range}(A)$, then Assumption˜4 holds for all $t$, since
$0\preceq\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}\preceq\overline{\lambda}_{w}I-\hat{\Sigma}_{w}\implies\mathrm{range}(\Sigma_{w,t}^{*}-\hat{\Sigma}_{w})\subseteq\mathrm{range}(\overline{\lambda}_{w}I-\hat{\Sigma}_{w})\subseteq\mathrm{range}(A).$
The form in (20) is structurally analogous to the risk-sensitive Riccati equation [4, 34]. Levy et al. [35] showed that, under Assumption˜3, the $N$-fold composition of the risk-sensitive Riccati mapping is a contraction in a suitable Riemannian metric for $N\geq n_{x}$, while Zorzi [36] extended this result to robust filters under the Thompson part metric. These results motivate the use of contraction-based arguments for analyzing the DR Riccati mapping.
An alternative interpretation of (20) arises by embedding the DR filter in a Krein space [37]. Augmenting the nominal system with a synthetic uncertainty $u_{t}$ and enforcing the constraint $0=\hat{x}_{t}+u_{t}$ yields an indefinite inner product on $u_{t}$ weighted by $\Phi_{t}^{-1}$. This perspective links distributional robustness to risk-sensitive filtering and provides a natural foundation for contraction-based analyses.
To formalize this idea, fix $N\in\mathbb{N}$ and consider the downsampled state $\hat{x}_{t}^{d}=\hat{x}_{tN}$. Downsampling lifts the time-varying recursion to an $N$-block time-invariant map, yielding block-diagonal perturbations and enabling a contraction argument similar to [37]. This induces an equivalent block model with matrices $(\mathcal{R}_{N},\mathcal{O}_{N},\mathcal{L}_{N},\mathcal{H}_{N})$ and a block-diagonal perturbation
$\bar{\Phi}_{N,t}=\mathrm{blkdiag}(\Phi_{tN+N-2},\ldots,\Phi_{tN-1}).$
The details are provided in Appendix C.1. Letting $\Sigma_{x,t}^{-,d}\coloneq\Sigma_{x,tN}^{-}$ denote the downsampled prior covariance matrix, the resulting Riccati recursion becomes
$\Sigma_{x,t+1}^{-,d}=r_{\mathbb{D},t}^{d}(\Sigma_{x,t}^{-,d}),$ (23)
where $r_{\mathbb{D},t}^{d}$ denotes the $N$-fold composition of $r_{\mathbb{D}}$. If $r_{\mathbb{D},t}^{d}$ is a contraction for sufficiently large $t$, then so is $r_{\mathbb{D}}$. By Banach’s fixed-point theorem [38], this ensures the existence and uniqueness of a fixed point $\Sigma_{x,\infty}^{-}\succ 0$, and hence of a steady-state gain $K_{\infty}$.
The following proposition provides a tractable sufficient condition for ensuring contraction.
Proposition 2.
Suppose Assumptions˜1, 2, 3 and 4 hold. Define
$\tilde{\phi}_{N}\;=\;\frac{1}{\lambda_{\max}\left(\mathcal{L}_{N}\bigl(I_{Nn_{x}}+\mathcal{H}_{N}^{\top}(\mathcal{D}_{N}\mathcal{D}_{N}^{\top})^{-1}\mathcal{H}_{N}\bigr)^{-1}\mathcal{L}_{N}^{\top}\right)}.$
Then, for some $N\geq n_{x}$, there exists $\phi_{N}\in(0,\tilde{\phi}_{N})$ such that if
$0\preceq\bar{\Phi}_{N,t}\preceq\phi_{N}I_{Nn_{x}},\qquad\forall t\geq q,$ (24)
for some $q>0$, the mapping $r_{\mathbb{D},t}^{d}$ is a contraction with respect to the Thompson part metric [39] on the cone of positive definite matrices for all $t\geq q$, and hence $r_{\mathbb{D}}$ is also a contraction in the same metric.
The proof is provided in Appendix C.2. A practical procedure for determining a suitable $\phi_{N}$ is described in [36]. One begins with $\phi_{N}=\tilde{\phi}_{N}$ and checks whether the associated Gramian $\Omega_{\phi_{N}I_{Nn_{x}}}$ is positive definite. If not, $\phi_{N}$ is iteratively reduced until $\Omega_{\phi_{N}I_{Nn_{x}}}$ becomes positive semidefinite but singular. At this threshold, if $\bar{\Phi}_{N,t}\prec\phi_{N}I_{Nn_{x}}$ for all $t\geq\tilde{q}$, then both Gramians $\Omega_{\bar{\Phi}_{N,t}}$ and $W_{\bar{\Phi}_{N,t}}$ are positive definite for $t\geq\tilde{q}$, which guarantees that the downsampled DR Riccati mapping (23) is a contraction.
Proposition˜2 establishes convergence provided that $\bar{\Phi}_{N,t}$ remains uniformly bounded by $\phi_{N}I_{Nn_{x}}$. In practice, however, $\bar{\Phi}_{N,t}$ depends on the sequence of SDP solutions in (7) and is therefore difficult to verify a priori. This motivates the derivation of explicit, tractable sufficient conditions on the ambiguity radii $(\theta_{v},\theta_{w})$ that guarantee the bound $\bar{\Phi}_{N,t}\preceq\phi_{N}I_{Nn_{x}}$ uniformly across time. Here, our noise-centric modeling becomes essential, as it leads to closed-form, computationally tractable expressions for these radii.
Theorem 3.
Suppose Assumptions˜3, 1, 2 and 4 hold. Let $\phi_{v},\phi_{x}\geq 0$ with $\phi_{v}+\phi_{x}\leq\phi_{N}$, where $\phi_{N}$ is defined in Proposition˜2. Then, the following properties hold.
  1. (ii)
    (Measurement uncertainty bound) For any $\phi_{v}\geq 0$, define
    $\delta_{v}^{\max}(\phi_{v})\coloneqq\frac{\phi_{v}(\lambda_{\min}(\hat{\Sigma}_{v}))^{2}}{\|C\|_{2}^{2}+\phi_{v}\lambda_{\min}(\hat{\Sigma}_{v})}<\lambda_{\min}(\hat{\Sigma}_{v}).$ (25)
    Then, any $\theta_{v}$ satisfying
    $0\leq\theta_{v}\leq\sqrt{\lambda_{\max}(\hat{\Sigma}_{v})+\delta_{v}^{\max}(\phi_{v})}-\sqrt{\lambda_{\max}(\hat{\Sigma}_{v})}$ (26)
    guarantees that, for all $t\geq q$,
    $C^{\top}\!\bigl(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\bigr)C\;\preceq\;\phi_{v}I.$ (27)
  2. (iiii)
    (Process uncertainty bound) Define
    $\delta_{x}^{\max}\;\coloneqq\;\frac{1}{\lambda_{\min}(\hat{\Sigma}_{w})}+\frac{\|C\|_{2}^{2}}{\lambda_{\min}(\hat{\Sigma}_{v})}.$ (28)
    If $0\leq\phi_{x}<\delta_{x}^{\max}$, then any $\theta_{w}$ satisfying
    $0\leq\theta_{w}\leq\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})+\frac{\phi_{x}}{\|A^{+}\|_{2}^{2}\,\delta_{x}^{\max}(\delta_{x}^{\max}-\phi_{x})}}-\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}$ (29)
    guarantees that, for all $t\geq q$,
    $X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}\preceq\phi_{x}I.$ (30)
Moreover, any pair $(\theta_{v},\theta_{w})$ satisfying (26) and (29) ensures that the DR Riccati mapping (19) admits a unique fixed point $\Sigma_{x,\infty}^{-}\succ 0$, and the corresponding DR Kalman gain converges: $K_{t}\to K_{\infty}$.
The proof of this theorem is provided in Appendix C.4. The bounds (26) and (29) yield tractable, closed-form radii $(\theta_{v},\theta_{w})$ that certify convergence, directly in terms of nominal covariances and operator norms. This explicit tuning rule represents a practical consequence of our noise-centric ambiguity design.
Remark 5.
The radius $\theta_{x_{0}}$ is arbitrary and affects only the transient phase. The sufficient conditions (26) and (29) depend only on $(\theta_{v},\theta_{w})$ and the nominal noise covariances. In particular, for any $\Sigma_{x,0}^{-}\in\mathbb{S}^{n_{x}}_{++}$, the first prediction satisfies $\Sigma_{x,1}^{-}\succeq\Sigma_{w,0}^{\star}\succeq\hat{\Sigma}_{w}$. Thereafter, the contraction bounds are independent of the prior. Hence, if  (26) and (29) hold uniformly, one may take $q=0$ in Theorem˜3; otherwise, convergence is still guaranteed after a finite transient $q\geq 1$.
An immediate implication of Theorem˜3 is that the steady-state solution can be computed directly through the single stationary SDP (13) offline, rather than via online Riccati iterations.
Corollary 3.
Suppose that Assumptions˜1, 2, 3 and 4 hold. Then, the steady-state covariance matrices $(\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*})$, together with the least-favorable noise covariances $\Sigma_{w,\infty}^{*}$ and $\Sigma_{v,\infty}^{*}$, solve the stationary SDP (13). Consequently, $K_{\infty}^{*}=K_{\infty}=\lim_{t\to\infty}K_{t}$.
Under Assumption˜3, the Riccati recursions of the standard KFs in Theorem˜2 converge. When the DRKF covariances also converge, we can take limits in the matrix inequality of Theorem˜2 and, using that limits preserve positive semidefiniteness, obtain steady-state upper and lower bounds on the DRKF covariance, stated in the following corollary.
Corollary 4.
Suppose that Assumptions˜1, 2, 3 and 4 hold and the DRKF converges. Then, the steady-state prior and posterior error covariances $\Sigma_{x,\infty}^{-}$ and $\Sigma_{x,\infty}$ are bounded by the limits of the lower- and upper-bound Kalman recursions described in Theorem˜2:
$\underline{\Sigma}_{x,\infty}^{-}$ $\preceq\Sigma_{x,\infty}^{-}\preceq\overline{\Sigma}_{x,\infty}^{-}$
$\underline{\Sigma}_{x,\infty}$ $\preceq\Sigma_{x,\infty}\preceq\overline{\Sigma}_{x,\infty},$
where $\underline{\Sigma}_{x,\infty}^{-}\coloneqq\lim_{t\to\infty}\underline{\Sigma}_{x,t}^{-}$ and $\overline{\Sigma}_{x,\infty}^{-}\coloneqq\lim_{t\to\infty}\overline{\Sigma}_{x,t}^{-}$ denote the steady-state bounds for the prior covariances, and the posterior bounds $\underline{\Sigma}_{x,\infty}$ and $\overline{\Sigma}_{x,\infty}$ are defined analogously.
Finally, the DR Riccati recursion (18) reveals a structural connection to existing robust filtering paradigms. In particular, it resembles the risk-sensitive Riccati equation [4, 34]. This correspondence implies that the time-varying DRKF can be interpreted as a risk-sensitive filter with an effective risk parameter $\theta$ and a time-varying weighting matrix $Q_{t}\in\mathbb{S}^{n_{x}}_{+}$ satisfying $\theta Q_{t}^{\top}Q_{t}=\Phi_{t-1}$. At steady state, DRKF reduces to a risk-sensitive filter with a constant weighting matrix $Q$ chosen such that $\theta Q^{\top}Q=\Phi_{\infty}$, where $\Phi_{\infty}$ denotes the limiting perturbation.
4.3 Stability
A fundamental property of the classical KF is the asymptotic stability of its estimation error dynamics under standard controllability and observability assumptions. Specifically, the closed-loop error matrix (IKC)A(I-KC)A is Schur, ensuring boundedness of the estimation error covariance and decay of the error mean in the absence of bias.
In the DR setting, however, the estimator competes against the least-favorable noise distributions within the ambiguity set. A natural question is therefore whether the DRKF preserves this stability property. The following result shows that, despite the adversarial nature of the noise, the DRKF error dynamics remain asymptotically stable in the same sense as in the classical case.
Theorem 4.
Suppose Assumptions˜3, 1, 2 and 4 hold, and the ambiguity radii $(\theta_{v},\theta_{w})$ satisfy the conditions of Theorem˜3. Let $(\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*})$ and $(\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*})$ denote the steady-state DRKF covariances, and let $K_{\infty}$ be the corresponding steady-state DR gain. Define the closed-loop error matrix $F_{\infty}\coloneq(I-K_{\infty}C)A$. Then, $F_{\infty}$ is Schur stable, and the posterior estimation error $e_{t}:=x_{t}-\bar{x}_{t}$ evolves according to
[…]
Consequently, under the least-favorable noise distributions,
$\lim_{t\to\infty}\mathbb{E}\|e_{t}\|^{2}\;=\;\operatorname{Tr}[\Sigma_{x,\infty}^{*}]\quad\text{and}\quad\sup_{t\geq 0}\,\mathbb{E}\|e_{t}\|^{2}\;<\;\infty.$
The proof is provided in Appendix D.1.
While Theorem˜4 guarantees mean-square stability, it does not by itself characterize the estimation-error mean in the presence of bias. When the least-favorable distributions (or the true noise processes) have nonzero means, the filter tracks a steady bias determined by the mean mismatch. The following corollary characterizes this bias.
Corollary 5.
Under the assumptions of Theorem˜4, let $m_{t}:=\mathbb{E}[e_{t}]=\mathbb{E}[x_{t}-\bar{x}_{t}]$ denote the posterior error mean under the least-favorable distributions. Let $\mathbb{E}[w_{t}]=\mu_{w}$ and $\mathbb{E}[v_{t}]=\mu_{v}$. Then,
[…]
and hence
[…]
In particular, if the noise means used by the filter coincide with the true means ($\hat{w}=\mu_{w}$ and $\hat{v}=\mu_{v}$), then $\mathbb{E}[e_{t}]\to 0$.
The proof of this corollary is provided in Appendix D.2. Corollary˜5 shows that our DRKF mirrors the classical KF in its treatment of bias: if the true noise means coincide with those used by the filter ($\mu_{w}=\hat{w}$ and $\mu_{v}=\hat{v}$), then the estimation-error mean vanishes asymptotically. If instead either mean is misspecified ($\mu_{w}\neq\hat{w}$ or $\mu_{v}\neq\hat{v}$), the filter remains stable and the bias converges to the constant vector in the right-hand side above, which depends only on $F_{\infty}$, $K_{\infty}$, and the mean mismatch.
4.4 Optimality
A key property of the classical KF is that it is the unique linear, causal estimator minimizing the MSE when the Gaussian noise distributions are known exactly. In the DR setting, however, the true noise covariances are unknown and only assumed to lie within Wasserstein ambiguity sets. In this case, it is natural to consider minimax optimality: among all causal estimators, the filter should minimize the worst-case one-step MSE over all distributions in the ambiguity sets.
The following theorem establishes that DRKF satisfies this minimax optimality in an asymptotic sense.
Theorem 5.
Suppose Assumptions˜1, 2, 3 and 4 hold. Let $(\theta_{v},\theta_{w})$ satisfy the conditions of Theorem˜3, and let $\Sigma_{x,\infty}^{-,*}\in\mathbb{S}^{n_{x}}_{++}$ denote the unique fixed point of the DR Riccati map $r_{\mathbb{D}}$. Let $\Sigma_{x,\infty}^{*}$ be the corresponding steady-state posterior covariance. Then, for any causal state estimator sequence $\{\psi_{t}\}_{t\geq 0}$ with $\psi_{t}\in\mathcal{F}_{t}$, we have
$\liminf_{t\to\infty}\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})$ $\geq\operatorname{Tr}[\Sigma_{x,\infty}^{*}],$ (31)
$\liminf_{T\to\infty}\ \frac{1}{T}\sum_{t=1}^{T}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})$ $\geq\operatorname{Tr}[\Sigma_{x,\infty}^{*}],$ (32)
almost surely. Moreover, the steady-state DRKF $\psi_{\infty}$ with gain $K_{\infty}$ achieves these bounds with equality:
$\begin{split}\lim_{t\to\infty}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})&=\operatorname{Tr}[\Sigma_{x,\infty}^{*}],\\ \lim_{T\to\infty}\ \frac{1}{T}\sum_{t=1}^{T}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})&=\operatorname{Tr}[\Sigma_{x,\infty}^{*}],\end{split}$
almost surely. Hence, the steady-state DRKF is asymptotically minimax-optimal with respect to the worst-case one-step and long-run average MSE.
The proof of this theorem is provided in Appendix D.3. This result shows that the steady-state DRKF attains the smallest possible worst-case asymptotic MSE among all causal estimators, thereby extending the fundamental optimality of the classical KF to the DR setting. Importantly, the result concerns asymptotic performance rather than global finite-horizon minimax optimality, which is generally not well-posed under stage-wise ambiguity.
Remark 6.
When $\theta_{v}=\theta_{w}=0$, the ambiguity sets contain only the nominal distributions. The least-favorable covariances then coincide with the nominal ones, so the DRKF reduces exactly to the classical KF. Thus, Theorem˜5 recovers the standard optimality of the KF as a special case, demonstrating that the DRKF is a strict generalization of the classical optimal estimator to the DR regime.
Overall, the results in this section confirm that the proposed noise-centric DRKF preserves the essential analytical structure of the classical KF while extending it to systems subject to distributional ambiguity. The filter remains stable, convergent, and minimax optimal, thereby bridging stochastic state estimation and DRO. These theoretical guarantees demonstrate that DRKF is not only practical but also provably reliable under distributional and long-horizon uncertainty.
5 Experiments
We conducted numerical experiments to evaluate the proposed filters and validate their theoretical properties. In Section˜5.1, we compare our methods against a wide range of time-varying and steady-state filters, assessing estimation accuracy, closed-loop performance in a trajectory-tracking task, and computational efficiency. We also examine how the ambiguity radii and the accuracy of the nominal noise distributions affect estimation performance. In Section˜5.2.2, we empirically verify the theoretical results of Section˜4.888All experiments were performed on a laptop equipped with an Intel Core Ultra 7 155H @ 3.80 GHz and 32 GB of RAM. Source code is available at https://github.com/jangminhyuk/DRKF2025.
Figure 1: Effect of $\theta$ on the average regret MSE under a) Gaussian and b) U-Quadratic noise distributions, averaged over 20 simulation runs.
5.1 Comparison of Robust State Estimators
We evaluate both the time-varying and steady-state DRKFs against the following benchmark state estimators:
  1. (A)
    Standard time-varying KF
  2. (B)
    Steady-state KF
  3. (C)
    Risk-sensitive filter (risk-averse) [4, 34, 35, 36]
  4. (D)
    Risk-sensitive filter (risk-seeking) [4, 34, 35, 36]
  5. (E)
    DRKF with Wasserstein ambiguity on the joint law of $(x_{t},y_{t})$ [16]
  6. (F)
    DRKF with bicausal Wasserstein ambiguity [19]
  7. (G)
    Time-varying DRKF with Wasserstein ambiguity on the prior state and measurement noise [21, Theorem 1]
  8. (H)
    Time-invariant version of (G) with stationary nominals [21, Section 4]
  9. (I)
    Time-varying DRKF (ours) [Algorithm˜1]
  10. (J)
    Steady-state DRKF (ours) [Algorithm˜2]
We use the same symbol $\theta$ for the robustness parameter in all filters, although its role and scale differ across methods. For (C) and (D), $\theta$ is the risk-sensitivity coefficient, with (D) using $\theta$ as the risk-sensitivity parameter. For (C), we select $\theta$ up to the largest value for which the risk-sensitive Riccati mapping remains well defined [35, 36], i.e., $((\Sigma_{x,t}^{-})^{-1}-\theta D^{\top}D)$ remains positive definite with $D=I_{n_{x}}$. For (E) and (F), $\theta$ specifies the Wasserstein radius, and for (F), we follow [19] and cap the number of internal iterations at 20. For (G) and (H), we set $\theta_{x}=\theta_{v}=\theta$. In our proposed formulations, we set $\theta_{x_{0}}=\theta_{w}=\theta_{v}=\theta$ for (I), and $\theta_{w}=\theta_{v}=\theta$ for (J). To ensure the fairness, we sweep each method’s admissible range of $\theta$ and report performance at its best tuning.
5.1.1 Estimation Accuracy
Figure 2: Effect of the number of samples used to construct the nominal distributions on the average MSE under Gaussian noise, averaged over 10 runs.
We begin with a linear time-invariant system (no control input) subject to inaccuracies in both the process and measurement noise distributions:
[…]
We consider two noise models: $(i)$ Gaussian, with $x_{0}\sim\mathcal{N}(0,0.01I_{4})$, $w_{t}\sim\mathcal{N}(0,0.01I_{4})$, $v_{t}\sim\mathcal{N}(0,0.01I_{2})$, and $(ii)$ U-Quadratic, with $x_{0}\sim\mathcal{UQ}([-0.1,0.1]^{4})$, $w_{t}\sim\mathcal{UQ}([-0.1,0.1]^{4})$, $v_{t}\sim\mathcal{UQ}([-0.1,0.1]^{2})$. The nominal covariances $\hat{\Sigma}_{w}$ and $\hat{\Sigma}_{v}$ are learned from only 10 input-output samples using the expectation-maximization (EM) method in [19]. This induces a significant distributional mismatch, highlighting the challenge of learning accurate nominal models from limited data. The estimation horizon is set to 50 time steps.
Fig.˜1 illustrates the effect of $\theta\in[10^{-2},10^{1}]$ on the average regret MSE for each filter under Gaussian and U-Quadratic noise, averaged over 20 runs. The regret MSE is defined as the difference between the filter’s MSE and that of the KF using the true noise distributions, which serves as the MMSE-optimal benchmark. The results show that both our time-varying and steady-state DRKFs consistently achieve the lowest tuned regret and remain competitive across a broad range of $\theta$. In particular, (I) achieves performance closest to the optimal estimator under both poorly estimated Gaussian and non-Gaussian noise.
In comparison, (F) exhibits numerical instability and performance discontinuities across $\theta$, even when the internal iteration limit is increased to 50. The risk-sensitive filters (C) and (D) become overly conservative for large $\theta$, leading to degraded performance. These results highlight the limitations of existing methods, whereas our DRKFs maintain stable performance and a favorable robustness-accuracy tradeoff.
Fig.˜2 shows the average MSE as a function of the number of samples used to construct the nominal noise distributions. Each filter is optimally tuned over $\theta\in[10^{-2},10^{1}]$, and we vary the number of input-output samples used in the EM procedure so that small sample sizes correspond to less accurate nominal distributions. As the number of samples increases, the average MSE decreases for all filters. Across all sample sizes, our DRKF consistently achieves the lowest error, with a particularly clear advantage when only limited data are available.
Figure 3: Average regret MSE of the proposed steady-state DRKF under Gaussian noise, as a function of $\theta_{w}$ and $\theta_{v}$, averaged over 100 runs.
Fig.˜3 shows the average MSE regret of the steady-state DRKF (J) as a function of the ambiguity radii $\theta_{w}$ and $\theta_{v}$ over $10^{-2},10^{1}]$, averaged over 100 runs. Both parameters significantly influence performance: the regret is large when both radii are small, reflecting insufficient robustness against distributional mismatch. Increasing either $\theta_{w}$ or $\theta_{v}$ improves performance up to an optimal range, beyond which excessively large radii make the estimator overly conservative. This surface highlights the need to balance $\theta_{w}$ and $\theta_{v}$ in practice.
5.1.2 Trajectory Tracking Task
To assess the practical performance of our filters, we applied them to a 2D trajectory-tracking task, following the setup in [21, Section 5.2]:
$\begin{split}x_{t+1}&=\begin{bmatrix}I_{2}&\Delta tI_{2}\\ \mathbf{0}_{2\times 2}&I_{2}\end{bmatrix}x_{t}+\begin{bmatrix}0.5(\Delta t)^{2}I_{2}\\ \Delta tI_{2}\end{bmatrix}u_{t}+w_{t}\\ y_{t}&=\begin{bmatrix}I_{2}&\mathbf{0}_{2\times 2}\end{bmatrix}x_{t}+v_{t},\end{split}$
where $\Delta t=0.2\,\text{s}$. The state $x_{t}=[p_{t}^{x},p_{t}^{y},v_{t}^{x},v_{t}^{y}]^{\top}$ collects the planar position and velocity, and the control input $u_{t}=[a_{t}^{x},a_{t}^{y}]^{\top}$ represents acceleration commands.
We consider two noise settings: $(i)$ Gaussian, with $x_{0}\sim\mathcal{N}(0,0.02I_{4})$, $w_{t}\sim\mathcal{N}(0,0.02I_{4})$, $v_{t}\sim\mathcal{N}(0,0.02I_{2})$, and $(ii)$ nonzero-mean U-Quadratic, with $x_{0}\sim\mathcal{UQ}([-0.1,0.1]^{4})$, $w_{t}\sim\mathcal{UQ}([-0.1,0.2]^{4})$, $v_{t}\sim\mathcal{UQ}([-0.2,0.1]^{2})$, simulated over $10\,\text{s}$. For the nominal distributions, we fit an EM model using only $1\,\text{s}$ of input-output data, yielding noticeably inaccurate nominal statistics and reflecting the practical challenge of estimating noise models from limited observations.
Figure 4: Effect of $\theta$ on the average MPC cost under a) Gaussian and b) nonzero-mean U-Quadratic noise, averaged over 20 simulation runs.
Figure 5: 2D trajectories averaged over 200 simulations under Gaussian noise. Each filter uses its optimal parameter $\theta\in[10^{-2},10^{1}]$ minimizing the MPC cost. The black dashed line denotes the desired trajectory; colored curves show the mean, and shaded tubes indicate $\pm 1$ standard deviation.
Table 1: Mean and standard deviation of the acceleration-command magnitude $\|u_{t}\|_{2}$ for each filter, averaged over 200 simulations.
(A) (B) (E) (F) (G) (H) (I) (J)
$\|u_{t}\|_{2}$ 5.69 (7.44) 5.91 (7.70) 3.29 (1.93) 2.70 (1.78) 3.37 (2.02) 3.30 (1.99) 2.18 (1.63) 2.20 (1.64)
All methods use the same observer-based MPC controller (e.g., [40]), with the current state estimate $\bar{x}_{t}$ used for prediction. The controller solves a finite-horizon tracking problem with horizon $10$ and weights $Q=\mathrm{diag}([10,10,1,1])$ and $R=0.1I_{2}$, where $Q$ penalizes tracking error and $R$ penalizes control effort. The first control input $u_{t}$ is applied, and the optimization is repeated at the next step using $\bar{x}_{t+1}$. The MPC model, horizon, and weights are identical across all methods; the only component that changes is the state estimator, where the Kalman filter is replaced with each candidate estimator.
Fig.˜4 shows the effect of $\theta$ on the average MPC cost, defined as the closed-loop quadratic sum of tracking error and input energy weighted by $Q$ and $R$. Consistent with Section˜5.1.1, both our time-varying and steady-state DRKFs achieve the lowest cost when optimally tuned, even in the presence of non-Gaussian disturbances and nonzero-mean noise.
Fig.˜5 displays the mean and dispersion of the closed-loop trajectories obtained with the best-tuned radius $\theta\in[10^{-2},10^{1}]$ for each filter. The standard KFs (A) and (B) produce wider uncertainty tubes, indicating higher estimation variance. The BCOT-based DRKF (F) reduces dispersion but introduces greater bias, especially during sharp turns. The remaining DRKFs–(E), (G), and (H)–exhibit accurate tracking with modest dispersion. Our time-varying and steady-state DRKFs deliver the most reliable performance, achieving the lowest dispersion and bias across runs.
Although mean trajectories appear similar across DRKFs, the required control effort differs substantially. Table˜1 reports the time- and run-averaged acceleration magnitude $\|u_{t}\|_{2}$, which shows that our DRKFs achieve lower control energy among all methods.
5.1.3 Computation Time
The online computational cost of all filters is similar to that of the classical KF, since each estimator computes its gains offline. A key advantage of the steady-state DRKF is that it requires solving only a single SDP offline, whereas time-varying DRKFs must solve $T$ sequential SDPs.
Figure 6: Average offline computation time as a function of the horizon length $T$ for a system with $n_{x}=4$, $n_{y}=2$, averaged over 10 runs.
Table 2: Average offline computation time of the steady-state DRKF as a function of $n_{x}=n_{y}=n$ (10 runs).
$\bm{n}$ 5 10 15 20 25 30 35 40
Time (s) 0.0600 0.1794 0.4736 1.2414 2.2905 4.8900 8.3890 12.4232
Fig.˜6 compares the offline computation time of (E)(J) as a function of the horizon length $T$. We use the same setup as in Section˜5.1.1 with $\theta=1.0$ and Gaussian noise. Time-varying DRKFs exhibit linear growth in computation time with respect to $T$, with (F) being particularly expensive. In contrast, the steady-state DRKFs (G) and (J) maintain constant runtime regardless of $T$, demonstrating excellent scalability for long-horizon tasks.
Finally, Table˜2 evaluates scalability with respect to the system dimension. For each $n$, we generate a random system with $n_{x}=n_{y}=n$ and measure the offline computation time. Even at $n=40$, the total runtime remains under $13$ seconds, confirming the practical efficiency of the steady-state DRKF for moderately sized systems.
5.2 Validation of Theoretical Properties
5.2.1 DRKF Sandwich Properties
Theorem˜2 states that, at each time step, the DRKF prior and posterior covariances lie between those of two KFs run with extremal noise levels from (17). We illustrate this sandwich property on a concrete 2D example. Consider the system
$A=0.95\begin{bmatrix}\cos(\pi/8)&-\sin(\pi/8)\\ \sin(\pi/8)&\cos(\pi/8)\end{bmatrix},\quad C=\begin{bmatrix}1.0&0.1\\ 0.1&1.0\end{bmatrix}$,
with horizon $T=20$ and ambiguity radii $\theta_{w}=\theta_{v}=\theta_{x_{0}}=0.1$. The initial prior covariance is
$\Sigma_{x,0}^{-}=\begin{bmatrix}0.2&0.05\\ 0.05&0.1\end{bmatrix}$.
We use time-varying nominal covariances of the form $\hat{\Sigma}_{w,t}=a_{t}\hat{\Sigma}_{w}$, $\hat{\Sigma}_{v,t}=a_{t}\hat{\Sigma}_{v}$, constructed by scaling the base matrices
$\hat{\Sigma}_{w}=\begin{bmatrix}0.3&0.1\\ 0.1&0.2\end{bmatrix},\quad\hat{\Sigma}_{v}=\begin{bmatrix}0.15&0.05\\ 0.05&0.1\end{bmatrix}$,
with piecewise-linear factors $a_{t}$:
[…]
These covariances are used at each step $t$ to generate the lower-bound KF (LOW-KF), upper-bound KF (HIGH-KF), and DRKF posterior covariance trajectories. To visualize the uncertainty, we use the $95\%$ confidence ellipse $\mathcal{E}(\Sigma)\coloneqq\{x\in\mathbb{R}^{2}:x^{\top}\Sigma^{-1}x\leq\chi^{2}_{2,0.95}\}$ with $\chi^{2}_{2,0.95}=5.991$. This region contains $95\%$ of a zero-mean Gaussian with covariance $\Sigma$. By Theorem˜2, $\mathcal{E}(\underline{\Sigma}_{x,t})\subseteq\mathcal{E}(\Sigma_{x,t})\subseteq\mathcal{E}(\overline{\Sigma}_{x,t}),$ so the DRKF posterior ellipse must lie between the LOW-KF and HIGH-KF ellipses at every time step.
Figure 7: 95% confidence ellipses of the posterior covariances for LOW-KF, DRKF, and HIGH-KF at $t\in\{0,1,5,20\}$.
Figure 8: Three-dimensional tube visualization obtained by stacking the $95\%$ ellipses of the posterior covariances of LOW-KF, DRKF, and HIGH-KF for $t=0,\dots,20$.
As shown in Figs.˜7 and 8, the ellipses at representative time steps confirm the predicted nesting. The DRKF posterior covariance is consistently sandwiched between the LOW-KF and HIGH-KF bounds, providing a clear and interpretable robustness envelope throughout the horizon.
5.2.2 Convergence Properties
To evaluate the convergence behavior of the proposed filter, we consider the linear system from [37]:
$A=\begin{bmatrix}0.1&1\\ 1&-1\end{bmatrix},\;C=\begin{bmatrix}1&-1\end{bmatrix}$.
Figure 9: Relative difference between $\operatorname{Tr}[\Sigma_{x,t}]$ and $\operatorname{Tr}[\Sigma_{x,\infty}^{*}]$ over time.
We set the nominal covariances to $\hat{\Sigma}_{w}=I_{2}$ and $\hat{\Sigma}_{v}=1$. Following the procedure in Section˜4.2, we first compute $\phi_{N}$ with $N=8$ and $q=20$, and set $\phi_{v}=\phi_{x}=\phi_{N}/2$. We then choose the largest pair $(\theta_{v},\theta_{w})$ permitted by Theorem˜3.
To quantify convergence of the time-varying DRKF toward the steady-state solution, we track the relative error $|(\operatorname{Tr}[\Sigma_{x,\infty}^{*}]-\operatorname{Tr}[\Sigma_{x,t}])/\operatorname{Tr}[\Sigma_{x,\infty}^{*}]|$. As shown in Fig.˜9, this quantity decays approximately linearly on a logarithmic scale, demonstrating exponential convergence of the time-varying DRKF to the steady-state DRKF.
6 Conclusions and Future Work
This work presented a noise-centric formulation of DRKF in which uncertainty in the process and measurement noises is modeled directly via Wasserstein ambiguity sets. In contrast to prior DRKF formulations, the proposed approach preserves dynamic consistency and yields tractable spectral bounds on the least-favorable covariances. The resulting recursion retains the analytical structure of the classical KF while providing robustness to distributional errors. We established the existence, uniqueness, and convergence of the steady-state DRKF and proved that it is asymptotically minimax optimal with respect to the worst-case mean-square error. Notably, the steady-state solution is obtained from a single stationary SDP, yielding a constant-gain filter with Kalman-level online complexity and strong robustness guarantees. Empirical studies corroborate the theory, demonstrating improved estimation accuracy, tighter and more reliable uncertainty quantification, and superior closed-loop performance in control tasks.
Future directions include extending the framework to nonlinear systems and developing adaptive mechanisms for online calibration of the ambiguity set radius. It is also of interest to relax some of the structural assumptions underlying our convergence analysis (e.g., independence and range conditions on the process noise), to incorporate simultaneous model uncertainty in the system matrices, and to explore temporally coupled ambiguity sets that better capture long-range noise correlations while preserving computational tractability.
Appendix A Eigenvalue Lower Bounds for DRSE Maximizers
Lemma 3.
For $t=0$, the DRSE formulation (4) admits a maximizer pair $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ such that
$\Sigma_{x,0}^{-,*}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\Sigma_{v,0}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}.$ (33)
For $t>0$, the DRSE formulation (5) admits a maximizer pair $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ such that
$\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}.$ (34)
Proof.
For $t=0$, (33) is shown in [17, Theorem 3.1]. We now prove (34) by reformulating (5) in a form to which [17, Lemma A.3] applies. The objective of (5) is the trace of the posterior covariance matrix:
$\begin{split}\operatorname{Tr}[\Sigma_{x,t}]&=\operatorname{Tr}[\Sigma_{x,t}^{-}-\Sigma_{x,t}^{-}C^{\top}(C\Sigma_{x,t}^{-}C^{\top}+\Sigma_{v,t})^{-1}C\Sigma_{x,t}^{-}]\\ &=\inf_{K_{t}}\operatorname{Tr}[(I_{n_{x}}-K_{t}C_{t})\Sigma_{x,t}^{-}(I_{n_{x}}-K_{t}C_{t})^{\top}+K_{t}\Sigma_{v,t}K_{t}^{\top}]\\ &=\inf_{K_{t}}(\langle(I_{n_{x}}-K_{t}C_{t})^{\top}(I_{n_{x}}-K_{t}C_{t}),\Sigma_{x,t}^{-}\rangle+\langle K_{t}^{\top}K_{t},\Sigma_{v,t}\rangle)\end{split}$
where the last equality uses the inner product $\langle A,B\rangle:=\operatorname{Tr}[A^{\top}B]$. Given $\Sigma_{x,t-1}$, the deterministic part $A_{t}\Sigma_{x,t-1}A_{t}^{\top}$ is fixed, so $\Sigma_{x,t}^{-}=\bar{\Sigma}_{t}+\Sigma_{w,t-1}$, $\bar{\Sigma}_{t}\coloneqq A_{t}\Sigma_{x,t-1}A_{t}^{\top}\succeq 0$. Substituting this into the objective yields
$\begin{split}\operatorname{Tr}[\Sigma_{x,t}]&=\inf_{K_{t}}\bigg(\langle(I_{n_{x}}-K_{t}C_{t})^{\top}(I_{n_{x}}-K_{t}C_{t}),\Sigma_{w,t-1}\rangle\\ &\qquad+\langle(I_{n_{x}}-K_{t}C_{t})^{\top}(I_{n_{x}}-K_{t}C_{t}),\bar{\Sigma}_{t}\rangle+\langle K_{t}^{\top}K_{t},\Sigma_{v,t}\rangle\bigg)\end{split}$ (35)
First, we fix $\Sigma_{v,t}$ and view the objective as a function of $\Sigma_{w,t-1}$. From (35) this dependence can be written as
$\inf_{K_{t}}\Big(\langle(I_{n_{x}}-K_{t}C_{t})^{\top}(I_{n_{x}}-K_{t}C_{t}),\Sigma_{w,t-1}\rangle+f_{1,t}(K_{t})\Big),$ (36)
where $f_{1,t}(K_{t})$ collects the remaining terms, and is convex and continuous in $K_{t}$. Define the convex set $\mathcal{C}\coloneqq\{I_{n_{x}}-K_{t}C_{t}:K_{t}\in\mathbb{R}^{n_{x}\times n_{y}}\}$. For each $L\in\mathcal{C}$, define
$f(L)\coloneqq\inf_{K_{t}:I_{n_{x}}-K_{t}C_{t}=L}f_{1,t}(K_{t}).$
The function $f$ is convex and continuous since $f_{1,t}$ is quadratic in $K_{t}$ and we take the infimum over an affine subspace. Then (36) can be rewritten as
$\inf_{L\in\mathcal{C}}\big(\langle L^{\top}L,\Sigma_{w,t-1}\rangle+f(L)\big).$
Together with the Bures-Wasserstein constraint on $\Sigma_{w,t-1}$ in (5), this matches the setting of [17, Lemma A.3] with variable $\Sigma=\Sigma_{w,t-1}$. Hence there exists a maximizer $\Sigma_{w,t-1}^{*}$ such that $\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}}$.
Next, we fix $\Sigma_{w,t-1}$ (and thus $\Sigma_{x,t}^{-}$) and view the objective as a function of $\Sigma_{v,t}$. From (35) the dependence on $\Sigma_{v,t}$ is
$\inf_{K_{t}}\Big(\langle K_{t}^{\top}K_{t},\Sigma_{v,t}\rangle+f_{2,t}(K_{t})\Big),$
where $f_{2,t}(K_{t})$ collects the remaining terms and is convex and continuous in $K_{t}$. This is exactly of the form in [17, Lemma A.3] with variable $\Sigma=\Sigma_{v,t}$. Applying [17, Lemma A.3] again yields a maximizer $\Sigma_{v,t}^{*}$ such that $\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}$.
By choosing such partial maximizers when forming a joint maximizer of (5), we obtain (34), completing the proof.
Appendix B Proofs for Spectral Boundedness
B.1 Proof of Proposition˜1
Proof.
By Lemma˜2, there exists $U\in O(n)$ such that $\|X^{1/2}-\hat{X}^{1/2}U\|_{F}\leq\theta.$ Since $\|\cdot\|_{2}\leq\|\cdot\|_{F}$, we also have
$\|X^{1/2}-\hat{X}^{1/2}U\|_{2}\leq\theta.$ (37)
For any matrices $A,B\in\real{n\times n}$, the singular-value perturbation bounds [41, Corollary 7.3.5] state that
$\begin{split}&|\sigma_{\max}(A)-\sigma_{\max}(B)|\leq\|A-B\|_{2},\\ &|\sigma_{\min}(B)-\sigma_{\min}(A)|\leq\|A-B\|_{2}.\end{split}$ (38)
Apply (LABEL:eqn:sig_per) with $A=X^{1/2}$ and $B=\hat{X}^{1/2}U$, and combine with (37) to obtain
$\begin{split}&|\sigma_{\max}(X^{1/2})-\sigma_{\max}(\hat{X}^{1/2}U)|\leq\theta,\\ &|\sigma_{\min}(\hat{X}^{1/2}U)-\sigma_{\min}(X^{1/2})|\leq\theta\end{split}$
Because $U$ is orthogonal, $X\succeq 0$ and $\hat{X}\succeq 0$, we have
[…]
Substituting these into the inequalities above yields
$\sqrt{\lambda_{\max}(X)}=\sigma_{\max}(X^{1/2})\leq\sqrt{\lambda_{\max}(\hat{X})}+\theta,$
$\sqrt{\lambda_{\min}(X)}=\sigma_{\min}(X^{1/2})\geq\sqrt{\lambda_{\min}(\hat{X})}-\theta$
Squaring both sides and clipping the lower bound at zero yields the claim. ∎
B.2 Proof of Corollary˜2
Proof.
By Lemma˜1, the matrices $\Sigma_{v,t}^{*},\Sigma_{w,t}^{*}$ and $\Sigma_{x,0}^{-,*}$ belong to the Bures–Wasserstein balls centered at $\hat{\Sigma}_{v,t},\hat{\Sigma}_{w,t}$ and $\hat{\Sigma}_{x,0}^{-}$ with radii $\theta_{v},\theta_{w}$ and $\theta_{x_{0}}$, respectively. Hence, Proposition˜1 gives the corresponding upper bounds.
The lower bounds follow directly from the constraints
$\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}},\quad\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\quad\Sigma_{x,0}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},$
which are imposed in (4) and (5). Thus, any maximizers $\Sigma_{v,t}^{*}$, $\Sigma_{w,t-1}^{*}$, and $\Sigma_{x,0}^{-,*}$ necessarily satisfy these inequalities, giving the claimed lower bounds. ∎
B.3 Proof of Theorem˜2
Proof.
Define the posterior and prediction maps
$\begin{split}&\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t})\coloneqq\Sigma_{x,t}^{-}-\Sigma_{x}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t})^{-1}C_{t}\Sigma_{x,t}^{-}\\ &\Phi(\Sigma_{x,t},\Sigma_{w,t})\coloneqq A_{t}\Sigma_{x,t}A_{t}^{\top}+\Sigma_{w,t}.\end{split}$
The DRKF recursion is obtained by evaluating these maps at the least-favorable noise covariances:
$\Sigma_{x,t}=\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t}^{*})\quad\text{and}\quad\Sigma_{x,t+1}^{-}=\Phi(\Sigma_{x,t},\Sigma_{w,t}^{*}).$
Using the identity $\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t})=((\Sigma_{x,t}^{-})^{-1}+C_{t}^{\top}\Sigma_{v,t}^{-1}C_{t})^{-1}$ and the fact that $P_{1}\succeq P_{2}$ implies $P_{1}^{-1}\preceq P_{2}^{-1}$, the map $\Psi$ is monotone increasing in each argument. The map $\Phi$ is also monotone since congruence with $A_{t}$ and addition of a PSD matrix preserve the Loewner order.
Corollary˜2, together with the monotonicity of the posterior and prediction maps $\Psi$ and $\Phi$, implies by induction that the KF driven by $(\underline{\lambda}_{w,t}I_{n_{x}},\underline{\lambda}_{v,t}I_{n_{y}})$ produces covariances no larger than those of the DRKF, and the KF driven by $(\overline{\lambda}_{w,t}I_{n_{x}},\overline{\lambda}_{v,t}I_{n_{y}})$ produces covariances no smaller. ∎
Appendix C Proofs for Convergence
C.1 Details of the Downsampled System
For the convergence analysis, we define a downsampled state as $x_{t}^{d}=x_{tN}$ for some $N\in\mathbb{N}$. Then, the downsampled dynamics take the form
$\begin{split}x_{t+1}^{d}&=A^{N}x_{t}^{d}+\mathcal{R}_{N}\bm{w}_{t}^{N}\\ y_{t}^{N}&=\mathcal{O}_{N}x_{t}^{d}+\mathcal{D}_{N}\bm{v}_{t}^{N}+\mathcal{H}_{N}\bm{w}_{t}^{N}\\ \bm{0}&=\mathcal{O}_{N}^{R}x_{t}^{d}+\bm{u}_{t}^{N}+\mathcal{L}_{N}\bm{w}_{t}^{N},\end{split}$
where the block matrices collecting process and measurement noise are
$\begin{split}\mathcal{R}_{N}&\coloneqq\left[\hat{\Sigma}_{w}^{\frac{1}{2}},\,A\hat{\Sigma}_{w}^{\frac{1}{2}},\,\ldots,\,A^{N-1}\hat{\Sigma}_{w}^{\frac{1}{2}}\right]\\ \mathcal{O}_{N}&\coloneqq\left[(CA^{N-1})^{\top},\,\ldots,\,(CA)^{\top},\,C^{\top}\right]^{\top}\\ \mathcal{O}_{N}^{R}&\coloneqq\left[(A^{N-1})^{\top},\,\ldots,\,(A)^{\top},\,I_{n_{x}}\right]^{\top}\\ \mathcal{D}_{N}&\coloneqq I_{N}\otimes\hat{\Sigma}_{v}^{\frac{1}{2}}.\end{split}$
Here, $\mathcal{R}_{N}$ and $\mathcal{O}_{N}$ represent the $N$-block controllability and observability matrices. The block Hankel matrices $\mathcal{L}_{N}$ and $\mathcal{H}_{N}\coloneqq(I_{N}\otimes C)\mathcal{L}_{N}$ encode cross-correlations between observations and noise terms, with entries of $\mathcal{L}_{N}$ defined as $\mathcal{L}_{N}]_{ij}=A^{j-i-1}\hat{\Sigma}_{w}^{\frac{1}{2}}$ for $j>i$ and zero otherwise.
Adopting the procedure in [36] and denoting the downsampled prior covariance matrix by $\Sigma_{x,t}^{-,d}:=\Sigma_{x,tN}^{-}$, the DR Riccati recursion associated with the downsampled model is given by (23), where
$r_{\mathbb{D},t}^{d}(\Sigma_{x,t}^{-,d})\coloneqq\alpha_{N,t}\left[(\Sigma_{x,t}^{-,d})^{-1}+\Omega_{\bar{\Phi}_{N,t}}\right]^{-1}\alpha_{N,t}+W_{\bar{\Phi}_{N,t}},$
with auxiliary matrices given by
$\alpha_{N,t}$ $\coloneqq A^{N}-\mathcal{R}_{N}(\mathcal{H}_{N}^{\top}\mathcal{K}_{{\bar{\Phi}}_{N,t}}^{-1}\mathcal{O}_{N}+\mathcal{L}_{N}^{\top}\mathcal{K}_{{\bar{\Phi}}_{N,t}}^{-1}\mathcal{O}_{N}^{R})$
$W_{\bar{\Phi}_{N,t}}$ $\coloneqq\mathcal{R}_{N}\mathcal{Q}_{\bar{\Phi}_{N,t}}\mathcal{R}_{N}^{\top}$
$\mathcal{Q}_{\bar{\Phi}_{N,t}}$ $\coloneqq\left[I_{Nn_{x}}+\mathcal{H}_{N}^{\top}(\mathcal{D}_{N}\mathcal{D}_{N}^{\top})^{-1}\mathcal{H}_{N}-\mathcal{L}_{N}^{\top}\bar{\Phi}_{N,t}\mathcal{L}_{N}\right]^{-1}$
$\Omega_{\bar{\Phi}_{N,t}}$ $\coloneqq\Omega_{N}+\mathcal{J}_{N}^{\top}S_{\bar{\Phi}_{N,t}}^{-1}\mathcal{J}_{N}$
$\mathcal{J}_{N}$ $\coloneqq\mathcal{O}_{N}^{R}-\mathcal{L}_{N}\mathcal{H}_{N}^{\top}\left[\mathcal{D}_{N}\mathcal{D}_{N}^{\top}+\mathcal{H}_{N}\mathcal{H}_{N}^{\top}\right]^{-1}\mathcal{O}_{N}$
$\Omega_{N}$ $\coloneqq\mathcal{O}_{N}^{\top}(\mathcal{D}_{N}\mathcal{D}_{N}^{\top}+\mathcal{H}_{N}\mathcal{H}_{N}^{\top})^{-1}\mathcal{O}_{N}$
$S_{\bar{\Phi}_{N,t}}$ $\coloneqq\mathcal{L}_{N}(I_{Nn_{x}}+\mathcal{H}_{N}^{\top}(\mathcal{D}_{N}\mathcal{D}_{N}^{\top})^{-1}\mathcal{H}_{N})^{-1}\mathcal{L}_{N}^{\top}-\bar{\Phi}_{N,t}^{-1}$
$\mathcal{K}_{\bar{\Phi}_{N,t}}$ $\coloneqq\begin{bmatrix}\mathcal{D}_{N}\mathcal{D}_{N}^{\top}+\mathcal{H}_{N}\mathcal{H}_{N}^{\top}&\mathcal{H}_{N}\mathcal{L}_{N}^{\top}\\ \mathcal{L}_{N}\mathcal{H}_{N}^{\top}&\mathcal{L}_{N}\mathcal{L}_{N}^{\top}-\bar{\Phi}_{N,t}^{-1}\end{bmatrix}$.
C.2 Proof of Proposition˜2
Proof.
The proof follows the argument in [36, Proposition 3.1]. The downsampled DR Riccati mapping in (23) has the same structure as the robust Riccati mapping analyzed in [42, Theorem 5.3]. According to this result, if the matrices $\Sigma_{x,t}^{-,d},\Omega_{\bar{\Phi}_{N,t}}$, and $W_{\bar{\Phi}_{N,t}}$ are positive definite, then $r_{\mathbb{D},t}^{d}$ is a contraction mapping.
By Assumption˜1, $\hat{\Sigma}_{w}\succ 0$, which ensures that $\Sigma_{x,t}^{-,d}\succ 0$. Moreover, $\mathcal{Q}_{N,t}\succ 0$ whenever $0\preceq\bar{\Phi}_{N,t}\prec\tilde{\phi}_{N}I_{Nn_{x}}$, which, under the controllability assumption, is sufficient for the positive definiteness of $W_{\bar{\Phi}_{N,t}}$ for $N\geq n_{x}$.
Similarly, $S_{\bar{\Phi}_{N,t}}\prec 0$ and the map $\bar{\Phi}_{N,t}\mapsto\Omega_{\bar{\Phi}_{N,t}}$ is monotonically non-increasing for $0\prec\bar{\Phi}_{N,t}\prec\tilde{\phi}_{N}I_{Nn_{x}}$. Under the observability assumption, the matrix $\Omega_{0}=\Omega_{N}$ is positive definite for $N\geq n_{x}$. Hence, by continuity, there exists $\phi_{N}\in(0,\tilde{\phi}_{N})$ such that $\Omega_{\bar{\Phi}_{N,t}}\succ 0$ for all $\bar{\Phi}_{N,t}$ satisfying (24).
Thus all three matrices $\Sigma_{x,t}^{-,d},W_{\bar{\Phi}_{N,t}}$, and $\Omega_{\bar{\Phi}_{N,t}}$ are positive definite under (24), and therefore $r_{\mathbb{D},t}^{d}$ is a contraction for $t\geq q$. Since $r_{\mathbb{D}}$ is the corresponding time-unrolled operator, it inherits the same contraction property. ∎
C.3 Operator-norm bounds in a Bures–Wasserstein ball
We bound the operator-norm deviation between a feasible covariance matrix and its nominal counterpart when both lie in a Bures–Wasserstein ball.
Lemma 4.
Fix $\hat{X}\succ 0$. For any $X\succ 0$ with $\mathcal{B}(X,\hat{X})\leq\theta$,
$\|X-\hat{X}\|_{2}\;\leq\;\theta^{2}+2\theta\sqrt{\lambda_{\max}(\hat{X})}$.
Proof.
By Lemma˜2, there exists $U\in O(n)$ such that
$\|X^{1/2}-\hat{X}^{1/2}U\|_{2}\leq\|X^{1/2}-\hat{X}^{1/2}U\|_{F}\leq\theta.$
Let $E:=X^{1/2}-\hat{X}^{1/2}U$, so that $X^{1/2}=E+\hat{X}^{1/2}U$. Then,
$X=\hat{X}+EE^{\top}+\hat{X}^{1/2}UE^{\top}+EU^{\top}\hat{X}^{1/2},$
since $U$ is orthogonal. Taking operator norms and applying the triangle inequality and submultiplicativity of $|\cdot|_{2}$, we obtain
$\|X-\hat{X}\|_{2}\leq\|E\|_{2}^{2}+2\|\hat{X}^{1/2}\|_{2}\|E\|_{2}.$
Since $\|E\|_{2}\leq\theta$ and $\|\hat{X}^{1/2}\|_{2}=\sqrt{\lambda_{\max}(\hat{X})}$, the result follows. ∎
C.4 Proof of Theorem˜3
Proof.
From (21) and the matrix inversion lemma,
$\begin{split}\Phi_{t}&=C^{\top}\bigl(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\bigr)C+X_{t}-(I+X_{t}\Delta_{t})^{-1}X_{t}\\ &=C^{\top}(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1})C+X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}.\end{split}$
We bound the two terms separately.
$(i)$ Measurement uncertainty bound. Let $\delta_{v}\coloneq\|\Sigma_{v,t}^{*}-\hat{\Sigma}_{v}\|_{2}$. Since both matrices are invertible,
$\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}=\hat{\Sigma}_{v}^{-1}(\Sigma_{v,t}^{*}-\hat{\Sigma}_{v})(\Sigma_{v,t}^{*})^{-1}.$
By submultiplicativity, we have that
$\|\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\|_{2}\leq\delta_{v}\|\hat{\Sigma}_{v}^{-1}\|_{2}\|(\Sigma_{v,t}^{*})^{-1}\|_{2}.$ (39)
On the other hand, Weyl’s inequality gives
$\lambda_{\min}(\Sigma_{v,t}^{*})\geq\lambda_{\min}(\hat{\Sigma}_{v})-\|\Sigma_{v,t}^{*}-\hat{\Sigma}_{v}\|_{2}=\lambda_{\min}(\hat{\Sigma}_{v})-\delta_{v},$
so for $\delta_{v}<\lambda_{\min}(\hat{\Sigma}_{v})$,
$\|\hat{\Sigma}_{v}^{-1}\|_{2}=\frac{1}{\lambda_{\min}(\hat{\Sigma}_{v})},$ (40)
and
$\|(\Sigma_{v,t}^{*})^{-1}\|_{2}\ =\frac{1}{\lambda_{\min}(\Sigma_{v,t}^{*})}\leq\frac{1}{\lambda_{\min}(\hat{\Sigma}_{v})-\delta_{v}}.$ (41)
Substituting into (39) yields
$\|\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\|_{2}\leq\frac{\delta_{v}}{\lambda_{\min}(\hat{\Sigma}_{v})\,(\lambda_{\min}(\hat{\Sigma}_{v})-\delta_{v})}.$
Congruence with $C$ then gives
$\bigl\|C^{\top}(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1})C\bigr\|_{2}\leq\|C\|_{2}^{2}\,\|\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\|_{2}\leq\frac{\|C\|_{2}^{2}\,\delta_{v}}{\lambda_{\min}(\hat{\Sigma}_{v})\,(\lambda_{\min}(\hat{\Sigma}_{v})-\delta_{v})}.$ (42)
To guarantee $C^{\top}(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1})C\preceq\phi_{v}I$, it is sufficient that the right-hand side of (42) does not exceed $\phi_{v}$. This requirement is equivalent to $\delta_{v}\leq\delta_{v}^{\max}(\phi_{v})$ with $\delta_{v}^{\max}(\phi_{v})$ defined in (25). Moreover, any $\delta_{v}$ satisfying this bound automatically satisfies the condition $\delta_{v}<\lambda_{\min}(\hat{\Sigma}_{v})$ used in (40) and (41), ensuring consistency of the argument.
$\delta_{v}\leq\theta_{v}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{v})}\theta_{v}.$
Thus, it is sufficient to require that this upper bound does not exceed $\delta_{v}^{\max}(\phi_{v})$, which guarantees (27). Solving the resulting quadratic inequality in $\theta_{v}$ yields the explicit condition in (26).
$(ii)$ Process uncertainty bound. Define $Y_{t}:=X_{t}^{1/2}\Delta_{t}X_{t}^{1/2}\succeq 0$. Then,
$X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}=X_{t}^{1/2}Y_{t}(I+Y_{t})^{-1}X_{t}^{1/2}.$
and since $Y_{t}\succeq 0$,
$\|Y_{t}(I+Y_{t})^{-1}\|_{2}=\lambda_{\max}(Y_{t})/(1+\lambda_{\max}(Y_{t}))=\|Y_{t}\|_{2}/(1+\|Y_{t}\|_{2}).$
Therefore,
$\|X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}\|_{2}\leq\|X_{t}^{1/2}\|_{2}^{2}\cdot\frac{\|Y_{t}\|_{2}}{1+\|Y_{t}\|_{2}}.$
However,
$\|Y_{t}\|_{2}=\|X_{t}^{1/2}\Delta_{t}X_{t}^{1/2}\|_{2}\leq\|X_{t}\|_{2}\|\Delta_{t}\|_{2},$
resulting in the bound
$\|X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}\|_{2}\leq\frac{\|X_{t}\|_{2}^{2}\,\|\Delta_{t}\|_{2}}{1+\|X_{t}\|_{2}\|\Delta_{t}\|_{2}}.$ (43)
Now, for the minimum-norm construction of $\Delta_{t}$ in (22),
$\|\Delta_{t}\|_{2}\leq\|A^{+}\|_{2}^{2}\,\|\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}\|_{2}.$ (44)
By Lemma˜4 in Appendix C.3, $\|\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}\|_{2}\leq\theta_{w}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}\theta_{w}$. On the other hand,
$\begin{split}\|X_{t}\|_{2}&\leq\|(\Sigma_{x,t}^{-})^{-1}\|_{2}+\|C\|_{2}^{2}\|(\Sigma_{v,t}^{*})^{-1}\|_{2}\\ &=\frac{1}{\lambda_{\min}(\Sigma_{x,t}^{-})}+\frac{\|C\|_{2}^{2}}{\lambda_{\min}(\Sigma_{v,t}^{*})}.\end{split}$
By Corollary˜2, $\lambda_{\min}(\Sigma_{v,t}^{*})\geq\lambda_{\min}(\hat{\Sigma}_{v})$. Moreover, since by definition
$\Sigma_{x,t}^{-}=A\Sigma_{x,t-1}A^{\top}+\Sigma_{w,t-1}^{*}\succeq\Sigma_{w,t-1}^{*},$
we have $\lambda_{\min}(\Sigma_{x,t}^{-})\geq\lambda_{\min}(\Sigma_{w,t-1}^{*})\geq\lambda_{\min}(\hat{\Sigma}_{w})$. Thus
$\|X_{t}\|_{2}\leq\delta_{x}^{\max}$ (45)
with $\delta_{x}^{\max}$ as defined in (28).
By the monotonicity of $f(x,y):=x^{2}y/(1+xy)$ on $\real{2}_{+}$, replacing $\|X_{t}\|_{2}$ and $\|\Delta_{t}\|_{2}$ in (43) by their respective upper bounds from (44) and (45) preserves the inequality. We thus obtain
$\|X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}\|_{2}\leq\frac{(\delta_{x}^{\max})^{2}\|A^{+}\|_{2}^{2}\,\bigl(\theta_{w}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}\theta_{w}\bigr)}{1+\delta_{x}^{\max}\|A^{+}\|_{2}^{2}\,\bigl(\theta_{w}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}\,\theta_{w}\bigr)}.$
Requiring the right-hand side to be $\leq\phi_{x}$ is equivalent to
$\theta_{w}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}\theta_{w}\leq\frac{\phi_{x}}{\|A^{+}\|_{2}^{2}\,\delta_{x}^{\max}\,(\delta_{x}^{\max}-\phi_{x})},$
which ensures (30). Solving the quadratic inequality yields the explicit bound (29).
Combining $(i)$ and $(ii)$ gives $\Phi_{t}\preceq\phi_{v}I+\phi_{x}I\preceq\phi_{N}I,t\geq q$. Since $\phi_{v}+\phi_{x}\leq\phi_{N}$Proposition˜2 applies: the downsampled DR Riccati map is a contraction in the Thompson part metric. By Banach’s fixed-point theorem, the recursion admits a unique fixed point, and the DR Kalman gain converges.
Appendix D Proofs for Stability and Optimality
D.1 Proof of Theorem˜4
Proof.
Let $\tilde{e}_{t}:=x_{t}-\bar{x}^{-}_{t}$ denote the prediction error. From the DRKF update,
[…]
The prediction dynamics satisfy
$\tilde{e}_{t}=Ae_{t-1}+w_{t-1}-\hat{w}$.
Eliminating $\tilde{e}_{t}$ yields the posterior error recursion
[…]
At steady state, the DRKF covariances satisfy
$\Sigma_{x,\infty}^{-,*}=A\Sigma_{x,\infty}^{*}A^{\top}+\Sigma_{w,\infty}^{*}$
with
$K_{\infty}=\Sigma_{x,\infty}^{-,*}C^{\top}\!\bigl(C\,\Sigma_{x,\infty}^{-,*}C^{\top}+\Sigma_{v,\infty}^{*}\bigr)^{-1}.$
Using the Joseph form,
$\Sigma_{x,\infty}^{*}\;=\;(I-K_{\infty}C)\,\Sigma_{x,\infty}^{-,*}\,(I-K_{\infty}C)^{\top}+K_{\infty}\Sigma_{v,\infty}^{*}K_{\infty}^{\top}.$
Substituting $\Sigma_{x,\infty}^{-,*}=A\Sigma_{x,\infty}^{*}A^{\top}+\Sigma_{w,\infty}^{*}$ gives
$\Sigma_{x,\infty}^{*}$ $=F_{\infty}\,\Sigma_{x,\infty}^{*}\,F_{\infty}^{\top}+(I-K_{\infty}C)\,\Sigma_{w,\infty}^{*}\,(I-K_{\infty}C)^{\top}+K_{\infty}\Sigma_{v,\infty}^{*}K_{\infty}^{\top}.$
Rearranging,
$\Sigma_{x,\infty}^{*}-F_{\infty}\Sigma_{x,\infty}^{*}F_{\infty}^{\top}=(I-K_{\infty}C)\,\Sigma_{w,\infty}^{*}\,(I-K_{\infty}C)^{\top}+K_{\infty}\Sigma_{v,\infty}^{*}K_{\infty}^{\top}\succeq 0.$
Since $\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*}\succ 0$, the right-hand side is in fact $\succ 0$, implying the strict Lyapunov inequality $\Sigma_{x,\infty}^{*}\succ F_{\infty}\Sigma_{x,\infty}^{*}F_{\infty}^{\top}.$ In fact, it implies that $F_{\infty}$ is Schur stable. Indeed, if $F_{\infty}$ had an eigenvalue on or outside the unit circle, no $\Sigma_{x,\infty}^{*}\succ 0$ could satisfy $\Sigma_{x,\infty}^{*}\succ F_{\infty}\Sigma_{x,\infty}^{*}F_{\infty}^{\top}$. [^9] This is the standard equivalence: $F$ is Schur $\Leftrightarrow$ there exists $P\succ 0$ with $P-FPF^{\top}\succ 0$.
With $F_{\infty}$ Schur, the homogeneous error dynamics $e_{t}^{0}=F_{\infty}e_{t-1}^{0}$ thus converge to 0. With noises, the covariance of $e_{t}$ under the least-favorable distributions obeys the steady-state equation derived above and thus converges to $\Sigma_{x,\infty}^{*}$. Taking traces gives the claimed convergence and boundedness.
D.2 Proof of Corollary˜5
Proof.
Taking expectations in (46) yields
[…]
Since $F_{\infty}$ is Schur by Theorem˜4, the affine recursion converges, yielding
[…]
If $\hat{w}=\mu_{w}$ and $\hat{v}=\mu_{v}$, the constant input term vanishes, and thus $m_{t}\to 0$. ∎
D.3 Proof of Theorem˜5
Proof.
By the strong duality of Wasserstein DRO and the orthogonality principle for quadratic losses, the one-step minimax problem admits a saddle point $(\psi_{t}^{*},\mathbb{P}_{e,t}^{*})$ such that
$\inf_{\psi_{t}\in\mathcal{F}_{t}}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})=\operatorname{Tr}[\Sigma_{x,t}^{*}],$
where $\Sigma_{x,t}^{*}=\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t}^{*})$ is the DRKF posterior covariance.
Thus, for any admissible $\psi_{t}$,
$\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})\geq\operatorname{Tr}[\Sigma_{x,t}^{*}],$
almost surely. Since $\Sigma_{x,t}^{*}\to\Sigma_{x,\infty}^{*}$ by Theorem˜3, both lower bounds (31) and (32) follow. For the steady-state DRKF $\psi_{\infty}$ with gain $K_{\infty}$, the least-favorable posterior covariance is exactly $\Sigma_{x,\infty}^{*}$. Hence, the worst-case one-step MSE equals $\operatorname{Tr}[\Sigma_{x,\infty}^{*}]$ for all sufficiently large $t$, establishing equality in the pointwise limit and in the long-run average. ∎

### v3

@@
 Abstract
-In this work, we propose a noise-centric formulation of the distributionally robust Kalman filter (DRKF) for discrete-time linear stochastic systems with uncertain noise statistics.
-By placing Wasserstein ambiguity sets directly on the process and measurement noise distributions, the proposed DRKF preserves the analytical structure of the classical Kalman filter while providing a priori spectral bounds on all feasible covariances.
-In the time-invariant setting, we derive a steady-state DRKF from a single stationary semidefinite program, yielding a constant-gain estimator with the same per-step computational complexity as the standard Kalman filter.
-We establish conditions guaranteeing the existence, uniqueness, and convergence of this steady-state solution, and we prove its asymptotic minimax optimality with respect to the worst-case mean-square error.
-Numerical experiments validate the theory and demonstrate that the proposed DRKF improves estimation accuracy under unknown or uncertain noise models while offering computational advantages over existing robust and distributionally robust filters.
+We study state estimation for discrete-time linear stochastic systems under distributional ambiguity in the initial state, process noise, and measurement noise.
+We propose a noise-centric distributionally robust Kalman filter (DRKF) based on Wasserstein ambiguity sets imposed directly on these distributions.
+This formulation excludes dynamically unreachable priors and yields a Kalman-type recursion driven by least-favorable covariances computed via semidefinite programs (SDP).
+In the time-invariant case, the steady-state DRKF is obtained from a single stationary SDP, producing a constant gain with Kalman-level online complexity.
+We establish the convergence of the DR Riccati covariance iteration to the stationary SDP solution, together with an explicit sufficient condition for a prescribed convergence rate.
+We further show that the proposed noise-centric model induces a priori spectral bounds on all feasible covariances and a Kalman filter sandwiching property for the DRKF covariances.
+Finally, we prove that the steady-state error dynamics are Schur stable, and the steady-state DRKF is asymptotically minimax optimal with respect to worst-case mean-square error.
 1 Introduction
-State estimation is a central problem in control, perception, signal processing, and robotics, where the goal is to infer system states from noisy and partial measurements.
-The Kalman filter (KF) [1] is optimal for linear–Gaussian systems when the noise statistics are known.
-In practice, however, these statistics are rarely available a priori; they are estimated from limited data, subject to modeling error, and often vary over time.
-Even mild misspecification can degrade estimation accuracy, while severe mismatch may lead to filter divergence.
-To address this sensitivity, several robust estimation frameworks have been proposed.
-The $H_{\infty}$ filter [2] bounds the disturbance-to-error energy gain and guarantees finite error under worst-case noise, but it is often conservative in stochastic settings due to its pessimistic nature.
-Risk-sensitive filters [3, 4] penalize large errors exponentially, but their performance depends on a risk parameter that lacks a clear statistical interpretation and is difficult to tune, frequently resulting in excessive conservatism or instability.
-Recent advances in distributionally robust optimization (DRO) have inspired the development of distributionally robust state estimators (DRSE).
-Although distributional robustness has already evolved into a rich research area in control design (e.g., [5, 6, 7, 8, 9, 10, 11]), its exploration in state estimation remains relatively limited.
-DRSE methods generally assume that the true noise distributions lie within an ambiguity set centered at a nominal distribution and seek estimators that minimize the worst-case expected error over this set.
-Within this framework, a variety of ambiguity sets have been considered, including $\phi$-divergence balls, moment-based ambiguity sets [12, 13, 14, 15], and Wasserstein balls [16, 17, 18, 19, 20].
-Importantly, DRSE formulations differ not only in the type of ambiguity set employed, but also in the distributions to which the ambiguity is applied.
-For example, [16] imposes a Wasserstein ambiguity set on the joint state–measurement distribution; [17] places separate ambiguity sets on the prior state and measurement-noise covariances; and [19] introduces a bicausal optimal-transport ambiguity set that explicitly encodes temporal causality.
-Despite this progress, steady-state formulations of DRSE remain underdeveloped.
-Analyzing the infinite-horizon regime is crucial for long-term operation and computational tractability: time-varying DRKFs typically require solving an optimization problem at every time step or a large optimization problem whose size grows with the horizon.
-The frequency-domain minimax formulation of [20] characterizes the optimal linear time-invariant estimator, but the resulting filter is non-rational and does not admit a finite-dimensional state-space realization.
-Our previous work [21] developed a steady-state DRKF that placed ambiguity sets on the prior state and measurement-noise distributions.
-In this paper, we instead adopt a noise-centric formulation, placing Wasserstein ambiguity sets directly on the initial state, process noise, and measurement noise distributions.
-Unlike existing approaches (e.g., [20]), the resulting steady-state DRKF is obtained from a single stationary semidefinite program (SDP) that directly returns a constant-gain solution.
-This noise-centric formulation ensures a form of dynamic consistency—all admissible priors are generated by propagating the system dynamics under feasible noise sequences—and enables several theoretical and computational advantages.
+State estimation is a fundamental problem in systems and control, where the objective is to infer the system state from noisy and partial measurements.
+For linear systems with Gaussian noise and known statistics, the Kalman filter (KF) [12] provides the optimal minimum mean-square error (MMSE) estimator.
+In practice, however, noise statistics are rarely known a priori: they are typically learned from limited data, subject to modeling error, and may vary over time.
+Even moderate misspecification can significantly degrade estimation performance, while severe mismatch may lead to divergence.
+To mitigate this sensitivity, several robust estimation frameworks have been developed.
+The $H_{\infty}$ filter [27] guarantees bounded estimation error under worst-case disturbances, but it is often conservative in stochastic settings.
+Risk-sensitive filters (e.g., [28] and [32]) penalize large errors exponentially via an entropic risk measure, but their performance relies on an accurate specification of the underlying noise distributions, which may be difficult to justify in practice.
+In contrast, recent advances in distributionally robust optimization (DRO) have motivated growing interest in distributionally robust state estimation (DRSE).
+In DRSE, the true noise distributions are assumed to lie in an ambiguity set centered at a nominal model, and the estimator minimizes the worst-case expected error over this set.
+While distributional robustness has been extensively studied in control design (e.g., [29, 33, 25, 17, 7, 19, 3]), its application to state estimation remains comparatively limited.
+Existing DRSE formulations vary along two key dimensions: (i) the geometry of the ambiguity set, such as $\phi$-divergence balls, moment-based sets [30, 31, 35, 15], or Wasserstein balls [26, 23, 18, 8, 13]; and (ii) the location where ambiguity is imposed, e.g., on the joint state–measurement distribution [26], on the prior state and measurement noise [23], or via bicausal optimal transport models that encode temporal dependence [8].
+These choices have significant implications for tractability, interpretability, and long-term behavior.
+Despite this progress, steady-state DRSE remains underdeveloped.
+For long-horizon operation, constant-gain estimators with fixed offline complexity are essential.
+However, many existing time-varying DRKFs require solving an optimization problem at each time step, or lead to horizon-dependent formulations whose size grows with time.
+While the frequency-domain minimax approach of [13] characterizes an optimal linear time-invariant estimator, the resulting filter is generally non-rational and does not admit a finite-dimensional state-space realization.
+The preliminary version [11] of this work derived a steady-state DRKF by placing ambiguity sets on the prior state and measurement noise distributions.
+In this paper, we substantially extend this line of work by adopting a noise-centric formulation, in which Wasserstein ambiguity sets are imposed directly on the initial state, process noise, and measurement noise distributions.
+This modeling choice enforces a form of dynamic consistency: all admissible prior distributions arise through propagation of the system dynamics under admissible noise realizations.
+Moreover, it admits a finite-dimensional steady-state characterization and yields a constant-gain DRKF from a single stationary semidefinite program (SDP), while providing several theoretical and computational advantages.
 Our main contributions are summarized as follows.
   • Noise-centric ambiguity modeling.
-Wasserstein ambiguity sets are defined directly on the initial state, process noise, and measurement noise distributions.
-This design $(i)$ enforces a form of dynamic consistency by ensuring that all admissible priors are reachable through the system dynamics under admissible noise realizations; $(ii)$ provides a priori spectral bounds on the DRKF covariance recursion, unlike ambiguity sets using joint state–measurement laws, which may include unreachable priors and typically lack such spectral guarantees; and $(iii)$ yields a tractable time-varying DRKF whose stage-wise gains are precomputed via small convex SDPs and applied online with Kalman-level per-step complexity.
-  • Steady-state DRKF via a single SDP.
-A constant-gain DRKF is obtained by solving one stationary SDP offline.
-Its online computational cost matches that of the classical KF.
-The resulting steady-state covariances solve a time-invariant convex program, and the gains of the time-varying DRKF converge to this stationary gain.
-  • Theoretical guarantees.
-$(i)$ Spectral boundedness and KF sandwiching: Wasserstein balls impose explicit eigenvalue bounds on feasible covariances, ensuring that the DRKF’s prior and posterior covariances lie between those of two appropriately chosen KFs.
-This property prevents arbitrarily inflated or singular covariances under worst-case deviations in the noise statistics.
-$(ii)$ Explicit convergence radii: we derive closed-form sufficient conditions on the ambiguity radii guaranteeing contraction of the DR Riccati recursion and convergence of the time-varying DRKF to its steady-state counterpart.
-$(iii)$ Stability and optimality: under standard controllability and observability assumptions, the estimation error dynamics are Schur stable, and the steady-state DRKF achieves asymptotic minimax mean-square error (MSE) optimality within the specified ambiguity sets.
-The remainder of this paper is organized as follows.
-Section˜2 reviews preliminaries on DRSE and Wasserstein ambiguity sets.
-Section˜3 presents the time-varying and steady-state DRKF algorithms and establishes their key properties.
-Section˜4 develops convergence, stability, and optimality results.
-Section˜5 demonstrates the performance and utility of the proposed framework through numerical experiments.
+We define Wasserstein ambiguity sets directly on the initial state, process noise, and measurement noise distributions.
+This modeling choice enforces dynamic consistency, ensuring that all admissible prior distributions arise through propagation of the system dynamics under admissible noise realizations, and forms the structural basis for the subsequent theoretical and algorithmic developments.
+  • Steady-state DRKF via a single stationary SDP.
+In the time-invariant setting, we obtain a constant-gain steady-state DRKF by solving a single stationary SDP offline, yielding Kalman-level per-step online complexity.
+  • Convergence with explicit rate conditions.
+We establish the convergence of the DR Riccati iteration governing the state-error covariance to a unique steady-state solution characterized by the stationary SDP.
+Moreover, we derive an explicit sufficient condition on the ambiguity radii that guarantees a prescribed contraction rate.
+  • Spectral structure and steady-state guarantees.
+We show that Wasserstein ambiguity induces explicit a priori eigenvalue bounds (an eigenvalue tube) and a KF-sandwich property for the DRKF covariances.
+We further establish Schur stability of the steady-state estimation error dynamics and asymptotic minimax optimality with respect to the worst-case mean-square error (MSE).
+The remainder of the paper is organized as follows.
+Section˜2 introduces the problem formulation and Wasserstein ambiguity sets.
+Section˜3 presents the finite-horizon and steady-state DRKF constructions and their SDP reformulations.
+Section˜4 develops spectral boundedness, convergence, stability, and optimality results.
+Section˜5 illustrates the performance of the proposed methods through numerical experiments.
 2 Problem Setup
-We begin by introducing notation.
-Let $\mathbb{S}^{n}$ denote the set of symmetric matrices in $\real{n\times n}$, with $\mathbb{S}^{n}_{+}$ and $\mathbb{S}^{n}_{++}$ denoting the positive semidefinite and positive definite cones.
-For $A,B\in\mathbb{S}^{n}$, the relations $A\succeq B$ and $A\succ B$ indicate that $A-B\in\mathbb{S}^{n}_{+}$ and $A-B\in\mathbb{S}^{n}_{++}$, respectively.
-For $A\in\real{m\times n}$, $\|A\|_{2}$ and $\|A\|_{F}$ denote the spectral and Frobenius norms.
-For $A\in\mathbb{S}^{n}$, $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denote its smallest and largest eigenvalues, and $A^{+}$ denotes its Moore–Penrose pseudoinverse.
-Let $\mathcal{P}(\mathcal{W})$ be the set of Borel probability measures supported on a set $\mathcal{W}\subseteq\real{n}$, and let $\mathcal{P}_{2}(\mathcal{W})$ denote the subset of measures with finite second-order moments.
-Given two probability measures $\mathbb{P}$ and $\mathbb{Q}$, their product measure is denoted by $\mathbb{P}\otimes\mathbb{Q}$.
-For a sequence $\{w_{k}\}_{k\geq 0}$ with $w_{k}\in\real{n}$, the stacked block vector of length $N$ at time $t$ is defined as $\bm{w}_{t}^{N}\coloneqq[\,w_{tN+N-1}^{\top},\dots,w_{tN}^{\top}\,]^{\top}$.
+Notation.
+Let $\mathbb{S}^{n}$ denote the set of symmetric matrices in $\real{n\times n}$, and let $\mathbb{S}^{n}_{+}$ and $\mathbb{S}^{n}_{++}$ be the positive semidefinite and positive definite cones.
+For $A,B\in\mathbb{S}^{n}$, write $A\succeq B$ (resp.
+$A\succ B$) if $A-B\in\mathbb{S}^{n}_{+}$ (resp.
+$\mathbb{S}^{n}_{++}$).
+For a matrix AA, $\|A\|_{2}$ and $\|A\|_{F}$ denote the spectral and Frobenius norms.
+For $A\in\mathbb{S}^{n}$, $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ denote its smallest and largest eigenvalues, while $\sigma_{\min}(A)$ and $\sigma_{\max}(A)$ denote its smallest and largest singular values.
+Let $\mathcal{P}(\mathcal{W})$ be the set of Borel probability measures supported on $\mathcal{W}\subseteq\real{n}$, and let $\mathcal{P}_{2}(\mathcal{W})$ be those with finite second moments.
+For probability measures $\mathbb{P},\mathbb{Q}$, their product measure is denoted by $\mathbb{P}\otimes\mathbb{Q}$.
+Let $\mathbf{1}_{N}\in\mathbb{R}^{N}$ denote the vector of all ones, and let $\operatorname{blkdiag}(A_{1},\dots,A_{N})$ denote the block-diagonal matrix with diagonal blocks $A_{1},\dots,A_{N}$.
 2.1 Distributionally Robust State Estimation Problem
-Consider the discrete-time linear stochastic system [^1] Throughout, we assume that the system matrices $A$ and $C$ are known, and we focus exclusively on uncertainty in the noise statistics.
+Consider the discrete-time linear stochastic system [^1] Throughout, we assume that the system matrices $A_{t}$ and $C_{t}$ are known, and we focus exclusively on uncertainty in the noise statistics.
 $\begin{split}x_{t+1}&=A_{t}x_{t}+w_{t},\\ y_{t}&=C_{t}x_{t}+v_{t},\end{split}$ (1)
-where $x_{t}\in\real{n_{x}}$ and $y_{t}\in\real{n_{y}}$ denote the system state and the output.
-The process noise $w_{t}\in\real{n_{x}}$ and the measurement noise $v_{t}\in\real{n_{y}}$ follow distributions $\mathbb{Q}_{w,t}\in\mathcal{P}(\real{n_{x}})$ and $\mathbb{Q}_{v,t}\in\mathcal{P}(\real{n_{y}})$, while the initial state $x_{0}$ follows $\mathbb{Q}_{x,0}\in\mathcal{P}(\real{n_{x}})$.
-We assume that $x_{0}$, $\{w_{t}\}$, and $\{v_{t}\}$ are mutually independent, and that each of $\{w_{t}\}$ and $\{v_{t}\}$ is temporally independent; that is, $w_{t}$ and $w_{t^{\prime}}$ (and likewise $v_{t}$ and $v_{t^{\prime}}$) are independent for all $t\neq t^{\prime}$.
-[^2] Such independence assumptions are standard in the control and estimation literature; see, e.g., [22, Sec.
-2.2] and [2, Sec.
+where $x_{t}\in\real{n_{x}}$ and $y_{t}\in\real{n_{y}}$ denote the system state and output.
+The process noise $w_{t}\in\real{n_{x}}$ and measurement noise $v_{t}\in\real{n_{y}}$ follow distributions $\mathbb{Q}_{w,t}\in\mathcal{P}(\real{n_{x}})$ and $\mathbb{Q}_{v,t}\in\mathcal{P}(\real{n_{y}})$, while the initial state $x_{0}$ follows $\mathbb{Q}_{x,0}\in\mathcal{P}(\real{n_{x}})$.
+We assume that $x_{0}$, $\{w_{t}\}$, and $\{v_{t}\}$ are mutually independent, and that each of $\{w_{t}\}$ and $\{v_{t}\}$ is temporally independent.
+[^2] Such independence assumptions are standard in the control and estimation literature; see, e.g., [1, Sect.
+2.2] and [27, Sect.
 5.1].
-At each time $t\geq 0$, let $\mathcal{Y}_{t}\coloneqq\{y_{0},\dots,y_{t}\}$ denote the measurement history.
+At each time $t\geq 0$, the estimator has access to the measurement history $\mathcal{Y}_{t}\coloneqq\{y_{0},\dots,y_{t}\}$.
 The goal is to estimate $x_{t}$ given $\mathcal{Y}_{t}$.
-Since outputs arrive sequentially, a natural approach is to compute the minimum mean-square error (MMSE) estimator recursively:
+A natural criterion is the conditional minimum mean-square error (MMSE),
 $\min_{\psi_{t}\in\mathcal{F}_{t}}J_{t}(\psi_{t},\mathbb{Q}_{e,t})\coloneqq\mathbb{E}\left[\|x_{t}-\psi_{t}(\mathcal{Y}_{t})\|^{2}\mid\mathcal{Y}_{t-1}\right],$ (2)
-where $\mathcal{F}_{t}=\{\,\psi_{t}:(\real{n_{y}})^{t+1}\rightarrow\real{n_{x}}:\psi_{t}\text{ measurable}\,\}$ denotes the set of admissible estimators.
-The expectation is taken with respect to the conditional joint distribution of $(x_{t},y_{t})$ given the past observations $\mathcal{Y}_{t-1}$.
-Under (1), this conditional distribution is fully determined, for $t>0$, by the posterior distribution $\mathbb{P}_{x,t-1}$ of $x_{t-1}$ given $\mathcal{Y}_{t-1}$ and the joint noise distribution $\mathbb{Q}_{e,t}\coloneqq\mathbb{Q}_{w,t-1}\otimes\mathbb{Q}_{v,t}$ of $(w_{t-1},v_{t})$.
+where $\mathcal{F}_{t}$ denotes the set of measurable estimators $\psi_{t}:(\real{n_{y}})^{t+1}\to\real{n_{x}}$.
+That is, $\psi_{t}$ minimizes the conditional mean-square estimation error given past observations.
+For $t>0$, the conditional distribution of $(x_{t},y_{t})$ given $\mathcal{Y}_{t-1}$ is fully determined by the posterior distribution $\mathbb{P}_{x,t-1}$ of $x_{t-1}$ and the joint noise distribution $\mathbb{Q}_{e,t}\coloneqq\mathbb{Q}_{w,t-1}\otimes\mathbb{Q}_{v,t}$.
 At $t=0$, no posterior from a previous step exists.
-By slight abuse of notation, we adopt the convention $\mathcal{Y}_{-1}=\emptyset$ and set $\mathbb{P}_{x,0}^{-}\coloneqq\mathbb{Q}_{x,0}$ and $\mathbb{Q}_{e,0}\coloneqq\mathbb{P}_{x,0}^{-}\otimes\mathbb{Q}_{v,0}$, so that the definition in (2) also covers $t=0$, where the expectation reduces to an unconditional expectation with respect to the joint law of $(x_{0},v_{0})$.
+We adopt the convention $\mathcal{Y}_{-1}=\emptyset$ and define $\mathbb{P}_{x,0}^{-}\coloneqq\mathbb{Q}_{x,0}$ and $\mathbb{Q}_{e,0}\coloneqq\mathbb{P}_{x,0}^{-}\otimes\mathbb{Q}_{v,0}$, so that (2) also covers the initial stage.
 [^3] We use the superscript “-” for prior quantities; e.g., $\bar{x}^{-}_{t}$, $\Sigma_{x,t}^{-}$, and $\mathbb{P}_{x,t}^{-}$ denote the prior mean, covariance, and distribution of $x_{t}$ conditioned on $\mathcal{Y}_{t-1}$.
-When $x_{0}\sim\mathcal{N}(\hat{x}_{0}^{-},\Sigma_{x,t}^{-})$ and $\{w_{t}\},\{v_{t}\}$ are independent Gaussian sequences with known covariances, the optimal causal MMSE estimator reduces to the classical KF.
-In practice, noise statistics are estimated from limited data and thus subject to ambiguity.
-Typically only nominal distributions $\hat{\mathbb{Q}}_{w,t}$, $\hat{\mathbb{Q}}_{v,t}$, and $\hat{\mathbb{Q}}_{x,0}$ are available from identification methods [23, 24, 25, 26].
+When $x_{0}\sim\mathcal{N}(\hat{x}_{0}^{-},\Sigma_{x,0}^{-})$ and $\{w_{t}\},\{v_{t}\}$ are independent Gaussian sequences with known covariances, the optimal causal MMSE estimator reduces to the classical KF.
+In practice, however, noise statistics are typically estimated from limited data and are thus subject to ambiguity.
+In practice, only nominal distributions $\hat{\mathbb{Q}}_{w,t}$, $\hat{\mathbb{Q}}_{v,t}$, and $\hat{\mathbb{Q}}_{x,0}$ are available from identification procedures (e.g., [20, 24]).
 [^4] A hat, $\hat{\cdot}$, denotes nominal quantities; e.g., $\hat{x}_{0}^{-}$ and $\hat{\mathbb{Q}}_{x,0}$ represent the nominal initial state mean and distribution, respectively.
-Even modest deviations between nominal and true distributions can degrade performance or cause divergence, especially when disturbances accumulate over time (e.g., marginally stable or unstable systems).
-To address this, we formulate a stage-wise minimax problem in which the estimator competes against an adversary that selects distributions from ambiguity sets $\mathbb{D}_{j,t}\subset\mathcal{P}(\real{n_{j}})$ for $j\in\{w,v,x_{0}\}$.
-Each ambiguity set is centered at its corresponding nominal distribution and parameterized by a robustness radius.
-The DRSE problem at time $t$ is then given by
+Even modest deviations between nominal and true distributions can significantly degrade performance or cause divergence, especially when disturbances accumulate over time (e.g., marginally stable or unstable systems).
+To address this, we formulate a stage-wise minimax problem in which the estimator competes against an adversary that selects distributions from ambiguity sets $\mathbb{D}_{j,t}\subset\mathcal{P}(\real{n_{j}})$ for $j\in\{w,v,x\}$ to be defined later.
+The resulting DRSE problem at time $t$ is
 $\min_{\psi_{t}\in\mathcal{F}_{t}}\max_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t}),$ (3)
 where $\mathbb{D}_{e,0}\coloneqq\mathbb{D}_{x,0}\times\mathbb{D}_{v,0}$ and $\mathbb{D}_{e,t}\coloneqq\mathbb{D}_{w,t-1}\times\mathbb{D}_{v,t}$ for $t>0$.
-The corresponding joint distributions are given by $\mathbb{P}_{e,0}\coloneqq\mathbb{P}_{x,0}^{-}\otimes\mathbb{P}_{v,0}$ and $\mathbb{P}_{e,t}\coloneqq\mathbb{P}_{w,t-1}\otimes\mathbb{P}_{v,t}$.
-This product form reflects the assumed independence between process and measurement noise at each stage.
+The corresponding joint distributions are $\mathbb{P}_{e,0}\coloneqq\mathbb{P}_{x,0}^{-}\otimes\mathbb{P}_{v,0}$ and $\mathbb{P}_{e,t}\coloneqq\mathbb{P}_{w,t-1}\otimes\mathbb{P}_{v,t}$, reflecting the assumed stage-wise independence between process and measurement noise.
 Remark 1.
-In contrast to [17] and [21], which define the ambiguity set around the joint prior-measurement distribution, our formulation places ambiguity directly on the noise distributions.
-This enforces a form of dynamic consistency: priors arise only through the system dynamics (1) and admissible noise realizations.
-By contrast, DR-MMSE formulations may allow priors that are not dynamically reachable.
-Our formulation also enables the adversary to couple estimation performance across time, capturing long-term accumulation of disturbances via the state recursion, even though the ambiguity sets themselves are stage-wise.
+In contrast to formulations that place ambiguity on a joint prior state–measurement distribution (e.g., [23, 11]), our approach imposes ambiguity directly on the noise distributions.
+This enforces a form of dynamic consistency: admissible priors arise only through the system dynamics (1) under admissible noise realizations.
+By contrast, previous DRSE formulations may allow priors that are not dynamically reachable.
+Our formulation also allows the adversary to couple estimation performance across time, capturing long-term accumulation of disturbances via the state recursion, even though the ambiguity sets themselves are stage-wise.
 For systems with slow or unstable dynamics, even small biases in $w_{t}$ can lead to unbounded state growth, making this sequential robustness essential.
-While DRSE formulations have been explored in several contexts [17, 18, 16], their application to control-oriented state estimation remains limited.
-Moreover, understanding their steady-state behavior as $t\to\infty$ is critical for assessing long-term robustness and computational tractability.
 2.2 Wasserstein Ambiguity Sets
-Wasserstein ambiguity sets are widely used in the DRO literature due to their strong analytical and computational properties, including convex reformulations (e.g., [27, 28]).
-Unlike $\phi$-divergence balls, which may include statistically irrelevant distributions, Wasserstein balls account for distances between support elements and thus capture meaningful perturbations.
-This property prevents pathological distributions and leads to more reliable solutions [29].
+We adopt Wasserstein ambiguity sets due to their favorable analytical properties and their ability to exclude statistically irrelevant distributions (e.g., [21, 5]).
 The type-2 Wasserstein distance between two probability distributions $\mathbb{P},\mathbb{Q}\in\mathcal{P}(\mathcal{W})$ is defined as
 $W_{2}(\mathbb{P},\mathbb{Q})\coloneqq\inf_{\tau\in\mathcal{T}(\mathbb{P},\mathbb{Q})}\Bigg\{\left(\int_{\mathcal{W}\times\mathcal{W}}\|x-y\|^{2}\,d\tau(x,y)\right)^{\frac{1}{2}}\Bigg\},$
-where $\mathcal{T}(\mathbb{P},\mathbb{Q})$ denotes the set of joint probability measures on $\mathcal{W}\times\mathcal{W}$ with marginals $\mathbb{P}$ and $\mathbb{Q}$.
+where $\mathcal{T}(\mathbb{P},\mathbb{Q})$ denotes the set of joint probability measures with marginals $\mathbb{P}$ and $\mathbb{Q}$.
 The optimization variable $\tau$ represents the transportation plan that redistributes probability mass from $\mathbb{P}$ to $\mathbb{Q}$.
 [^5] We use the Euclidean norm to measure the transportation cost $\|x-y\|$.
-We define the ambiguity sets as Wasserstein balls centered at nominal distributions:
-$\mathbb{D}_{j,t}\coloneqq\left\{\mathbb{P}_{j,t}\in\mathcal{P}(\real{n_{j}})\,|\,W_{2}(\mathbb{P}_{j,t},\hat{\mathbb{Q}}_{j,t})\leq\theta_{j}\right\}$
-for $j\in\{w,v,x_{0}\}$, where $\theta_{j}$ denotes the Wasserstein ball radius, which specifies the level of robustness.
-[^6] Although we model the process and measurement noise uncertainties using separate ambiguity sets, the formulation can be extended to correlated noise by placing a single Wasserstein ball around their joint distribution.
-Extending the theory to such joint ambiguity sets is conceptually straightforward, but it leads to higher-dimensional optimization problems for filter design.
-In general, computing $W_{2}$ explicitly is difficult.
+We define the ambiguity sets as Wasserstein balls centered at nominal distributions: for $j\in\{w,v,x_{0}\}$,
+$\mathbb{D}_{j,t}\coloneqq\big\{\mathbb{P}_{j,t}\in\mathcal{P}(\real{n_{j}})\,|\,W_{2}(\mathbb{P}_{j,t},\hat{\mathbb{Q}}_{j,t})\leq\theta_{j}\big\},$
+where $\theta_{j}$ specifies the robustness radius.
+[^6] The formulation can be extended to correlated noise by placing a single Wasserstein ball around the joint distribution, at the cost of higher-dimensional optimization problems.
+In general, computing $W_{2}$ is difficult.
 A useful surrogate is the Gelbrich distance.
-Specifically, the Gelbrich distance between $\mathbb{P}\in\mathcal{P}(\mathbb{R}^{n})$ and $\mathbb{Q}\in\mathcal{P}(\mathbb{R}^{n})$ with mean vectors $\mu_{1},\mu_{2}\in\real{n}$ and covariance matrices $\Sigma_{1},\Sigma_{2}\in\mathbb{S}^{n}_{+}$ is defined as
+Specifically, the Gelbrich distance between $\mathbb{P}\in\mathcal{P}(\real{n})$ and $\mathbb{Q}\in\mathcal{P}(\real{n})$ with mean vectors $\mu_{1},\mu_{2}\in\real{n}$ and covariance matrices $\Sigma_{1},\Sigma_{2}\in\mathbb{S}^{n}_{+}$ is defined as
 $G(\mathbb{P},\mathbb{Q})\coloneqq\sqrt{\|\mu_{1}-\mu_{2}\|_{2}^{2}+\mathcal{B}^{2}(\Sigma_{1},\Sigma_{2})},$
-where
-$\mathcal{B}(\Sigma_{1},\Sigma_{2})\coloneqq\sqrt{\operatorname{Tr}[\Sigma_{1}+\Sigma_{2}-2\big(\Sigma_{2}^{\frac{1}{2}}\Sigma_{1}\Sigma_{2}^{\frac{1}{2}}\big)^{\frac{1}{2}}]}$
-is the Bures–Wasserstein distance.
+where $\mathcal{B}(\Sigma_{1},\Sigma_{2})\coloneqq\sqrt{\operatorname{Tr}[\Sigma_{1}+\Sigma_{2}-2\big(\Sigma_{2}^{\frac{1}{2}}\Sigma_{1}\Sigma_{2}^{\frac{1}{2}}\big)^{\frac{1}{2}}]}$ is the Bures–Wasserstein distance.
 The Gelbrich distance always provides a lower bound on the type-2 Wasserstein distance, i.e., $G(\mathbb{P},\mathbb{Q})\leq W_{2}(\mathbb{P},\mathbb{Q})$.
-Moreover, equality holds when both distributions are Gaussian, or more generally, elliptical with the same density generator [30].
+Moreover, equality holds when both distributions are Gaussian, or, more generally, elliptical with the same density generator [6].
 3 Distributionally Robust Kalman Filters
-We now develop tractable DR state estimators consistent with the noise-centric ambiguity model introduced in Section˜2.
-We first present a finite-horizon formulation whose recursion mirrors that of the KF under the least-favorable noise covariances.
-We then specialize to the time-invariant case and derive a steady-state DRKF that uses a single stationary SDP, resulting in a constant-gain filter with Kalman-level online complexity.
+We develop tractable DR state estimators consistent with the noise-centric ambiguity model in Section˜2.
+We first present a finite-horizon formulation whose recursion mirrors that of the classical KF when driven by least-favorable noise covariances.
+We then specialize to the time-invariant case and derive a steady-state DRKF obtained from a single stationary SDP, yielding a constant-gain filter with Kalman-level online complexity.
 3.1 Finite-Horizon Case
-We begin with assumptions on the nominal distributions that guarantee tractability of the DRSE problem and preserve the analytical structure of the classical KF, specifically the Gaussian and affine relationships that enable closed-form recursive updates.
+We begin with a standard assumption on the nominal distributions that ensures tractability of the DRSE problem and preserves the affine–Gaussian structure underlying closed-form KF updates.
 Assumption 1.
@@
 Suppose Assumption˜1 holds. Then, the DRSE problem (3) satisfies the following properties.
-$(i)$ At the initial stage, the following convex optimization problem has the same optimal value as (3), and its optimal solutions are optimal for (3):
-$\begin{split}\max_{\Sigma_{x,0}^{-},\Sigma_{v,0}}&\operatorname{Tr}[\Sigma_{x,0}^{-}-\Sigma_{x,0}^{-}C_{0}^{\top}(C_{0}\Sigma_{x,0}^{-}C_{0}^{\top}+\Sigma_{v,0})^{-1}C_{0}\Sigma_{x,0}^{-}]\\ \mbox{s.t.}\;&\mathrm{Tr}[\Sigma_{x,0}^{-}+\hat{\Sigma}_{x,0}^{-}-2((\hat{\Sigma}_{x,0}^{-})^{\frac{1}{2}}\Sigma_{x,0}^{-}(\hat{\Sigma}_{x,0}^{-})^{\frac{1}{2}})^{\frac{1}{2}}]\leq\theta_{x_{0}}^{2}\\ &\mathrm{Tr}[\Sigma_{v,0}+\hat{\Sigma}_{v,0}-2\big(\hat{\Sigma}_{v,0}^{\frac{1}{2}}\Sigma_{v,0}\hat{\Sigma}_{v,0}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{v}^{2}\\ &\Sigma_{x,0}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\Sigma_{v,0}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}\\ &\Sigma_{x,0}^{-}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,0}\in\mathbb{S}^{n_{y}}_{+}.\end{split}$ (4)
+$(i)$ At the initial stage, the following convex optimization problem has the same optimal value as (3), and any of its optimal solutions are optimal for (3):
+$\begin{split}\max_{\Sigma_{x,0}^{-},\Sigma_{v,0}}&\operatorname{Tr}[\Sigma_{x,0}^{-}-\Sigma_{x,0}^{-}C_{0}^{\top}(C_{0}\Sigma_{x,0}^{-}C_{0}^{\top}+\Sigma_{v,0})^{-1}C_{0}\Sigma_{x,0}^{-}]\\ \mbox{s.t.}\;&\operatorname{Tr}[\Sigma_{x,0}^{-}+\hat{\Sigma}_{x,0}^{-}-2((\hat{\Sigma}_{x,0}^{-})^{\frac{1}{2}}\Sigma_{x,0}^{-}(\hat{\Sigma}_{x,0}^{-})^{\frac{1}{2}})^{\frac{1}{2}}]\leq\theta_{x_{0}}^{2}\\ &\operatorname{Tr}[\Sigma_{v,0}+\hat{\Sigma}_{v,0}-2\big(\hat{\Sigma}_{v,0}^{\frac{1}{2}}\Sigma_{v,0}\hat{\Sigma}_{v,0}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{v}^{2}\\ &\Sigma_{x,0}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\;\Sigma_{v,0}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}\\ &\Sigma_{x,0}^{-}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,0}\in\mathbb{S}^{n_{y}}_{+}.\end{split}$ (4)
 $(ii)$ For any $t>0$, fix the state distribution $\mathbb{P}_{x,t-1}=\mathcal{N}(\bar{x}_{t-1},\Sigma_{x,t-1})$.
-Then the following convex optimization problem has the same optimal value as (3), and its optimal solutions are optimal for (3):
-$\begin{split}\max_{\Sigma_{w,t-1},\Sigma_{v,t}}\;&\operatorname{Tr}[\Sigma_{x,t}^{-}-\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t})^{-1}C_{t}\Sigma_{x,t}^{-}]\\ \mbox{s.t.}\;&\operatorname{Tr}[\Sigma_{w,t-1}+\hat{\Sigma}_{w,t-1}-2\big(\hat{\Sigma}_{w,t-1}^{\frac{1}{2}}\Sigma_{w,t-1}\hat{\Sigma}_{w,t-1}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,t}+\hat{\Sigma}_{v,t}-2\big(\hat{\Sigma}_{v,t}^{\frac{1}{2}}\Sigma_{v,t}\hat{\Sigma}_{v,t}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{v}^{2}\\ &\Sigma_{x,t}^{-}=A_{t}\Sigma_{x,t-1}A_{t}^{\top}+\Sigma_{w,t-1}\\ &\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}\\ &\Sigma_{w,t-1}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,t}\in\mathbb{S}^{n_{y}}_{+}.\end{split}$ (5)
-$(iii)$ If $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ and $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ solve (4) and (5), respectively, then the maximum in (3) is attained by the Gaussian distributions $\mathbb{P}_{x,0}^{-,*}=\mathcal{N}(\hat{x}_{0}^{-},\Sigma_{x,0}^{-,*}),\mathbb{P}_{w,t}^{*}=\mathcal{N}(\hat{w}_{t},\Sigma_{w,t}^{*})$ and $\mathbb{P}_{v,t}^{*}=\mathcal{N}(\hat{v}_{t},\Sigma_{v,t}^{*})$ for each $t\geq 0$.
-Furthermore, at any time $t$, the optimal state estimator achieving the minimum in (3) is affine and given by
-$\psi_{t}^{*}(\mathcal{Y}_{t})=\bar{x}^{-}_{t}+\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t}^{*})^{-1}(y_{t}-C_{t}\bar{x}^{-}_{t}-\hat{v}_{t}),$ (6)
-with $\bar{x}^{-}_{t}=A_{t-1}\bar{x}_{t-1}+\hat{w}_{t-1},\Sigma_{x,t}^{-}=A_{t}\Sigma_{x,t-1}A_{t}^{\top}+\Sigma_{w,t}^{*}$, and $\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*},\bar{x}^{-}_{0}=\hat{x}_{0}^{-}$.
-The problems (4) and (5) define the stage-wise minimax interaction: the estimator selects $\psi_{t}$, while the adversary selects noise covariances within Wasserstein balls.
-Specifically, problem (4) yields the least-favorable prior and measurement covariances $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ at $t=0$, whereas (5) yields the least-favorable process and measurement covariances $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ for $t\geq 1$.
+Then the following convex optimization problem has the same optimal value as (3), and any of its optimal solutions are optimal for (3):
+$\begin{split}\max_{\begin{subarray}{c}\Sigma_{w,t-1},\\ \Sigma_{v,t},\Sigma_{x,t}^{-}\end{subarray}}\;&\operatorname{Tr}[\Sigma_{x,t}^{-}-\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t})^{-1}C_{t}\Sigma_{x,t}^{-}]\\ \mbox{s.t.}\;&\operatorname{Tr}[\Sigma_{w,t-1}+\hat{\Sigma}_{w,t-1}-2\big(\hat{\Sigma}_{w,t-1}^{\frac{1}{2}}\Sigma_{w,t-1}\hat{\Sigma}_{w,t-1}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,t}+\hat{\Sigma}_{v,t}-2\big(\hat{\Sigma}_{v,t}^{\frac{1}{2}}\Sigma_{v,t}\hat{\Sigma}_{v,t}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{v}^{2}\\ &\Sigma_{x,t}^{-}=A_{t-1}\Sigma_{x,t-1}A_{t-1}^{\top}+\Sigma_{w,t-1}\\ &\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\;\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}\\ &\Sigma_{x,t}^{-},\Sigma_{w,t-1}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,t}\in\mathbb{S}^{n_{y}}_{+}.\end{split}$ (5)
+$(iii)$ If $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ and $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ solve (4) and (5), respectively, then the maximum in (3) is attained by the Gaussian distributions $\mathbb{P}_{x,0}^{-,*}=\mathcal{N}(\hat{x}_{0}^{-},\Sigma_{x,0}^{-,*}),\mathbb{P}_{w,t-1}^{*}=\mathcal{N}(\hat{w}_{t-1},\Sigma_{w,t-1}^{*})$ and $\mathbb{P}_{v,t}^{*}=\mathcal{N}(\hat{v}_{t},\Sigma_{v,t}^{*})$ for each $t$.
+Furthermore, the minimax-optimal estimator is affine and given by
+$\begin{split}\psi_{t}^{*}&(\mathcal{Y}_{t})=\bar{x}^{-}_{t}+\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t}^{*})^{-1}(y_{t}-C_{t}\bar{x}^{-}_{t}-\hat{v}_{t}),\end{split}$ (6)
+with $\bar{x}^{-}_{t}=A_{t-1}\bar{x}_{t-1}+\hat{w}_{t-1},\Sigma_{x,t}^{-}=A_{t-1}\Sigma_{x,t-1}A_{t-1}^{\top}+\Sigma_{w,t-1}^{*}$, and $\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*},\bar{x}^{-}_{0}=\hat{x}_{0}^{-}$.
+Problems (4) and (5) define the stage-wise minimax interaction: the estimator selects $\psi_{t}$, while the adversary selects noise covariances within Wasserstein balls.
+Problem (4) yields the least-favorable prior and measurement covariances $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ at $t=0$, whereas (5) yields the least-favorable process and measurement covariances $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ for $t\geq 1$.
 The corresponding optimal values represent the least-favorable posterior MSE at each stage.
-Lemma˜1 extends [17, Theorem 3.1] to the noise-centric setting by replacing ambiguity on the prior state (for $t>0$) with ambiguity on the process noise.
-Importantly, although the Wasserstein constraints are expressed in terms of first- and second-order moments, this does not restrict the adversary in a meaningful way under Assumption˜1.
-For Gaussian nominal distributions, the least-favorable distributions are Gaussian, and their worst-case behavior is fully characterized by their means and covariances.
+Lemma˜1 extends [23, Thm.
+3.1] to the noise-centric setting by replacing ambiguity on the prior state (for $t>0$) with ambiguity on the process noise.
+Although the Wasserstein constraints are equivalently expressible in terms of first and second moments, this does not meaningfully restrict the adversary under Assumption˜1.
+For Gaussian nominal distributions, the least-favorable distributions are Gaussian, and the worst-case behavior is fully characterized by their means and covariances.
 Furthermore, even without Gaussian nominal assumptions, (6) remains minimax optimal within the class of affine estimators [23].
 Remark 2.
-By adapting the structural argument used in [17, Lemma A.3], DRSE problem (3) admits at least one maximizer whose covariance matrices satisfy $\Sigma_{x,0}^{-,*}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\Sigma_{v,0}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}$, for $t=0$, and $\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}$ for all $t>0$ (see Appendix A).
-Because such a maximizer always exists, the explicit eigenvalue lower-bound constraints included in (4) and (5) are redundant: they do not alter the optimal value but simply restrict attention to maximizers that satisfy these bounds.
-In particular, the versions of (4) and (5) without these explicit bounds are equivalent to the original DRSE problem (3).
+Although the ambiguity sets are defined as $W_{2}$ balls over full distributions, the maximization problems in (4) and (5) admit least-favorable distributions that preserve the nominal noise means.
+Indeed, for any fixed distribution, the minimizing estimator is the conditional mean, and the conditional MSE depends only on the posterior covariance through its trace.
+Mean shifts consume Wasserstein budget without increasing the worst-case MSE as effectively as perturbing the covariance.
+Hence, mean shifts are suboptimal for the adversary, and the ambiguity budget is used entirely to perturb covariances, as reflected in Lemma˜1 and Theorem˜1.
 Remark 3.
-In the inner maximization problems (4) and (5), the least-favorable distributions retain the nominal noise means.
-Any shift of a noise mean would introduce a systematic bias in the estimation error while also increasing the Wasserstein distance, making such a shift never advantageous for the adversary.
-Because of this trade-off, the ambiguity budget is used entirely to perturb the covariances, which is reflected in Lemma˜1 and Theorem˜1.
+By adapting the structural argument used in [23, Lemma A.3], the DRSE problem (3) admits at least one maximizer whose covariance matrices satisfy $\Sigma_{x,0}^{-,*}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\Sigma_{v,0}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}$, for $t=0$, and $\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}$ for all $t>0$ (see Appendix A).
+Because such a maximizer always exists, the explicit eigenvalue lower-bound constraints in (4) and (5) are redundant: they do not alter the optimal value but only restrict attention to maximizers that satisfy these bounds.
+Equivalently, removing these constraints yields problems with the same optimal value as (3).
 Both problems (4) and (5) are solvable for $\hat{\Sigma}_{v,t}\in\mathbb{S}^{n_{y}}_{++}$ since their objectives are continuous over compact feasible sets.
-Applying a standard Schur complement argument (e.g., [31, Appendix 5.5]), these problems reduce to SDPs that are efficiently solvable using off-the-shelf solvers.
+Applying the standard Schur complement argument, these problems reduce to SDPs that can be solved efficiently using off-the-shelf solvers.
 Corollary 1.
 Suppose Assumption˜1 holds. Then, for any time $t>0$, the optimization problem (5) is equivalent to the following tractable convex SDP:
-$\begin{split}\max_{\begin{subarray}{c}\Sigma_{x,t}^{-},\Sigma_{x,t},\\ \Sigma_{w,t-1},\Sigma_{v,t},\\ Y,Z\end{subarray}}\;&\operatorname{Tr}[\Sigma_{x,t}]\\ \mbox{s.t.}\;&\begin{bmatrix}\Sigma_{x,t}^{-}-\Sigma_{x,t}&\Sigma_{x,t}^{-}C_{t}^{\top}\\ C_{t}\Sigma_{x,t}^{-}&C_{t}\Sigma_{x,t}^{-}C^{\top}+\Sigma_{v,t}\end{bmatrix}\succeq 0\\ &\begin{bmatrix}\hat{\Sigma}_{w,t-1}&Y\\ Y^{\top}&\Sigma_{w,t-1}\end{bmatrix}\succeq 0,\;\begin{bmatrix}\hat{\Sigma}_{v,t}&Z\\ Z^{\top}&\Sigma_{v,t}\end{bmatrix}\succeq 0\\ &\operatorname{Tr}[\Sigma_{w,t-1}+\hat{\Sigma}_{w,t-1}-2Y]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,t}+\hat{\Sigma}_{v,t}-2Z]\leq\theta_{v}^{2}\\ &\Sigma_{x,t}^{-}=A_{t}\Sigma_{x,t-1}A_{t}^{\top}+\Sigma_{w,t-1}\\ &\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}\\ &\Sigma_{x,t}^{-},\Sigma_{x,t},\Sigma_{w,t-1}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,t}\in\mathbb{S}^{n_{y}}_{+}\\ &Y\in\real{n_{x}\times n_{x}},\;Z\in\real{n_{y}\times n_{y}}.\end{split}$ (7)
+$\begin{split}\max_{\begin{subarray}{c}\Sigma_{x,t}^{-},\Sigma_{x,t},\Sigma_{w,t-1},\\ \Sigma_{v,t},Y,Z\end{subarray}}\;&\operatorname{Tr}[\Sigma_{x,t}]\\ \mbox{s.t.}\;&\begin{bmatrix}\Sigma_{x,t}^{-}-\Sigma_{x,t}&\Sigma_{x,t}^{-}C_{t}^{\top}\\ C_{t}\Sigma_{x,t}^{-}&C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t}\end{bmatrix}\succeq 0\\ &\begin{bmatrix}\hat{\Sigma}_{w,t-1}&Y\\ Y^{\top}&\Sigma_{w,t-1}\end{bmatrix}\succeq 0,\;\begin{bmatrix}\hat{\Sigma}_{v,t}&Z\\ Z^{\top}&\Sigma_{v,t}\end{bmatrix}\succeq 0\\ &\operatorname{Tr}[\Sigma_{w,t-1}+\hat{\Sigma}_{w,t-1}-2Y]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,t}+\hat{\Sigma}_{v,t}-2Z]\leq\theta_{v}^{2}\\ &\Sigma_{x,t}^{-}=A_{t-1}\Sigma_{x,t-1}A_{t-1}^{\top}+\Sigma_{w,t-1}\\ &\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\;\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}\\ &\Sigma_{x,t}^{-},\Sigma_{x,t},\Sigma_{w,t-1}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,t}\in\mathbb{S}^{n_{y}}_{+}\\ &Y\in\real{n_{x}\times n_{x}},\;Z\in\real{n_{y}\times n_{y}}.\end{split}$ (7)
 Similar results hold for $t=0$.
-Solving (7) offline for all $t$ yields the least-favorable covariances $(\Sigma_{w,t}^{*},\Sigma_{v,t}^{*})$.
+Solving (7) offline for all $t>0$ yields the least-favorable covariances $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$.
 The next theorem shows that these covariances can be used directly in a Kalman-style recursion, resulting in a DR state estimator that is robust to distributional uncertainty.
 Theorem 1 (DR Kalman Filter).
-Under Assumption˜1, the minimax DR state estimate $\psi^{*}_{t}(\mathcal{Y}_{t})$ at any time $t$ coincides with the conditional mean $\bar{x}_{t}$ under the least-favorable distributions.
+Under Assumption˜1, the minimax-optimal DR state estimate $\psi^{*}_{t}(\mathcal{Y}_{t})$ at any time $t$ coincides with the conditional mean $\bar{x}_{t}$ under the least-favorable distributions.
 Moreover, the least-favorable prior and posterior remain Gaussian: $\mathbb{P}_{x,t}^{-,*}=\mathcal{N}(\bar{x}^{-}_{t},\Sigma_{x,t}^{-})$, $\mathbb{P}_{x,t}^{*}=\mathcal{N}(\bar{x}_{t},\Sigma_{x,t})$.
@@
     $K_{t}=\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}{\Sigma_{x,t}^{-}}C_{t}^{\top}+\Sigma_{v,t}^{*})^{-1},$ (8)
-    where $\Sigma_{v,t}^{*}$ is the maximizer of (5) for $t>0$ and (4) for $t=0$. Then, update $\bar{x}_{t}$ and $\Sigma_{x,t}$ based on the measurement $y_{t}$ as follows:
+    where $\Sigma_{v,t}^{*}$ is the maximizer of (5) for $t>0$ and (4) for $t=0$. Then, update $\bar{x}_{t}$ and $\Sigma_{x,t}$ as
     $\bar{x}_{t}=K_{t}(y_{t}-C_{t}\bar{x}^{-}_{t}-\hat{v}_{t})+\bar{x}^{-}_{t}$ (9)
     $\Sigma_{x,t}=(I-K_{t}C_{t})\Sigma_{x,t}^{-}.$ (10)
-    The initial values are $\bar{x}^{-}_{0}=\hat{x}_{0}^{-},\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*}$, where $\Sigma_{x,0}^{-,*}$ is the maximizer of (4) .
+    The initial values are $\bar{x}^{-}_{0}=\hat{x}_{0}^{-},\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*}$, where $\Sigma_{x,0}^{-,*}$ is the maximizer of (4) at $t=0$.
   • (State Prediction) Predict $\bar{x}^{-}_{t+1}$ and $\Sigma_{x,t+1}^{-}$ as
@@
     $\Sigma_{x,t+1}^{-}$ $=A_{t}\Sigma_{x,t}A_{t}^{\top}+\Sigma_{w,t}^{*},$ (12)
-    where $\Sigma_{w,t}^{*}$ is the maximizer of (5) for $t>0$.
-Notably, the measurement update and state prediction equations (9)–(12) mirror those of the classical KF when the system is driven by the least-favorable distributions $\mathbb{P}_{x,0}^{-,*},\mathbb{P}_{w,t}^{*}$, and $\mathbb{P}_{v,t}^{*}$.
+    where $\Sigma_{w,t}^{*}$ is the maximizer of (5) at time $t+1$.
+Notably, the measurement update and state prediction equations (9)–(12) mirror those of the classical KF when the system is driven by the least-favorable distributions.
 3.2 Infinite-Horizon Case
-While the DRKF in Theorem˜1 provides a tractable solution to the finite-horizon DRSE problem, our focus is on its long-term behavior.
-To study this regime, we assume that both the system dynamics and the nominal uncertainty statistics are time invariant.
+While the DRKF in Theorem˜1 provides a tractable solution to the finite-horizon DRSE problem, our primary interest lies in its long-term behavior.
+To study this regime, we assume that both the system dynamics and the nominal uncertainty statistics are time-invariant.
 Assumption 2.
-The system (1) is time invariant, i.e., $A_{t}\equiv A,C_{t}\equiv C$.
+The system (1) is time-invariant, i.e., $A_{t}\equiv A,C_{t}\equiv C$.
 Moreover, the nominal uncertainty distributions are stationary, so that $\hat{w}_{t}\equiv\hat{w},\hat{v}_{t}\equiv\hat{v}$ and $\hat{\Sigma}_{v,t}\equiv\hat{\Sigma}_{v},\hat{\Sigma}_{w,t}\equiv\hat{\Sigma}_{w}$ for all $t\geq 0$, where $\hat{\Sigma}_{w}\in\mathbb{S}^{n_{x}}_{++}$ and $\hat{\Sigma}_{v}\in\mathbb{S}^{n_{y}}_{++}$.
+The nominal initial state covariance satisfies $\hat{\Sigma}_{x,0}^{-}\in\mathbb{S}^{n_{x}}_{++}$.
 Under Assumption˜2, the stage-wise SDPs (5)–(7) reduce to a time-invariant recursion for the covariance matrices.
 The resulting sequences of least-favorable noise covariances $\{\Sigma_{w,t}^{*},\Sigma_{v,t}^{*}\}$ and state covariances $\{\Sigma_{x,t}^{-},\Sigma_{x,t}\}$ may converge to limiting values $\{\Sigma_{w,\infty},\Sigma_{v,\infty}\}$ and $\{\Sigma_{x,\infty}^{-},\Sigma_{x,\infty}\}$ as $t\to\infty$.
-Whether such convergence holds depends on additional conditions, which will be established later in Section˜4.2.
-Motivated by this possibility, we seek a steady-state DR estimator with constant gain.
-In the infinite-horizon setting, the DRSE problem reduces to finding these steady-state covariances and the corresponding steady-state DR Kalman gain.
-Since the dynamics and nominal statistics are fixed, the stage-wise SDP (5) simplifies to the following time-invariant convex program:
+Whether such convergence holds depends on additional conditions, which are established later in Section˜4.2.
+Motivated by this, we seek a steady-state DR estimator with a constant gain.
+In the infinite-horizon setting, the DRSE problem reduces to computing these steady-state covariances and the associated DR Kalman gain.
+Since the dynamics and nominal statistics are time-invariant, the stage-wise SDP (5) simplifies to the following convex program:
 $\begin{split}\max_{\begin{subarray}{c}\Sigma_{w,\infty},\Sigma_{v,\infty}\\ \Sigma_{x,\infty}^{-},\Sigma_{x,\infty}\end{subarray}}\;&\operatorname{Tr}[\Sigma_{x,\infty}]\\ \mbox{s.t.}\;&\Sigma_{x,\infty}=\Sigma_{x,\infty}^{-}-\Sigma_{x,\infty}^{-}C^{\top}(C\Sigma_{x,\infty}^{-}C^{\top}+\Sigma_{v,\infty})^{-1}C\Sigma_{x,\infty}^{-}\\ &\Sigma_{x,\infty}^{-}=A\Sigma_{x,\infty}A^{\top}+\Sigma_{w,\infty}\\ &\operatorname{Tr}[\Sigma_{w,\infty}+\hat{\Sigma}_{w}-2\big(\hat{\Sigma}_{w}^{\frac{1}{2}}\Sigma_{w,\infty}\hat{\Sigma}_{w}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,\infty}+\hat{\Sigma}_{v}-2\big(\hat{\Sigma}_{v}^{\frac{1}{2}}\Sigma_{v,\infty}\hat{\Sigma}_{v}^{\frac{1}{2}}\big)^{\frac{1}{2}}]\leq\theta_{v}^{2}\\ &\Sigma_{w,\infty}\succeq\lambda_{\min}(\hat{\Sigma}_{w})I_{n_{x}},\;\Sigma_{v,\infty}\succeq\lambda_{\min}(\hat{\Sigma}_{v})I_{n_{y}}\\ &\Sigma_{x,\infty}^{-},\Sigma_{x,\infty},\Sigma_{w,\infty}\in\mathbb{S}^{n_{x}}_{+},\Sigma_{v,\infty}\in\mathbb{S}^{n_{y}}_{+}.\end{split}$
-As in (7), the optimization problem above can be reformulated as an equivalent steady-state SDP:
+As in the finite-horizon case, this problem admits an equivalent SDP formulation. In particular, it can be written as the following single stationary SDP:
 $\begin{split}\max_{\begin{subarray}{c}\Sigma_{x,\infty}^{-},\Sigma_{x,\infty},\\ \Sigma_{w,\infty},\Sigma_{v,\infty},\\ Y,Z\end{subarray}}\;&\operatorname{Tr}[\Sigma_{x,\infty}]\\ \mbox{s.t.}\;&\begin{bmatrix}\Sigma_{x,\infty}^{-}-\Sigma_{x,\infty}&\Sigma_{x,\infty}^{-}C^{\top}\\ C\Sigma_{x,\infty}^{-}&C\Sigma_{x,\infty}^{-}C^{\top}+\Sigma_{v,\infty}\end{bmatrix}\succeq 0\\ &\begin{bmatrix}\hat{\Sigma}_{w}&Y\\ Y^{\top}&\Sigma_{w,\infty}\end{bmatrix}\succeq 0,\quad\begin{bmatrix}\hat{\Sigma}_{v}&Z\\ Z^{\top}&\Sigma_{v,\infty}\end{bmatrix}\succeq 0\\ &\Sigma_{x,\infty}^{-}=A\Sigma_{x,\infty}A^{\top}+{\Sigma}_{w,\infty}\\ &\operatorname{Tr}[\Sigma_{w,\infty}+\hat{\Sigma}_{w}-2Y]\leq\theta_{w}^{2}\\ &\operatorname{Tr}[\Sigma_{v,\infty}+\hat{\Sigma}_{v}-2Z]\leq\theta_{v}^{2}\\ &\Sigma_{w,\infty}\succeq\lambda_{\min}(\hat{\Sigma}_{w})I_{n_{x}},\;\Sigma_{v,\infty}\succeq\lambda_{\min}(\hat{\Sigma}_{v})I_{n_{y}}\\ &\Sigma_{x,\infty}^{-},\Sigma_{x,\infty},\Sigma_{w,\infty}\in\mathbb{S}^{n_{x}}_{+},\;\Sigma_{v,\infty}\in\mathbb{S}^{n_{y}}_{+}\\ &Y\in\real{n_{x}\times n_{x}},\;Z\in\real{n_{y}\times n_{y}}.\end{split}$ (13)
-The optimal solution $\{\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*},\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*}\}$ defines the least-favorable stationary covariances, from which the corresponding steady-state DR Kalman gain is
+The optimal solution $\{\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*},\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*}\}$ defines the least-favorable stationary covariances and yields the steady-state DR Kalman gain
 $K_{\infty}^{*}:=\Sigma_{x,\infty}^{-,*}C^{\top}(C\Sigma_{x,\infty}^{-,*}C^{\top}+\Sigma_{v,\infty}^{*})^{-1}.$ (14)
-The resulting time-invariant state estimator is
+The resulting time-invariant estimator is
 […]
-with prior
+with the prior update
 $\bar{x}^{-}_{t+1}=A\bar{x}_{t}+\hat{w}$. (16)
 Compared with the finite-horizon DRKF, the steady-state filter (15) requires solving only a single stationary SDP offline.
-This makes it well-suited for real-time implementation.
-More importantly, it optimizes directly for worst-case asymptotic performance, ensuring robustness to persistent disturbances and long-term distributional errors.
-These properties are particularly valuable in safety-critical systems with slow or unstable dynamics.
-3.3 Algorithms
-We now summarize the proposed DRKF methods for both the time-varying and steady-state settings.
-These algorithms follow directly from the tractable reformulations in Section˜3 and the recursive structure established in Theorem˜1.
-Although both versions share the same online recursion as the classical KF, they differ in how the covariance matrices, and hence the Kalman gains, are computed to ensure robustness against distributional ambiguity.
-3.3.1 Time-Varying DRKF
-As shown in Algorithm˜1, the time-varying DRKF is implemented by first solving the stage-wise SDP problems (4) and (5) offline for each time step.
-If the nominal covariance matrices are known a priori, the least-favorable noise covariances and the DR gains $\{K_{t}\}$ can be precomputed offline for a given horizon $T$.
-At each stage, the gain $K_{t}$ is computed using the least-favorable prior state covariance and the least-favorable measurement noise covariance.
-The online stage then consists only of the measurement update and state prediction steps (9) and (11).
-Upon receiving a new measurement $y_{t}$, the update and prediction steps are carried out exactly as in the classical KF, except that they use the DR gains computed offline and the nominal noise means.
+This makes it particularly suitable for real-time implementation.
+More importantly, it directly optimizes worst-case asymptotic performance, providing robustness to persistent disturbances and long-term distributional errors–an essential property for safety-critical systems with slow or unstable dynamics.
 Algorithm 1 Time-Varying DRKF
 1:System matrices $\{A_{t},C_{t}\}_{t=0}^{T-1}$, nominal means and covariances $(\hat{w}_{t},\hat{v}_{t},\hat{\Sigma}_{w,t},\hat{\Sigma}_{v,t})$, prior state mean $\hat{x}_{0}^{-}$ and covariance $\hat{\Sigma}_{x,0}^{-}$, ambiguity radii $(\theta_{w},\theta_{v},\theta_{x_{0}})$, horizon length $T$
-2:Solve (4) to obtain $(\Sigma_{x,0}^{-*},\Sigma_{v,0}^{*})$ $\triangleright$ Offline stage
-3:Compute DR gain $K_{0}$ via (8)
+2:Solve (4) to obtain $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})\triangleright$ Offline stage
+3:Compute DR gain $K_{0}$ via (8) and update the posterior covariance $\Sigma_{x,0}$ via (10)
 4:for $t=1,\dots,T-1$ do
@@
 13:end for
-3.3.2 Steady-State DRKF
-Under the time-invariant assumptions of Section˜3.2, the DRKF recursion may converge to a stationary solution.
-In this setting, the procedure in Algorithm˜1 simplifies considerably.
-Rather than solving an SDP at every stage, one solves the single stationary SDP (13) offline to obtain the least-favorable noise covariances, the steady-state prior and posterior covariances, and the resulting steady-state DR Kalman gain $K_{\infty}^{*}$.
-The online recursion is identical in structure to the classical steady-state KF, making the approach particularly suitable for long-horizon or real-time applications.
-The complete procedure is given in Algorithm˜2.
 Algorithm 2 Steady-State DRKF
@@
 10:end for
-In the steady state regime, ambiguity in the initial state distribution becomes irrelevant: the Wasserstein radius $\theta_{x_{0}}$ affects only the transient phase and does not influence the asymptotic solution.
-As discussed in Section˜4.2 and Remark˜5, this dependence disappears after finitely many steps.
-Thus, the steady-state DRKF depends only on $(\theta_{w},\theta_{v})$ and the nominal noise covariances.
-3.4 Discussion on Computational Complexity
-The main computational burden in both DRKF variants arises in the offline stage, where SDPs must be solved. The online recursion has the same per-step complexity as the classical KF.
+3.3 Algorithms and Complexity
+We summarize the proposed DRKF methods for both the time-varying and steady-state settings.
+The two variants differ only in how the noise covariances (and hence the Kalman gains) are computed offline to ensure robustness against distributional ambiguity.
+3.3.1 Time-Varying and Steady-State DRKF
+As shown in Algorithm˜1, the time-varying DRKF is implemented by solving the stage-wise SDP problems (4) and (5) offline over a horizon $T$.
+This yields the least-favorable noise covariances and the corresponding DR gains $\{K_{t}\}$.
+If the nominal covariances are known a priori, these gains can be fully precomputed.
+Under the time-invariant assumptions of Section˜3.2, the DRKF recursion may converge to a stationary solution.
+In this case, the offline stage reduces to solving the single stationary SDP (13), which yields the least-favorable noise covariances, the steady-state prior and posterior covariances, and the constant DR Kalman gain $K_{\infty}^{*}$.
+In both cases, the online stage performs only the standard KF measurement update and prediction, and thus has the same per-step computational complexity as the classical KF.
+3.3.2 Discussion on Computational Complexity
+The computational cost of the proposed DRKF methods is dominated by the offline SDP computations; the online recursion has the same per-step complexity as the classical KF.
 […]
 For comparison, the stage-wise DRKF of [26] and the DR-MMSE estimator of [23] also require solving a convex program at each stage.
-The former uses the joint state-measurement covariance of dimension $(n_{x}+n_{y})$, leading to similar polynomial scaling but with larger LMI blocks and constants.
-Like our time-varying DRKF, these methods scale linearly in $T$; however, neither admits a steady-state variant that reduces the offline computation to a single SDP.
-Temporally coupled ambiguity sets, as in [20], can model long-range dependence across the noise sequence but incur substantially higher computational cost: all $T$ stages are merged into a single horizon-level SDP whose number of variables and LMI block sizes both grow linearly with $T$.
-The resulting interior-point complexity, $\tilde{\mathcal{O}}(n_{x}^{6}T^{6})$, quickly becomes prohibitive even for moderate horizons.
+The former relies on joint state–measurement covariances of dimension $(n_{x}+n_{y})$, leading to similar polynomial scaling but with larger LMI blocks and constants.
+Like our time-varying DRKF, these methods scale linearly with $T$, but they do not admit a steady-state formulation that reduces the offline computation to a single SDP.
+Temporally coupled ambiguity sets, as in [13], model long-range dependence across the noise sequence at the cost of significantly higher complexity: all $T$ stages are merged into a single horizon-level SDP whose size grows linearly with $T$, resulting in interior-point complexity of $\tilde{\mathcal{O}}(n_{x}^{6}T^{6})$.
 By contrast, our noise-centric formulation solves a fixed-size SDP at each stage, so the offline cost grows only linearly in $T$.
@@
 No finite-dimensional formulation is available.
-By contrast, our steady-state DRKF is obtained from a single stationary SDP of size $(n_{x}+n_{y})$ with offline complexity $\tilde{\mathcal{O}}\left((n_{x}+n_{y})^{3.5}\right)$, directly yielding a constant gain $K_{\infty}^{*}$ and an online recursion matching the classical steady-state KF.
+In contrast, our steady-state DRKF is obtained from a single stationary SDP of size $(n_{x}+n_{y})$ with offline complexity $\tilde{\mathcal{O}}\left((n_{x}+n_{y})^{3.5}\right)$.
+It directly yields a constant gain $K_{\infty}^{*}$ and an online recursion matching the classical steady-state KF.
 Although this stage-wise ambiguity does not model temporal noise correlations, it preserves the separability of the Riccati recursion and keeps the DR filter tractable for real-time use.
-4 Theoretical Analyses
+4 Theoretical Analysis
 Having established the DRKF formulations in both the finite- and infinite-horizon settings, we now analyze their fundamental theoretical properties.
-Our goal is to show that DRKF retains key structural guarantees of the classical KF, such as stability and boundedness, while providing provable robustness under distributional uncertainty.
+Our goal is to show that the DRKF retains key structural guarantees of the classical KF, such as stability and boundedness, while providing provable robustness under distributional uncertainty.
 In particular, we demonstrate that the DRKF covariance recursion remains spectrally bounded and that its steady-state solution inherits standard convergence properties of the classical filter.
@@
 This analysis shows that the DRKF operates within a uniformly bounded spectral envelope, a property that underpins its stability and distinguishes it from several existing robust formulations.
-Our starting point is a well-known characterization of the Bures–Wasserstein distance.
+Our starting point is a well-known characterization of the (non-squared) Bures–Wasserstein distance.
 Lemma 2.
-[33, Theorem 1] Let $A,B\in\mathbb{S}^{n}_{+}$. The Bures–Wasserstein distance satisfies
+[2, Thm. 1] Let $A,B\in\mathbb{S}^{n}_{+}$. The Bures–Wasserstein distance satisfies
 $\mathcal{B}(A,B)\;=\;\min_{U\in O(n)}\|A^{1/2}-B^{1/2}U\|_{F},$
-where $O(n)$ denotes the orthogonal group. The minimizer exists and is given by the unitary factor in the polar decomposition of $B^{1/2}A^{1/2}$.
+where $O(n)$ denotes the orthogonal group. The minimizer exists and is given by the orthogonal factor in the polar decomposition of $B^{1/2}A^{1/2}$.
 This characterization enables sharp spectral bounds for matrices within a Bures–Wasserstein ball.
@@
 The proof is provided in Appendix B.1.
-This result shows that Bures–Wasserstein balls impose both lower and upper eigenvalue bounds on any feasible covariance matrix.
-Equivalently, the Bures–Wasserstein ball defines an eigenvalue tube around the nominal covariance.
-This geometric interpretation is essential: it guarantees that our DRKF cannot produce arbitrarily inflated or degenerate covariances, even under worst-case perturbations of the noise distributions allowed by the ambiguity sets.
-In contrast, KL-based ambiguity sets can admit pathological least-favorable distributions [29].
+This result shows that a Bures–Wasserstein ball induces both upper and lower spectral bounds on feasible covariance matrices, defining an eigenvalue tube around the nominal covariance.
+This geometric perspective is essential: it rules out pathological behavior in which worst-case covariances can become arbitrarily inflated or collapse in directions that are inconsistent with the nominal statistics.
+As a result, the DRKF operates within a controlled spectral envelope under distributional perturbations allowed by the ambiguity sets.
+By contrast, KL-based ambiguity sets may admit least-favorable distributions with highly distorted or ill-conditioned covariance structure, leading to potentially pathological estimators [5].
 We now translate these spectral bounds to the least-favorable noise covariances that enter the DRKF recursion.
@@
 For ambiguity radii $\theta_{v},\theta_{w},\theta_{x_{0}}$, define
-$\begin{split}\underline{\lambda}_{v,t}&:=\lambda_{\min}(\hat{\Sigma}_{v,t}),\qquad\overline{\lambda}_{v,t}:=\overline{\lambda}(\hat{\Sigma}_{v,t},\theta_{v}),\\ \underline{\lambda}_{w,t}&:=\lambda_{\min}(\hat{\Sigma}_{w,t}),\qquad\overline{\lambda}_{w,t}:=\overline{\lambda}(\hat{\Sigma}_{w,t},\theta_{w})\\ \underline{\lambda}_{x,0}&:=\lambda_{\min}(\hat{\Sigma}_{x,0}^{-}),\qquad\overline{\lambda}_{x,0}:=\overline{\lambda}(\hat{\Sigma}_{x,0}^{-},\theta_{x_{0}}).\end{split}$ (17)
-Then, for all $t>0$, the least-favorable covariances $\Sigma_{v,t}^{*},\Sigma_{w,t}^{*}$ and $\Sigma_{x,0}^{-,*}$ obtained from (4)–(5) satisfy
+$\begin{split}\underline{\lambda}_{v,t}&:=\lambda_{\min}(\hat{\Sigma}_{v,t}),\qquad\overline{\lambda}_{v,t}:=\overline{\lambda}(\hat{\Sigma}_{v,t},\theta_{v}),\\ \underline{\lambda}_{w,t}&:=\lambda_{\min}(\hat{\Sigma}_{w,t}),\qquad\overline{\lambda}_{w,t}:=\overline{\lambda}(\hat{\Sigma}_{w,t},\theta_{w}),\\ \underline{\lambda}_{x,0}&:=\lambda_{\min}(\hat{\Sigma}_{x,0}^{-}),\qquad\overline{\lambda}_{x,0}:=\overline{\lambda}(\hat{\Sigma}_{x,0}^{-},\theta_{x_{0}}).\end{split}$ (17)
+Then, the least-favorable covariances $\Sigma_{v,t}^{*},\Sigma_{w,t}^{*}$ and $\Sigma_{x,0}^{-,*}$ obtained from (4)–(5) satisfy
 $\underline{\lambda}_{v,t}I_{n_{y}}\preceq\Sigma_{v,t}^{*}\preceq\overline{\lambda}_{v,t}I_{n_{y}},$
@@
 The proof is provided in Appendix B.2.
-In the time-invariant setting, the constants $\underline{\lambda}_{v},\overline{\lambda}_{v},\underline{\lambda}_{w},\overline{\lambda}_{w}$ are fixed.
+In the time-invariant setting, the constants $\underline{\lambda}_{w}\coloneq\lambda_{\min}(\hat{\Sigma}_{w})$, $\underline{\lambda}_{v}\coloneq\lambda_{\min}(\hat{\Sigma}_{v})$, $\overline{\lambda}_{w}\coloneq\overline{\lambda}(\hat{\Sigma}_{w},\theta_{w})$, and $\overline{\lambda}_{v}\coloneq\overline{\lambda}(\hat{\Sigma}_{v},\theta_{v})$ are fixed.
 Thus, Corollary˜2 ensures that the DRKF operates within a compact spectral envelope determined solely by the nominal statistics and the Wasserstein radii.
@@
 Consider two classical KFs with noise covariances chosen according to the lower and upper bounds in (17).
-Let $(\underline{\Sigma}_{x,t}^{-},\underline{\Sigma}_{x,t})$ and $(\overline{\Sigma}_{x,t}^{-},\overline{\Sigma}_{x,t})$ denote their respective prior and posterior error covariances, with initial conditions $\underline{\Sigma}_{x,0}^{-}=\underline{\lambda}_{x,0}I$ and $\overline{\Sigma}_{x,0}^{-}=\overline{\lambda}_{x,0}I$.
+Let $(\underline{\Sigma}_{x,t}^{-},\underline{\Sigma}_{x,t})$ and $(\overline{\Sigma}_{x,t}^{-},\overline{\Sigma}_{x,t})$ denote their respective prior and posterior error covariances, with initial conditions $\underline{\Sigma}_{x,0}^{-}=\underline{\lambda}_{x,0}I_{n_{x}}$ and $\overline{\Sigma}_{x,0}^{-}=\overline{\lambda}_{x,0}I_{n_{x}}$.
 Then, for all $t\geq 0$, the DRKF covariances $(\Sigma_{x,t}^{-},\Sigma_{x,t})$ satisfy
@@
 Having established the key theoretical properties of the least-favorable covariance matrices, we now derive conditions under which the DRKF admits a unique steady-state solution, and the sequence of time-varying gains converges to its steady-state value.
-Specifically, recall that the steady-state DR Kalman gain $K_{\infty}$ is defined in (14).
+Specifically, recall that the stationary DR Kalman gain $K_{\infty}^{*}$ is defined in (14).
 Our goal is to show that
-$\lim_{t\to\infty}K_{t}=K_{\infty}.$
-Under the standing assumptions, the steady-state SDP (13) has at least one optimal solution, since its objective is continuous and the feasible set is nonempty and compact due to the Bures-Wasserstein spectral bounds.
-Similarly, the time-varying SDP problems (4) and (5) admit optimal solutions, and therefore the sequence of DR Kalman gains $\{K_{t}\}$ is well defined.
+$\lim_{t\to\infty}K_{t}=K_{\infty}^{*}.$
+Under the standing assumptions, the time-varying SDP problems (4) and (5) admit optimal solutions (the feasible sets are nonempty and compact under the Bures–Wasserstein bounds), and therefore the sequence of DR Kalman gains $\{K_{t}\}$ is well defined.
+In the time-invariant setting, the steady-state SDP (13) is feasible (e.g., by choosing the nominal covariances), and below we show that the DR Riccati recursion converges to a unique fixed point characterized by this SDP.
 The prior covariance recursion (12) can be equivalently expressed as
@@
 with $\Sigma_{x,0}^{-}=\Sigma_{x,0}^{-,*}$.
-Since $\Sigma_{x,0}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}}$ and $\Sigma_{x,t}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{w})I_{n_{x}}$ for $t>0$ at the optimum, $\Sigma_{x,t}^{-}$ is positive definite and the recursion is well defined.
+If $\hat{\Sigma}_{x,0}^{-}\succ 0$, then $\Sigma_{x,0}^{-}\succ 0$ and the recursion is well defined.
+[^8] If $\hat{\Sigma}_{x,0}^{-}\succeq 0$, the uniform bound $\Sigma_{w,t}^{*}\succeq\underline{\lambda}_{w}I$ ensures that $\Sigma_{x,1}^{-}\in\mathbb{S}^{n_{x}}_{++}$, so the information-form recursion is well defined from time $t=1$ onward.
 This motivates the following definition of the DR Riccati mapping:
 $r_{\mathbb{D}}(\Sigma_{x,t}^{-}):=A\bigl((\Sigma_{x,t}^{-})^{-1}+C^{\top}(\Sigma_{v,t}^{*})^{-1}C\bigr)^{-1}A^{\top}+\Sigma_{w,t}^{*}.$ (19)
-As in the classical KF, controllability and observability are needed to guarantee a unique steady state.
+Under Assumption˜2, the problem data of the stage-wise SDP are time-invariant.
+Accordingly, the sequence $\{(\Sigma_{w,t}^{*},\Sigma_{v,t}^{*})\}$ is generated by repeatedly applying the same SDP template along the state covariance trajectory.
+Observing the structural similarity with Riccati equations arising in robust and risk-sensitive filtering (e.g., [32, 10]), we pursue contraction-based analysis in this subsection.
+We begin by imposing the standard observability assumption.
+[^9] Since $\hat{\Sigma}_{w}\succ 0$ and $\Sigma_{w,t}^{*}\succeq\underline{\lambda}_{w}I$, the process noise covariance is uniformly positive definite; hence no additional controllability assumption is required.
 Assumption 3.
-The pair $(A,\hat{\Sigma}_{w}^{1/2})$ is controllable and the pair $(A,C)$ is observable.
-To isolate the effect of distributional robustness in the process channel, we parameterize the perturbation through $\Delta_{t}$.
-Assumption 4.
-There exists $\Delta_{t}\in\mathbb{S}^{n_{x}}_{+}$ such that $A\Delta_{t}A^{\top}=\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}$.
-Under Assumption˜4, the recursion (18) can be rewritten as
-$r_{\mathbb{D}}(\Sigma_{x,t}^{-})=A\bigl((\Sigma_{x,t}^{-})^{-1}+C^{\top}\hat{\Sigma}_{v}^{-1}C-\Phi_{t}\bigr)^{-1}A^{\top}+\hat{\Sigma}_{w},$ (20)
+The pair $(A,C)$ is observable.
+To establish strict contraction under the observability assumption, we employ a downsampled (lifted) representation of the DR Riccati recursion, since the one-step mapping may fail to be contractive due to rank-deficient measurement operators, whereas observability guarantees full column rank of the lifted observability matrix.
+Fix $N\in\mathbb{N}$ and consider the system (1) under the least-favorable distributions.
+Define the downsampled state as $x_{t}^{d}\coloneq x_{tN}$.
+Downsampling yields an NN-block lifted model with block-diagonal uncertainty structure, enabling a contraction argument similar to [34, 36].
+Let $\bm{w}_{t}^{N}\coloneqq[w_{tN+N-1}^{\top},\dots,w_{tN}^{\top}]^{\top}$, and $\bm{v}_{t}^{N}\coloneqq[v_{tN+N-1}^{\top},\dots,v_{tN}^{\top}]^{\top}$ denote the stacked process and measurement noise vectors, with means $\mathbf{1}_{N}\otimes\hat{w}$ and $\mathbf{1}_{N}\otimes\hat{v}$, and covariances $\Sigma_{w,t}^{N,*}\coloneq\operatorname{blkdiag}(\Sigma_{w,tN+N-1}^{*},\dots,\Sigma_{w,tN}^{*})$ and $\Sigma_{v,t}^{N,*}\coloneq\operatorname{blkdiag}(\Sigma_{v,tN+N-1}^{*},\dots,\Sigma_{v,tN}^{*})$, respectively.
+The downsampled dynamics then take the form
+$\begin{split}x_{t+1}^{d}&=A^{N}x_{t}^{d}+\mathcal{R}_{N}\bm{w}_{t}^{N}\\ \bm{y}_{t}^{N}&=\mathcal{O}_{N}x_{t}^{d}+\mathcal{H}_{N}\bm{w}_{t}^{N}+\bm{v}_{t}^{N},\end{split}$
+where $\bm{y}_{t}^{N}\coloneqq[y_{tN+N-1}^{\top},\dots,y_{tN}^{\top}]^{\top}$ is the stacked observation vector, while the block reachability and observability matrices are defined as
+$\mathcal{R}_{N}$ $\coloneqq\left[I_{n_{x}},\,A,\,\ldots,\,A^{N-1}\right]$ (20)
+$\mathcal{O}_{N}$ $\coloneqq\left[(CA^{N-1})^{\top},\,\ldots,\,(CA)^{\top},\,C^{\top}\right]^{\top}.$ (21)
+The block Toeplitz matrix $\mathcal{H}_{N}\in\real{Nn_{y}\times Nn_{x}}$ captures the propagation of process noise into future outputs and is defined elementwise by $\mathcal{H}_{N}]_{ij}=CA^{j-i-1},j>i$, and zero otherwise, for $i,j=1,\ldots,N$.
+Let $\bm{u}_{t}^{N}\coloneq\mathcal{H}_{N}\bm{w}_{t}^{N}+\bm{v}_{t}^{N}$.
+Its covariance is given by $\Sigma_{u,t}^{N}=\mathcal{H}_{N}\Sigma_{w,t}^{N,*}\mathcal{H}_{N}^{\top}+\Sigma_{v,t}^{N,*}$.
+Adopting the procedure in [16] and denoting the downsampled prior covariance matrix by $\Sigma_{x,t}^{-,d}:=\Sigma_{x,tN}^{-}$, the DR Riccati recursion associated with the downsampled model becomes
+$\Sigma_{x,t+1}^{-,d}=r_{\mathbb{D},t}^{d}(\Sigma_{x,t}^{-,d}),$ (22)
 where
-$\Phi_{t}\coloneqq(\Sigma_{x,t}^{-})^{-1}+C^{\top}\hat{\Sigma}_{v}^{-1}C-\left(I+X_{t}\Delta_{t}\right)^{-1}X_{t},$ (21)
-and
-$X_{t}:=(\Sigma_{x,t}^{-})^{-1}+C^{\top}(\Sigma_{v,t}^{*})^{-1}C.$
-Assumption˜4 is a structural compatibility condition ensuring that the perturbation in the process-noise covariance lies in the range of the dynamics matrix $A$.
-While the DR Riccati iteration often converges numerically even when this condition is mildly violated, our formal convergence and uniqueness guarantees rely on it.
+$r_{\mathbb{D},t}^{d}(\Sigma_{x,t}^{-,d})=\alpha_{N,t}\Bigl((\Sigma_{x,t}^{-,d})^{-1}+\Omega_{N,t}\Bigr)^{-1}(\alpha_{N,t})^{\top}+W_{N,t},$ (23)
+with
+$\alpha_{N,t}$ $\coloneq A^{N}-\mathcal{R}_{N}\mathcal{G}_{N,t}\mathcal{O}_{N}$
+$\Omega_{N,t}$ $\coloneq\mathcal{O}_{N}^{\top}(\Sigma_{u,t}^{N})^{-1}\mathcal{O}_{N}$
+$W_{N,t}$ $\coloneq\mathcal{R}_{N}\mathcal{Q}_{N,t}\mathcal{R}_{N}^{\top}$
+$\mathcal{Q}_{N,t}$ $\coloneq\Sigma_{w,t}^{N,*}-\Sigma_{w,t}^{N,*}\mathcal{H}_{N}^{\top}(\Sigma_{u,t}^{N})^{-1}\mathcal{H}_{N}\Sigma_{w,t}^{N,*}$
+$\mathcal{G}_{N,t}$ $\coloneq\Sigma_{w,t}^{N,*}\mathcal{H}_{N}^{\top}(\Sigma_{u,t}^{N})^{-1}.$
+Under Assumption˜2, the one-step DR Riccati update is time-invariant, and the corresponding downsampled mapping $r_{\mathbb{D},t}^{d}$ coincides with the NN-fold iterate $r_{\mathbb{D}}^{N}$ of a single Riccati operator $r_{\mathbb{D}}$.
+We therefore analyze contraction properties of the DR Riccati recursion through this lifted representation, via a downsampling contraction argument in the spirit of [36].
+We first establish basic structural properties of $\Omega_{N,t}$ and $W_{N,t}$, which are required for the contraction analysis.
+Lemma 3.
+Suppose Assumptions˜3, 1 and 2 hold. Then, $W_{N,t}\in\mathbb{S}^{n_{x}}_{++}$ for any $N\geq 1$ and $\Omega_{N,t}\in\mathbb{S}^{n_{x}}_{++}$ for any $N\geq n_{x}$.
+The proof is provided in Appendix C.1.
+Using the above properties, we establish contraction of the (downsampled) DR Riccati mapping with respect to Thompson’s part metric, defined as follows: for $P,Q\in\mathbb{S}^{n_{x}}_{++}$,
+$d_{T}(P,Q)\coloneq\log\left(\max\{\lambda_{\max}(P^{-1}Q),\lambda_{\max}(Q^{-1}P)\}\right).$
+This metric is invariant under congruence transformations and is well suited for analyzing Riccati-type mappings.
+Lemma 4.
+Suppose Assumptions˜3, 1 and 2 hold and let $N\geq n_{x}$, so that $\mathcal{O}_{N}$ has full column rank.
+Then, for any $t\geq 0$, the downsampled DR Riccati mapping $r_{\mathbb{D},t}^{d}$ (23) is strictly contractive on $\mathbb{S}^{n_{x}}_{++}$ with respect to $d_{T}$: there exists $\xi_{N,t}\in(0,1)$ such that
+$d_{T}(r_{\mathbb{D},t}^{d}(P),r_{\mathbb{D},t}^{d}(Q))\leq\xi_{N,t}\,d_{T}(P,Q),\quad\forall P,Q\in\mathbb{S}^{n_{x}}_{++}.$
+Moreover, the contraction factor satisfies the bound
+$\xi_{N,t}\leq\left(\frac{\sqrt{\|M_{N,t}\|_{2}}}{1+\sqrt{1+\|M_{N,t}\|_{2}}}\right)^{2},$ (24)
+where $M_{N,t}\coloneq\Omega_{N,t}^{-1/2}\alpha_{N,t}^{\top}W_{N,t}^{-1}\alpha_{N,t}\Omega_{N,t}^{-1/2}$.
+This follows from [14, Thm. 5.3] together with Lemma˜3.
+We next show that the contraction factor can be bounded uniformly over time by exploiting the spectral envelope induced by Wasserstein ambiguity.
+Proposition 2.
+Suppose that Assumptions˜1, 2 and 3 hold. Fix $N\geq n_{x}$ and define
+$\begin{split}\kappa_{N}(\theta_{v},\theta_{w})\coloneq\;&\frac{\overline{\lambda}(\hat{\Sigma}_{v},\theta_{v})+\overline{\lambda}(\hat{\Sigma}_{w},\theta_{w})\|\mathcal{H}_{N}\|_{2}^{2}}{\sigma_{\min}(\mathcal{O}_{N})^{2}}\\ &\times\left(\|A^{N}\|_{2}+\frac{\overline{\lambda}(\hat{\Sigma}_{w},\theta_{w})\|\mathcal{H}_{N}\|_{2}\|\mathcal{R}_{N}\|_{2}\|\mathcal{O}_{N}\|_{2}}{\lambda_{\min}(\hat{\Sigma}_{v})}\right)^{2}\left(\frac{1}{\lambda_{\min}(\hat{\Sigma}_{w})}+\frac{\|\mathcal{H}_{N}\|_{2}^{2}}{\lambda_{\min}(\hat{\Sigma}_{v})}\right),\end{split}$ (25)
+where $\overline{\lambda}(\cdot,\cdot)$ is defined as in Proposition˜1. Then, for all $t\geq 0$, the contraction factor in (24) satisfies
+$\xi_{N,t}\leq\bar{\xi}_{N}(\theta_{v},\theta_{w})\coloneq\left(\frac{\sqrt{\kappa_{N}(\theta_{v},\theta_{w})}}{1+\sqrt{1+\kappa_{N}(\theta_{v},\theta_{w})}}\right)^{2}\in(0,1).$
+In particular, for any prescribed $\rho\in(0,1)$, any radii $(\theta_{v},\theta_{w})$ satisfying
+$\kappa_{N}(\theta_{v},\theta_{w})\leq\frac{4\rho}{(1-\rho)^{2}}$ (26)
+ensure $\xi_{N,t}\leq\rho$ for all $t$.
+The proof is provided in Appendix C.2.
+Combining the pointwise contraction in Lemma˜4 with the uniform bound in Proposition˜2 yields convergence of the DR Riccati recursion.
+Theorem 3.
+Suppose Assumptions˜3, 1 and 2 hold and let $N\geq n_{x}$.
+Then, for any $t\geq 0$, the downsampled DR Riccati mapping $r_{\mathbb{D},t}^{d}$ is strictly contractive on $\mathbb{S}^{n_{x}}_{++}$ with respect to Thompson’s part metric, with contraction factor $\xi_{N,t}$ satisfying
+$\xi_{N,t}\leq\bar{\xi}_{N}(\theta_{v},\theta_{w})<1\qquad\forall t\geq 0.$
+Consequently, the full DR Riccati recursion (18) converges to a unique fixed point $\Sigma_{x,\infty}^{-}\succ 0$, and the associated DR Kalman gain $K_{t}$ converges.
+The proof is provided in Appendix C.3.
+As a consequence of Theorem˜3, the steady-state solution can be obtained by solving the single stationary SDP (13) offline, rather than via online Riccati iterations.
+Corollary 3.
+Under the assumptions of Theorem˜3, the steady-state covariance matrices $(\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*})$, together with the least-favorable noise covariances $\Sigma_{w,\infty}^{*}$ and $\Sigma_{v,\infty}^{*}$, solve the stationary SDP (13).
+Consequently, $K_{\infty}^{*}=\lim_{t\to\infty}K_{t}$.
+The proof can be found in Appendix C.4.
+Under Assumption˜3, the Riccati recursions of the standard KFs in Theorem˜2 converge.
+Taking limits in the matrix inequalities $A\preceq B$ (in the sense of the Loewner partial order) appearing in Theorem˜2, and using that the positive semidefinite cone is closed, yields the following steady-state bounds.
+Corollary 4.
+Under the assumptions of Theorem˜3, the steady-state prior and posterior error covariances $\Sigma_{x,\infty}^{-}$ and $\Sigma_{x,\infty}$ satisfy
+$\underline{\Sigma}_{x,\infty}^{-}\preceq\Sigma_{x,\infty}^{-}\preceq\overline{\Sigma}_{x,\infty}^{-},\qquad\underline{\Sigma}_{x,\infty}\preceq\Sigma_{x,\infty}\preceq\overline{\Sigma}_{x,\infty},$
+where $\underline{\Sigma}_{x,\infty}^{-}\coloneq\lim_{t\to\infty}\underline{\Sigma}_{x,t}^{-}$ and $\overline{\Sigma}_{x,\infty}^{-}\coloneq\lim_{t\to\infty}\overline{\Sigma}_{x,t}^{-}$ (and similarly for $\underline{\Sigma}_{x,\infty},\overline{\Sigma}_{x,\infty}$) are the unique steady-state limits of the lower- and upper-bound KF recursions in Theorem˜2.
 Remark 4.
-Assumption˜4 holds whenever $\mathrm{range}(\Sigma_{w,t}^{*}-\hat{\Sigma}_{w})\subseteq\mathrm{range}(A)$. In this case, a canonical choice of $\Delta_{t}$ is the minimum-Frobenius-norm solution
-$\Delta_{t}:=A^{+}(\Sigma_{w,t}^{*}-\hat{\Sigma}_{w})(A^{+})^{\top}.$ (22)
-This matrix $\Delta_{t}$ is positive semidefinite, satisfies $A\Delta_{t}A^{\top}=\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}$, and obeys the bound
-$\|\Delta_{t}\|_{2}\leq\|A^{+}\|_{2}^{2}\|\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}\|_{2}.$
-In practice, verifying $\mathrm{range}(\Sigma_{w,t}^{*}-\hat{\Sigma}_{w})\subseteq\mathrm{range}(A)$ requires knowledge of the time-varying covariance $\Sigma_{w,t}^{*}$.
-A convenient sufficient condition that can be checked a priori uses the spectral bound $\overline{\lambda}_{w}$ from Corollary˜2: if $\mathrm{range}(\overline{\lambda}_{w}I-\hat{\Sigma}_{w})\subseteq\mathrm{range}(A)$, then Assumption˜4 holds for all $t$, since
-$0\preceq\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}\preceq\overline{\lambda}_{w}I-\hat{\Sigma}_{w}\implies\mathrm{range}(\Sigma_{w,t}^{*}-\hat{\Sigma}_{w})\subseteq\mathrm{range}(\overline{\lambda}_{w}I-\hat{\Sigma}_{w})\subseteq\mathrm{range}(A).$
-The form in (20) is structurally analogous to the risk-sensitive Riccati equation [4, 34].
-Levy et al.
-[35] showed that, under Assumption˜3, the $N$-fold composition of the risk-sensitive Riccati mapping is a contraction in a suitable Riemannian metric for $N\geq n_{x}$, while Zorzi [36] extended this result to robust filters under the Thompson part metric.
-These results motivate the use of contraction-based arguments for analyzing the DR Riccati mapping.
-An alternative interpretation of (20) arises by embedding the DR filter in a Krein space [37].
-Augmenting the nominal system with a synthetic uncertainty $u_{t}$ and enforcing the constraint $0=\hat{x}_{t}+u_{t}$ yields an indefinite inner product on $u_{t}$ weighted by $\Phi_{t}^{-1}$.
-This perspective links distributional robustness to risk-sensitive filtering and provides a natural foundation for contraction-based analyses.
-To formalize this idea, fix $N\in\mathbb{N}$ and consider the downsampled state $\hat{x}_{t}^{d}=\hat{x}_{tN}$.
-Downsampling lifts the time-varying recursion to an $N$-block time-invariant map, yielding block-diagonal perturbations and enabling a contraction argument similar to [37].
-This induces an equivalent block model with matrices $(\mathcal{R}_{N},\mathcal{O}_{N},\mathcal{L}_{N},\mathcal{H}_{N})$ and a block-diagonal perturbation
-$\bar{\Phi}_{N,t}=\mathrm{blkdiag}(\Phi_{tN+N-2},\ldots,\Phi_{tN-1}).$
-The details are provided in Appendix C.1. Letting $\Sigma_{x,t}^{-,d}\coloneq\Sigma_{x,tN}^{-}$ denote the downsampled prior covariance matrix, the resulting Riccati recursion becomes
-$\Sigma_{x,t+1}^{-,d}=r_{\mathbb{D},t}^{d}(\Sigma_{x,t}^{-,d}),$ (23)
-where $r_{\mathbb{D},t}^{d}$ denotes the $N$-fold composition of $r_{\mathbb{D}}$.
-If $r_{\mathbb{D},t}^{d}$ is a contraction for sufficiently large $t$, then so is $r_{\mathbb{D}}$.
-By Banach’s fixed-point theorem [38], this ensures the existence and uniqueness of a fixed point $\Sigma_{x,\infty}^{-}\succ 0$, and hence of a steady-state gain $K_{\infty}$.
-The following proposition provides a tractable sufficient condition for ensuring contraction.
-Proposition 2.
-Suppose Assumptions˜1, 2, 3 and 4 hold. Define
-$\tilde{\phi}_{N}\;=\;\frac{1}{\lambda_{\max}\left(\mathcal{L}_{N}\bigl(I_{Nn_{x}}+\mathcal{H}_{N}^{\top}(\mathcal{D}_{N}\mathcal{D}_{N}^{\top})^{-1}\mathcal{H}_{N}\bigr)^{-1}\mathcal{L}_{N}^{\top}\right)}.$
-Then, for some $N\geq n_{x}$, there exists $\phi_{N}\in(0,\tilde{\phi}_{N})$ such that if
-$0\preceq\bar{\Phi}_{N,t}\preceq\phi_{N}I_{Nn_{x}},\qquad\forall t\geq q,$ (24)
-for some $q>0$, the mapping $r_{\mathbb{D},t}^{d}$ is a contraction with respect to the Thompson part metric [39] on the cone of positive definite matrices for all $t\geq q$, and hence $r_{\mathbb{D}}$ is also a contraction in the same metric.
-The proof is provided in Appendix C.2.
-A practical procedure for determining a suitable $\phi_{N}$ is described in [36].
-One begins with $\phi_{N}=\tilde{\phi}_{N}$ and checks whether the associated Gramian $\Omega_{\phi_{N}I_{Nn_{x}}}$ is positive definite.
-If not, $\phi_{N}$ is iteratively reduced until $\Omega_{\phi_{N}I_{Nn_{x}}}$ becomes positive semidefinite but singular.
-At this threshold, if $\bar{\Phi}_{N,t}\prec\phi_{N}I_{Nn_{x}}$ for all $t\geq\tilde{q}$, then both Gramians $\Omega_{\bar{\Phi}_{N,t}}$ and $W_{\bar{\Phi}_{N,t}}$ are positive definite for $t\geq\tilde{q}$, which guarantees that the downsampled DR Riccati mapping (23) is a contraction.
-Proposition˜2 establishes convergence provided that $\bar{\Phi}_{N,t}$ remains uniformly bounded by $\phi_{N}I_{Nn_{x}}$.
-In practice, however, $\bar{\Phi}_{N,t}$ depends on the sequence of SDP solutions in (7) and is therefore difficult to verify a priori.
-This motivates the derivation of explicit, tractable sufficient conditions on the ambiguity radii $(\theta_{v},\theta_{w})$ that guarantee the bound $\bar{\Phi}_{N,t}\preceq\phi_{N}I_{Nn_{x}}$ uniformly across time.
-Here, our noise-centric modeling becomes essential, as it leads to closed-form, computationally tractable expressions for these radii.
+Since $\overline{\lambda}(\hat{X},\theta)$ is nondecreasing in $\theta$ for $\theta\geq 0$, the bound $\kappa_{N}(\theta_{v},\theta_{w})$ in (25) is monotone in both ambiguity radii.
+Consequently, a practical tuning procedure for $(\theta_{v},\theta_{w})$ consists of enforcing the rate constraint (26) using one of the following strategies: (i) fix $\theta_{v}$ and determine the largest admissible $\theta_{w}$; (ii) fix $\theta_{w}$ and determine the largest admissible $\theta_{v}$; or (iii) restrict $(\theta_{v},\theta_{w})$ to a one-dimensional curve (e.g., $\theta_{w}=\gamma\theta_{v}$ for some $\gamma>0$) and select the maximal radii satisfying (26).
+These procedures allow one to balance robustness and convergence speed in a transparent and computationally tractable manner.
+4.3 Stability
+A key property of the classical Kalman filter is the stability of its steady-state estimation error dynamics: under detectability of $(A,C)$ and positive definite process noise, the closed-loop error matrix (IKC)A(I-KC)A is Schur stable, ensuring bounded error covariance and asymptotic convergence of the estimation error in the unbiased case.
+In the distributionally robust setting, however, the estimator is designed against least-favorable noise distributions within ambiguity sets.
+It is thus nontrivial whether the resulting DRKF, whose steady-state gain is characterized by the minimax design in Theorem˜3, preserves the stability guarantees of the classical KF.
+The following result shows that, despite the adversarial nature of the noise model, the steady-state DRKF remains asymptotically stable in the same sense as the classical KF.
 Theorem 4.
-Suppose Assumptions˜3, 1, 2 and 4 hold. Let $\phi_{v},\phi_{x}\geq 0$ with $\phi_{v}+\phi_{x}\leq\phi_{N}$, where $\phi_{N}$ is defined in Proposition˜2. Then, the following properties hold.
-  1. (ii)
-    (Measurement uncertainty bound) For any $\phi_{v}\geq 0$, define
-    $\delta_{v}^{\max}(\phi_{v})\coloneqq\frac{\phi_{v}(\lambda_{\min}(\hat{\Sigma}_{v}))^{2}}{\|C\|_{2}^{2}+\phi_{v}\lambda_{\min}(\hat{\Sigma}_{v})}<\lambda_{\min}(\hat{\Sigma}_{v}).$ (25)
-    Then, any $\theta_{v}$ satisfying
-    $0\leq\theta_{v}\leq\sqrt{\lambda_{\max}(\hat{\Sigma}_{v})+\delta_{v}^{\max}(\phi_{v})}-\sqrt{\lambda_{\max}(\hat{\Sigma}_{v})}$ (26)
-    guarantees that, for all $t\geq q$,
-    $C^{\top}\!\bigl(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\bigr)C\;\preceq\;\phi_{v}I.$ (27)
-  2. (iiii)
-    (Process uncertainty bound) Define
-    $\delta_{x}^{\max}\;\coloneqq\;\frac{1}{\lambda_{\min}(\hat{\Sigma}_{w})}+\frac{\|C\|_{2}^{2}}{\lambda_{\min}(\hat{\Sigma}_{v})}.$ (28)
-    If $0\leq\phi_{x}<\delta_{x}^{\max}$, then any $\theta_{w}$ satisfying
-    $0\leq\theta_{w}\leq\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})+\frac{\phi_{x}}{\|A^{+}\|_{2}^{2}\,\delta_{x}^{\max}(\delta_{x}^{\max}-\phi_{x})}}-\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}$ (29)
-    guarantees that, for all $t\geq q$,
-    $X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}\preceq\phi_{x}I.$ (30)
-Moreover, any pair $(\theta_{v},\theta_{w})$ satisfying (26) and (29) ensures that the DR Riccati mapping (19) admits a unique fixed point $\Sigma_{x,\infty}^{-}\succ 0$, and the corresponding DR Kalman gain converges: $K_{t}\to K_{\infty}$.
-The proof of this theorem is provided in Appendix C.4.
-The bounds (26) and (29) yield tractable, closed-form radii $(\theta_{v},\theta_{w})$ that certify convergence, directly in terms of nominal covariances and operator norms.
-This explicit tuning rule represents a practical consequence of our noise-centric ambiguity design.
-Remark 5.
-The radius $\theta_{x_{0}}$ is arbitrary and affects only the transient phase.
-The sufficient conditions (26) and (29) depend only on $(\theta_{v},\theta_{w})$ and the nominal noise covariances.
-In particular, for any $\Sigma_{x,0}^{-}\in\mathbb{S}^{n_{x}}_{++}$, the first prediction satisfies $\Sigma_{x,1}^{-}\succeq\Sigma_{w,0}^{\star}\succeq\hat{\Sigma}_{w}$.
-Thereafter, the contraction bounds are independent of the prior.
-Hence, if  (26) and (29) hold uniformly, one may take $q=0$ in Theorem˜3; otherwise, convergence is still guaranteed after a finite transient $q\geq 1$.
-An immediate implication of Theorem˜3 is that the steady-state solution can be computed directly through the single stationary SDP (13) offline, rather than via online Riccati iterations.
-Corollary 3.
-Suppose that Assumptions˜1, 2, 3 and 4 hold.
-Then, the steady-state covariance matrices $(\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*})$, together with the least-favorable noise covariances $\Sigma_{w,\infty}^{*}$ and $\Sigma_{v,\infty}^{*}$, solve the stationary SDP (13).
-Consequently, $K_{\infty}^{*}=K_{\infty}=\lim_{t\to\infty}K_{t}$.
-Under Assumption˜3, the Riccati recursions of the standard KFs in Theorem˜2 converge.
-When the DRKF covariances also converge, we can take limits in the matrix inequality of Theorem˜2 and, using that limits preserve positive semidefiniteness, obtain steady-state upper and lower bounds on the DRKF covariance, stated in the following corollary.
-Corollary 4.
-Suppose that Assumptions˜1, 2, 3 and 4 hold and the DRKF converges.
-Then, the steady-state prior and posterior error covariances $\Sigma_{x,\infty}^{-}$ and $\Sigma_{x,\infty}$ are bounded by the limits of the lower- and upper-bound Kalman recursions described in Theorem˜2:
-$\underline{\Sigma}_{x,\infty}^{-}$ $\preceq\Sigma_{x,\infty}^{-}\preceq\overline{\Sigma}_{x,\infty}^{-}$
-$\underline{\Sigma}_{x,\infty}$ $\preceq\Sigma_{x,\infty}\preceq\overline{\Sigma}_{x,\infty},$
-where $\underline{\Sigma}_{x,\infty}^{-}\coloneqq\lim_{t\to\infty}\underline{\Sigma}_{x,t}^{-}$ and $\overline{\Sigma}_{x,\infty}^{-}\coloneqq\lim_{t\to\infty}\overline{\Sigma}_{x,t}^{-}$ denote the steady-state bounds for the prior covariances, and the posterior bounds $\underline{\Sigma}_{x,\infty}$ and $\overline{\Sigma}_{x,\infty}$ are defined analogously.
-Finally, the DR Riccati recursion (18) reveals a structural connection to existing robust filtering paradigms.
-In particular, it resembles the risk-sensitive Riccati equation [4, 34].
-This correspondence implies that the time-varying DRKF can be interpreted as a risk-sensitive filter with an effective risk parameter $\theta$ and a time-varying weighting matrix $Q_{t}\in\mathbb{S}^{n_{x}}_{+}$ satisfying $\theta Q_{t}^{\top}Q_{t}=\Phi_{t-1}$.
-At steady state, DRKF reduces to a risk-sensitive filter with a constant weighting matrix $Q$ chosen such that $\theta Q^{\top}Q=\Phi_{\infty}$, where $\Phi_{\infty}$ denotes the limiting perturbation.
-4.3 Stability
-A fundamental property of the classical KF is the asymptotic stability of its estimation error dynamics under standard controllability and observability assumptions.
-Specifically, the closed-loop error matrix (IKC)A(I-KC)A is Schur, ensuring boundedness of the estimation error covariance and decay of the error mean in the absence of bias.
-In the DR setting, however, the estimator competes against the least-favorable noise distributions within the ambiguity set.
-A natural question is therefore whether the DRKF preserves this stability property.
-The following result shows that, despite the adversarial nature of the noise, the DRKF error dynamics remain asymptotically stable in the same sense as in the classical case.
-Theorem 4.
-Suppose Assumptions˜3, 1, 2 and 4 hold, and the ambiguity radii $(\theta_{v},\theta_{w})$ satisfy the conditions of Theorem˜3.
-Let $(\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*})$ and $(\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*})$ denote the steady-state DRKF covariances, and let $K_{\infty}$ be the corresponding steady-state DR gain.
-Define the closed-loop error matrix $F_{\infty}\coloneq(I-K_{\infty}C)A$.
+Suppose Assumptions˜3, 1 and 2 hold.
+Let $(\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*})$ and $(\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*})$ denote the steady-state DRKF covariances, and let $K_{\infty}^{*}$ be the corresponding steady-state DR Kalman gain.
+Define the closed-loop error matrix $F_{\infty}\coloneq(I-K_{\infty}^{*}C)A$.
 Then, $F_{\infty}$ is Schur stable, and the posterior estimation error $e_{t}:=x_{t}-\bar{x}_{t}$ evolves according to
 […]
-Consequently, under the least-favorable noise distributions,
-$\lim_{t\to\infty}\mathbb{E}\|e_{t}\|^{2}\;=\;\operatorname{Tr}[\Sigma_{x,\infty}^{*}]\quad\text{and}\quad\sup_{t\geq 0}\,\mathbb{E}\|e_{t}\|^{2}\;<\;\infty.$
+Consequently, under the least-favorable noise distributions, the estimation error is mean-square stable with
+$\sup_{t\geq 0}\,\mathbb{E}[\|e_{t}\|^{2}]\;<\;\infty,\quad\text{and}\quad\lim_{t\to\infty}\mathrm{cov}(e_{t})=\Sigma_{x,\infty}^{*}.$
 The proof is provided in Appendix D.1.
-While Theorem˜4 guarantees mean-square stability, it does not by itself characterize the estimation-error mean in the presence of bias.
-When the least-favorable distributions (or the true noise processes) have nonzero means, the filter tracks a steady bias determined by the mean mismatch.
-The following corollary characterizes this bias.
+While Theorem˜4 establishes mean-square stability, it does not by itself characterize the behavior of the estimation-error mean under model mismatch.
+In practice, the true noise processes may have means that differ from the nominal values used by the filter, leading to a steady-state bias.
+The following corollary characterizes this bias under the true noise distributions.
 Corollary 5.
-Under the assumptions of Theorem˜4, let $m_{t}:=\mathbb{E}[e_{t}]=\mathbb{E}[x_{t}-\bar{x}_{t}]$ denote the posterior error mean under the least-favorable distributions.
-Let $\mathbb{E}[w_{t}]=\mu_{w}$ and $\mathbb{E}[v_{t}]=\mu_{v}$.
-Then,
 […]
-and hence
 […]
-In particular, if the noise means used by the filter coincide with the true means ($\hat{w}=\mu_{w}$ and $\hat{v}=\mu_{v}$), then $\mathbb{E}[e_{t}]\to 0$.
-The proof of this corollary is provided in Appendix D.2.
-Corollary˜5 shows that our DRKF mirrors the classical KF in its treatment of bias: if the true noise means coincide with those used by the filter ($\mu_{w}=\hat{w}$ and $\mu_{v}=\hat{v}$), then the estimation-error mean vanishes asymptotically.
-If instead either mean is misspecified ($\mu_{w}\neq\hat{w}$ or $\mu_{v}\neq\hat{v}$), the filter remains stable and the bias converges to the constant vector in the right-hand side above, which depends only on $F_{\infty}$, $K_{\infty}$, and the mean mismatch.
+In particular, if the nominal noise means used by the filter coincide with the true noise means ($\hat{w}=\mu_{w}$ and $\hat{v}=\mu_{v}$), then $\mathbb{E}[e_{t}]\to 0$.
+The proof is provided in Appendix D.2.
+As shown in Lemma˜1, mean shifts are suboptimal for the adversary under Wasserstein ambiguity at the design stage, so the least-favorable noise distributions preserve the nominal means.
+Therefore, when the nominal noise means used by the filter coincide with the true noise means, the steady-state DRKF is unbiased in operation.
+When this condition is violated, a constant steady-state bias appears, exactly as characterized in Corollary˜5.
 4.4 Optimality
-A key property of the classical KF is that it is the unique linear, causal estimator minimizing the MSE when the Gaussian noise distributions are known exactly.
-In the DR setting, however, the true noise covariances are unknown and only assumed to lie within Wasserstein ambiguity sets.
-In this case, it is natural to consider minimax optimality: among all causal estimators, the filter should minimize the worst-case one-step MSE over all distributions in the ambiguity sets.
-The following theorem establishes that DRKF satisfies this minimax optimality in an asymptotic sense.
+Having established convergence and stability of the proposed DRKF, we now address its performance optimality.
+In the presence of distributional uncertainty, a natural question is whether one can design a causal estimator that outperforms the DRKF in terms of worst-case estimation accuracy.
+In this subsection, we show that this is not possible asymptotically: among all causal estimators, the steady-state DRKF minimizes the worst-case one-step conditional mean-square error in the limit, as well as its long-run time average.
 Theorem 5.
-Suppose Assumptions˜1, 2, 3 and 4 hold.
-Let $(\theta_{v},\theta_{w})$ satisfy the conditions of Theorem˜3, and let $\Sigma_{x,\infty}^{-,*}\in\mathbb{S}^{n_{x}}_{++}$ denote the unique fixed point of the DR Riccati map $r_{\mathbb{D}}$.
-Let $\Sigma_{x,\infty}^{*}$ be the corresponding steady-state posterior covariance.
-Then, for any causal state estimator sequence $\{\psi_{t}\}_{t\geq 0}$ with $\psi_{t}\in\mathcal{F}_{t}$, we have
+Suppose Assumptions˜1, 2 and 3 hold.
+Let $\Sigma_{x,\infty}^{-,*}\in\mathbb{S}^{n_{x}}_{++}$ denote the unique fixed point of the DR Riccati map $r_{\mathbb{D}}$ and let $\Sigma_{x,\infty}^{*}$ be the corresponding steady-state posterior error covariance.
+Then, for any causal state estimator sequence $\{\psi_{t}\}_{t\geq 0}$ with $\psi_{t}\in\mathcal{F}_{t}$, the following inequalities hold almost surely with respect to $\sigma(\mathcal{Y}_{t-1})$:
 $\liminf_{t\to\infty}\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})$ $\geq\operatorname{Tr}[\Sigma_{x,\infty}^{*}],$ (27)
-$\liminf_{T\to\infty}\ \frac{1}{T}\sum_{t=1}^{T}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})$ $\geq\operatorname{Tr}[\Sigma_{x,\infty}^{*}],$ (32)
-almost surely. Moreover, the steady-state DRKF $\psi_{\infty}$ with gain $K_{\infty}$ achieves these bounds with equality:
-$\begin{split}\lim_{t\to\infty}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})&=\operatorname{Tr}[\Sigma_{x,\infty}^{*}],\\ \lim_{T\to\infty}\ \frac{1}{T}\sum_{t=1}^{T}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})&=\operatorname{Tr}[\Sigma_{x,\infty}^{*}],\end{split}$
-almost surely. Hence, the steady-state DRKF is asymptotically minimax-optimal with respect to the worst-case one-step and long-run average MSE.
-The proof of this theorem is provided in Appendix D.3.
+$\liminf_{T\to\infty}\ \frac{1}{T}\sum_{t=0}^{T-1}\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})$ $\geq\operatorname{Tr}[\Sigma_{x,\infty}^{*}].$ (28)
+Moreover, the steady-state DRKF $\psi_{\infty}$ with gain $K_{\infty}^{*}$ achieves these bounds with equality:
+$\lim_{t\to\infty}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})$ $=\operatorname{Tr}[\Sigma_{x,\infty}^{*}],$ (29)
+$\lim_{T\to\infty}\ \frac{1}{T}\sum_{t=0}^{T-1}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})$ $=\operatorname{Tr}[\Sigma_{x,\infty}^{*}].$ (30)
+Hence, the steady-state DRKF is asymptotically minimax-optimal with respect to the worst-case one-step and long-run average MSE.
+The proof of is provided in Appendix D.3.
 This result shows that the steady-state DRKF attains the smallest possible worst-case asymptotic MSE among all causal estimators, thereby extending the fundamental optimality of the classical KF to the DR setting.
-Importantly, the result concerns asymptotic performance rather than global finite-horizon minimax optimality, which is generally not well-posed under stage-wise ambiguity.
+Importantly, the theorem concerns asymptotic performance: establishing finite-horizon minimax optimality under stage-wise ambiguity leads to a substantially more intricate dynamic game and is beyond the scope of this work.
 Remark 5.
-When $\theta_{v}=\theta_{w}=0$, the ambiguity sets contain only the nominal distributions.
-The least-favorable covariances then coincide with the nominal ones, so the DRKF reduces exactly to the classical KF.
-Thus, Theorem˜5 recovers the standard optimality of the KF as a special case, demonstrating that the DRKF is a strict generalization of the classical optimal estimator to the DR regime.
-Overall, the results in this section confirm that the proposed noise-centric DRKF preserves the essential analytical structure of the classical KF while extending it to systems subject to distributional ambiguity.
-The filter remains stable, convergent, and minimax optimal, thereby bridging stochastic state estimation and DRO.
-These theoretical guarantees demonstrate that DRKF is not only practical but also provably reliable under distributional and long-horizon uncertainty.
+When $\theta_{v}=\theta_{w}=0$, the ambiguity sets reduce to the nominal noise distributions.
+In this case, the least-favorable covariances coincide with the nominal ones, and the DRKF reduces exactly to the classical KF.
+Thus, Theorem˜5 is consistent with the classical KF optimality and recovers it as a special case in the asymptotic regime.
 5 Experiments
-We conducted numerical experiments to evaluate the proposed filters and validate their theoretical properties.
-In Section˜5.1, we compare our methods against a wide range of time-varying and steady-state filters, assessing estimation accuracy, closed-loop performance in a trajectory-tracking task, and computational efficiency.
-We also examine how the ambiguity radii and the accuracy of the nominal noise distributions affect estimation performance.
-In Section˜5.2.2, we empirically verify the theoretical results of Section˜4.888All experiments were performed on a laptop equipped with an Intel Core Ultra 7 155H @ 3.80 GHz and 32 GB of RAM.
+Figure 1: Effect of $\theta$ on the average regret MSE under a) Gaussian and b) U-Quadratic noise distributions, averaged over 20 simulation runs.
+We evaluate the proposed DRKFs through numerical experiments that assess estimation accuracy, closed-loop performance, and computational efficiency, and that empirically validate the theoretical results in Section˜4.101010All experiments were performed on a laptop equipped with an Intel Core Ultra 7 155H @ 3.80 GHz and 32 GB of RAM.
 Source code is available at https://github.com/jangminhyuk/DRKF2025.
-Figure 1: Effect of $\theta$ on the average regret MSE under a) Gaussian and b) U-Quadratic noise distributions, averaged over 20 simulation runs.
 5.1 Comparison of Robust State Estimators
-We evaluate both the time-varying and steady-state DRKFs against the following benchmark state estimators:
+We begin by comparing the proposed time-varying and steady-state DRKFs with the following baseline estimators:
   1. (A)
@@
   5. (E)
-    DRKF with Wasserstein ambiguity on the joint law of $(x_{t},y_{t})$ [16]
+    DRKF with joint Wasserstein ambiguity on $(x_{t},y_{t})$ [26]
   6. (F)
-    DRKF with bicausal Wasserstein ambiguity [19]
+    DRKF with bicausal Wasserstein ambiguity [8] [^11] Following [8], we cap the number of internal iterations at 20.
   7. (G)
-    Time-varying DRKF with Wasserstein ambiguity on the prior state and measurement noise [21, Theorem 1]
+    Time-varying DRKF with Wasserstein ambiguity on the prior state and measurement noise [11, Thm. 1]
   8. (H)
-    Time-invariant version of (G) with stationary nominals [21, Section 4]
+    Time-invariant version of (G) with stationary nominals [11, Sect. 4]
   9. (I)
@@
     Steady-state DRKF (ours) [Algorithm˜2]
-We use the same symbol $\theta$ for the robustness parameter in all filters, although its role and scale differ across methods.
-For (C) and (D), $\theta$ is the risk-sensitivity coefficient, with (D) using $\theta$ as the risk-sensitivity parameter.
-For (C), we select $\theta$ up to the largest value for which the risk-sensitive Riccati mapping remains well defined [35, 36], i.e., $((\Sigma_{x,t}^{-})^{-1}-\theta D^{\top}D)$ remains positive definite with $D=I_{n_{x}}$.
-For (E) and (F), $\theta$ specifies the Wasserstein radius, and for (F), we follow [19] and cap the number of internal iterations at 20.
-For (G) and (H), we set $\theta_{x}=\theta_{v}=\theta$.
+We use the same symbol $\theta$ to denote the robustness parameter across all methods, although its interpretation differs.
+For risk-sensitive filters (C) and (D), $\theta$ is the risk parameter (with opposite signs for risk-averse/seeking variants), chosen within the stability range [16].
+For Wasserstein-based methods, $\theta$ denotes the ambiguity radius.
+For (G) and (H), we set $\theta_{x_{0}}=\theta_{v}=\theta$.
 In our proposed formulations, we set $\theta_{x_{0}}=\theta_{w}=\theta_{v}=\theta$ for (I), and $\theta_{w}=\theta_{v}=\theta$ for (J).
-To ensure the fairness, we sweep each method’s admissible range of $\theta$ and report performance at its best tuning.
+To ensure fairness, we sweep each method over its admissible range of $\theta$ and report performance at the best tuned value.
+[^12] In practice, DRKF radii $\theta$ can be selected from data, following standard approaches in the DRO literature, such as cross-validation or bootstrapping (e.g., [21, 4]).
 5.1.1 Estimation Accuracy
+We first consider a linear time-invariant system, subject to inaccuracies in both the process and measurement noise distributions:
+[…]
+We test two noise models: $(i)$ Gaussian, with $x_{0}\sim\mathcal{N}(0,0.01I_{4})$, $w_{t}\sim\mathcal{N}(0,0.01I_{4})$, $v_{t}\sim\mathcal{N}(0,0.01I_{2})$, and $(ii)$ U-Quadratic, with $x_{0}\sim\mathcal{UQ}([-0.1,0.1]^{4})$, $w_{t}\sim\mathcal{UQ}([-0.1,0.1]^{4})$, $v_{t}\sim\mathcal{UQ}([-0.1,0.1]^{2})$.
+Nominal covariances are learned from only 10 measurement samples using the expectation-maximization (EM) method in [8], creating a significant mismatch between nominal and true noise statistics.
+The estimation horizon is $T=50$.
+Fig.˜1 shows the effect of $\theta\in[10^{-2},10^{1}]$ on the average regret MSE under both noise models, averaged over 2020 runs.
+The regret MSE is defined as the difference between the filter’s MSE and that of the KF using the true noise distributions.
+The results show that both proposed DRKFs consistently achieve the lowest tuned regret and remain competitive across a wide range of $\theta$.
+In particular, (I) achieves performance closest to the optimal estimator under both poorly estimated Gaussian and non-Gaussian noise.
+In contrast, the BCOT-based DRKF (F) exhibits numerical instability and discontinuous performance across $\theta$, even when increasing the internal iteration limit to 50.
+The risk-sensitive filters (C) and (D) become overly conservative for large $\theta$, leading to degraded accuracy.
 Figure 2: Effect of the number of samples used to construct the nominal distributions on the average MSE under Gaussian noise, averaged over 10 runs.
-We begin with a linear time-invariant system (no control input) subject to inaccuracies in both the process and measurement noise distributions:
-[…]
-We consider two noise models: $(i)$ Gaussian, with $x_{0}\sim\mathcal{N}(0,0.01I_{4})$, $w_{t}\sim\mathcal{N}(0,0.01I_{4})$, $v_{t}\sim\mathcal{N}(0,0.01I_{2})$, and $(ii)$ U-Quadratic, with $x_{0}\sim\mathcal{UQ}([-0.1,0.1]^{4})$, $w_{t}\sim\mathcal{UQ}([-0.1,0.1]^{4})$, $v_{t}\sim\mathcal{UQ}([-0.1,0.1]^{2})$.
-The nominal covariances $\hat{\Sigma}_{w}$ and $\hat{\Sigma}_{v}$ are learned from only 10 input-output samples using the expectation-maximization (EM) method in [19].
-This induces a significant distributional mismatch, highlighting the challenge of learning accurate nominal models from limited data.
-The estimation horizon is set to 50 time steps.
-Fig.˜1 illustrates the effect of $\theta\in[10^{-2},10^{1}]$ on the average regret MSE for each filter under Gaussian and U-Quadratic noise, averaged over 20 runs.
-The regret MSE is defined as the difference between the filter’s MSE and that of the KF using the true noise distributions, which serves as the MMSE-optimal benchmark.
-The results show that both our time-varying and steady-state DRKFs consistently achieve the lowest tuned regret and remain competitive across a broad range of $\theta$.
-In particular, (I) achieves performance closest to the optimal estimator under both poorly estimated Gaussian and non-Gaussian noise.
-In comparison, (F) exhibits numerical instability and performance discontinuities across $\theta$, even when the internal iteration limit is increased to 50.
-The risk-sensitive filters (C) and (D) become overly conservative for large $\theta$, leading to degraded performance.
-These results highlight the limitations of existing methods, whereas our DRKFs maintain stable performance and a favorable robustness-accuracy tradeoff.
-Fig.˜2 shows the average MSE as a function of the number of samples used to construct the nominal noise distributions.
-Each filter is optimally tuned over $\theta\in[10^{-2},10^{1}]$, and we vary the number of input-output samples used in the EM procedure so that small sample sizes correspond to less accurate nominal distributions.
-As the number of samples increases, the average MSE decreases for all filters.
-Across all sample sizes, our DRKF consistently achieves the lowest error, with a particularly clear advantage when only limited data are available.
 Figure 3: Average regret MSE of the proposed steady-state DRKF under Gaussian noise, as a function of $\theta_{w}$ and $\theta_{v}$, averaged over 100 runs.
-Fig.˜3 shows the average MSE regret of the steady-state DRKF (J) as a function of the ambiguity radii $\theta_{w}$ and $\theta_{v}$ over $10^{-2},10^{1}]$, averaged over 100 runs.
-Both parameters significantly influence performance: the regret is large when both radii are small, reflecting insufficient robustness against distributional mismatch.
-Increasing either $\theta_{w}$ or $\theta_{v}$ improves performance up to an optimal range, beyond which excessively large radii make the estimator overly conservative.
-This surface highlights the need to balance $\theta_{w}$ and $\theta_{v}$ in practice.
+Figure 4: Effect of $\theta$ on the average MPC cost under a) Gaussian and b) nonzero-mean U-Quadratic noise, averaged over 20 simulation runs.
+Figure 5: 2D trajectories averaged over 200 simulations.
+Each filter uses its optimal $\theta\in[10^{-2},10^{1}]$ minimizing the MPC cost.
+The dashed black line is the desired trajectory; colored curves show the mean, and shaded tubes indicate $\pm 1$ standard deviation.
+Fig.˜2 further examines performance as a function of the number of samples used to estimate the nominal noise statistics.
+[^13] Each filter is optimally tuned over $\theta\in[10^{-2},10^{1}]$, and we vary the number of input-output samples used in the EM procedure so that small sample sizes correspond to less accurate nominal distributions.
+Across all sample sizes, the proposed DRKFs achieve the lowest MSE, with particularly clear advantages in the low-data regime.
+Fig.˜3 illustrates the sensitivity of the steady-state DRKF to the ambiguity radii $(\theta_{w},\theta_{v})$.
+Insufficient robustness (small radii) leads to large regret, while overly large radii induce excessive conservatism, highlighting the need to balance $\theta_{w}$ and $\theta_{v}$.
 5.1.2 Trajectory Tracking Task
-To assess the practical performance of our filters, we applied them to a 2D trajectory-tracking task, following the setup in [21, Section 5.2]:
+Table 1: Mean and standard deviation of the acceleration-command magnitude $\|u_{t}\|_{2}$ for each filter, averaged over 200 simulations.
+(A) (B) (E) (F) (G) (H) (I) (J)
+$\|u_{t}\|_{2}$ 5.69 (7.44) 5.91 (7.70) 3.29 (1.93) 2.70 (1.78) 3.37 (2.02) 3.30 (1.99) 2.18 (1.63) 2.20 (1.64)
+We next evaluate closed-loop performance on a 2D trajectory-tracking task, following the setup in [11, Sect. 5.2]:
 $\begin{split}x_{t+1}&=\begin{bmatrix}I_{2}&\Delta tI_{2}\\ \mathbf{0}_{2\times 2}&I_{2}\end{bmatrix}x_{t}+\begin{bmatrix}0.5(\Delta t)^{2}I_{2}\\ \Delta tI_{2}\end{bmatrix}u_{t}+w_{t}\\ y_{t}&=\begin{bmatrix}I_{2}&\mathbf{0}_{2\times 2}\end{bmatrix}x_{t}+v_{t},\end{split}$
@@
 The state $x_{t}=[p_{t}^{x},p_{t}^{y},v_{t}^{x},v_{t}^{y}]^{\top}$ collects the planar position and velocity, and the control input $u_{t}=[a_{t}^{x},a_{t}^{y}]^{\top}$ represents acceleration commands.
-We consider two noise settings: $(i)$ Gaussian, with $x_{0}\sim\mathcal{N}(0,0.02I_{4})$, $w_{t}\sim\mathcal{N}(0,0.02I_{4})$, $v_{t}\sim\mathcal{N}(0,0.02I_{2})$, and $(ii)$ nonzero-mean U-Quadratic, with $x_{0}\sim\mathcal{UQ}([-0.1,0.1]^{4})$, $w_{t}\sim\mathcal{UQ}([-0.1,0.2]^{4})$, $v_{t}\sim\mathcal{UQ}([-0.2,0.1]^{2})$, simulated over $10\,\text{s}$.
-For the nominal distributions, we fit an EM model using only $1\,\text{s}$ of input-output data, yielding noticeably inaccurate nominal statistics and reflecting the practical challenge of estimating noise models from limited observations.
-Figure 4: Effect of $\theta$ on the average MPC cost under a) Gaussian and b) nonzero-mean U-Quadratic noise, averaged over 20 simulation runs.
-Figure 5: 2D trajectories averaged over 200 simulations under Gaussian noise.
-Each filter uses its optimal parameter $\theta\in[10^{-2},10^{1}]$ minimizing the MPC cost.
-The black dashed line denotes the desired trajectory; colored curves show the mean, and shaded tubes indicate $\pm 1$ standard deviation.
-Table 1: Mean and standard deviation of the acceleration-command magnitude $\|u_{t}\|_{2}$ for each filter, averaged over 200 simulations.
-(A) (B) (E) (F) (G) (H) (I) (J)
-$\|u_{t}\|_{2}$ 5.69 (7.44) 5.91 (7.70) 3.29 (1.93) 2.70 (1.78) 3.37 (2.02) 3.30 (1.99) 2.18 (1.63) 2.20 (1.64)
-All methods use the same observer-based MPC controller (e.g., [40]), with the current state estimate $\bar{x}_{t}$ used for prediction.
+All methods use the same observer-based MPC controller, with the current state estimate $\bar{x}_{t}$ provided by each filter.
 The controller solves a finite-horizon tracking problem with horizon $10$ and weights $Q=\mathrm{diag}([10,10,1,1])$ and $R=0.1I_{2}$, where $Q$ penalizes tracking error and $R$ penalizes control effort.
 The first control input $u_{t}$ is applied, and the optimization is repeated at the next step using $\bar{x}_{t+1}$.
-The MPC model, horizon, and weights are identical across all methods; the only component that changes is the state estimator, where the Kalman filter is replaced with each candidate estimator.
-Fig.˜4 shows the effect of $\theta$ on the average MPC cost, defined as the closed-loop quadratic sum of tracking error and input energy weighted by $Q$ and $R$.
-Consistent with Section˜5.1.1, both our time-varying and steady-state DRKFs achieve the lowest cost when optimally tuned, even in the presence of non-Gaussian disturbances and nonzero-mean noise.
-Fig.˜5 displays the mean and dispersion of the closed-loop trajectories obtained with the best-tuned radius $\theta\in[10^{-2},10^{1}]$ for each filter.
-The standard KFs (A) and (B) produce wider uncertainty tubes, indicating higher estimation variance.
-The BCOT-based DRKF (F) reduces dispersion but introduces greater bias, especially during sharp turns.
-The remaining DRKFs–(E), (G), and (H)–exhibit accurate tracking with modest dispersion.
-Our time-varying and steady-state DRKFs deliver the most reliable performance, achieving the lowest dispersion and bias across runs.
-Although mean trajectories appear similar across DRKFs, the required control effort differs substantially.
-Table˜1 reports the time- and run-averaged acceleration magnitude $\|u_{t}\|_{2}$, which shows that our DRKFs achieve lower control energy among all methods.
-5.1.3 Computation Time
-The online computational cost of all filters is similar to that of the classical KF, since each estimator computes its gains offline.
-A key advantage of the steady-state DRKF is that it requires solving only a single SDP offline, whereas time-varying DRKFs must solve $T$ sequential SDPs.
-Figure 6: Average offline computation time as a function of the horizon length $T$ for a system with $n_{x}=4$, $n_{y}=2$, averaged over 10 runs.
+The true noise distributions are $x_{0}\sim\mathcal{N}(0,0.02I_{4})$, $w_{t}\sim\mathcal{N}(0,0.02I_{4})$, and $v_{t}\sim\mathcal{N}(0,0.02I_{2})$, simulated over $10\,\text{s}$.
+For the nominal distributions, we fit an EM model using only $1\,\text{s}$ of input-output data, yielding noticeably inaccurate nominal statistics and reflecting the practical challenge of estimating noise models from limited observations.
+Figure 6: Average offline computation time as a function of the horizon length $T$ for $n_{x}=4$, $n_{y}=2$, averaged over 10 runs.
 Table 2: Average offline computation time of the steady-state DRKF as a function of $n_{x}=n_{y}=n$ (10 runs).
@@
 Time (s) 0.0600 0.1794 0.4736 1.2414 2.2905 4.8900 8.3890 12.4232
-Fig.˜6 compares the offline computation time of (E)(J) as a function of the horizon length $T$.
-We use the same setup as in Section˜5.1.1 with $\theta=1.0$ and Gaussian noise.
-Time-varying DRKFs exhibit linear growth in computation time with respect to $T$, with (F) being particularly expensive.
-In contrast, the steady-state DRKFs (G) and (J) maintain constant runtime regardless of $T$, demonstrating excellent scalability for long-horizon tasks.
+Fig.˜4 shows the effect of $\theta$ on the average MPC cost, defined as the closed-loop quadratic sum of tracking error and input energy weighted by $Q$ and $R$.
+Consistent with Section˜5.1.1, both our time-varying and steady-state DRKFs achieve the lowest cost when optimally tuned, even in the presence of non-Gaussian disturbances and nonzero-mean noise.
+Fig.˜5 displays the mean trajectories and uncertainty envelopes obtained with the best-tuned radius $\theta\in[10^{-2},10^{1}]$ for each filter.
+The standard KFs (A) and (B) produce wider uncertainty tubes, indicating higher estimation variance.
+The BCOT-based DRKF (F) reduces dispersion but introduces greater bias, especially during sharp turns.
+Other DRKFs–(E), (G), and (H)–achieve improved tracking, while the proposed time-varying and steady-state DRKFs exhibit the lowest dispersion and bias.
+Although mean trajectories are similar across DRKFs, control effort differs substantially. As shown in Table˜1, the proposed DRKFs require the lowest average control energy among all methods.
+5.1.3 Computation Time
+Fig.˜6 compares the offline computation time of (E)(J) as a function of the horizon length TT.
+[^14] We use the same setup as in Section 5.1.1 with $\theta=1.0$ and Gaussian noise.
+Time-varying DRKFs exhibit linear growth in offline computation time with respect to $T$, with (F) being particularly expensive.
+In contrast, the steady-state DRKFs (H) and (J) maintain constant runtime, as they require solving only a single SDP offline.
+This demonstrates the scalability of the proposed steady-state DRKF for long-horizon applications.
 Finally, Table˜2 evaluates scalability with respect to the system dimension.
@@
 5.2 Validation of Theoretical Properties
-5.2.1 DRKF Sandwich Properties
-Theorem˜2 states that, at each time step, the DRKF prior and posterior covariances lie between those of two KFs run with extremal noise levels from (17).
-We illustrate this sandwich property on a concrete 2D example.
-Consider the system
-$A=0.95\begin{bmatrix}\cos(\pi/8)&-\sin(\pi/8)\\ \sin(\pi/8)&\cos(\pi/8)\end{bmatrix},\quad C=\begin{bmatrix}1.0&0.1\\ 0.1&1.0\end{bmatrix}$,
-with horizon $T=20$ and ambiguity radii $\theta_{w}=\theta_{v}=\theta_{x_{0}}=0.1$. The initial prior covariance is
-$\Sigma_{x,0}^{-}=\begin{bmatrix}0.2&0.05\\ 0.05&0.1\end{bmatrix}$.
+Figure 7: 95% confidence ellipses of the posterior covariances for LOW-KF, DRKF, and HIGH-KF at $t\in\{0,1,5,20\}$.
+Figure 8: Three-dimensional tube visualization obtained by stacking the $95\%$ ellipses of the posterior covariances of LOW-KF, DRKF, and HIGH-KF for $t=0,\dots,20$.
+5.2.1 Covariance Sandwich Property
+We illustrate the covariance sandwich property of Theorem˜2 using a 2D example with
+$A=0.95\begin{bmatrix}\cos(\pi/8)&-\sin(\pi/8)\\ \sin(\pi/8)&\cos(\pi/8)\end{bmatrix},\qquad C=\begin{bmatrix}1.0&0.1\\ 0.1&1.0\end{bmatrix}$,
+horizon $T=20$, and ambiguity radii $\theta_{w}=\theta_{v}=\theta_{x_{0}}=0.1$.
+The initial prior covariance is $\Sigma_{x,0}^{-}=\begin{bmatrix}0.2&0.05\\ 0.05&0.1\end{bmatrix}$.
 We use time-varying nominal covariances of the form $\hat{\Sigma}_{w,t}=a_{t}\hat{\Sigma}_{w}$, $\hat{\Sigma}_{v,t}=a_{t}\hat{\Sigma}_{v}$, constructed by scaling the base matrices
@@
 These covariances are used at each step $t$ to generate the lower-bound KF (LOW-KF), upper-bound KF (HIGH-KF), and DRKF posterior covariance trajectories.
-To visualize the uncertainty, we use the $95\%$ confidence ellipse $\mathcal{E}(\Sigma)\coloneqq\{x\in\mathbb{R}^{2}:x^{\top}\Sigma^{-1}x\leq\chi^{2}_{2,0.95}\}$ with $\chi^{2}_{2,0.95}=5.991$.
+To visualize the uncertainty, we use the $95\%$ confidence ellipse $\mathcal{E}(\Sigma)\coloneqq\{x\in\real{2}:x^{\top}\Sigma^{-1}x\leq\chi^{2}_{2,0.95}\}$ with $\chi^{2}_{2,0.95}=5.991$.
 This region contains $95\%$ of a zero-mean Gaussian with covariance $\Sigma$.
-By Theorem˜2, $\mathcal{E}(\underline{\Sigma}_{x,t})\subseteq\mathcal{E}(\Sigma_{x,t})\subseteq\mathcal{E}(\overline{\Sigma}_{x,t}),$ so the DRKF posterior ellipse must lie between the LOW-KF and HIGH-KF ellipses at every time step.
-Figure 7: 95% confidence ellipses of the posterior covariances for LOW-KF, DRKF, and HIGH-KF at $t\in\{0,1,5,20\}$.
-Figure 8: Three-dimensional tube visualization obtained by stacking the $95\%$ ellipses of the posterior covariances of LOW-KF, DRKF, and HIGH-KF for $t=0,\dots,20$.
-As shown in Figs.˜7 and 8, the ellipses at representative time steps confirm the predicted nesting.
-The DRKF posterior covariance is consistently sandwiched between the LOW-KF and HIGH-KF bounds, providing a clear and interpretable robustness envelope throughout the horizon.
-5.2.2 Convergence Properties
-To evaluate the convergence behavior of the proposed filter, we consider the linear system from [37]:
-$A=\begin{bmatrix}0.1&1\\ 1&-1\end{bmatrix},\;C=\begin{bmatrix}1&-1\end{bmatrix}$.
-Figure 9: Relative difference between $\operatorname{Tr}[\Sigma_{x,t}]$ and $\operatorname{Tr}[\Sigma_{x,\infty}^{*}]$ over time.
-We set the nominal covariances to $\hat{\Sigma}_{w}=I_{2}$ and $\hat{\Sigma}_{v}=1$.
-Following the procedure in Section˜4.2, we first compute $\phi_{N}$ with $N=8$ and $q=20$, and set $\phi_{v}=\phi_{x}=\phi_{N}/2$.
-We then choose the largest pair $(\theta_{v},\theta_{w})$ permitted by Theorem˜3.
-To quantify convergence of the time-varying DRKF toward the steady-state solution, we track the relative error $|(\operatorname{Tr}[\Sigma_{x,\infty}^{*}]-\operatorname{Tr}[\Sigma_{x,t}])/\operatorname{Tr}[\Sigma_{x,\infty}^{*}]|$.
-As shown in Fig.˜9, this quantity decays approximately linearly on a logarithmic scale, demonstrating exponential convergence of the time-varying DRKF to the steady-state DRKF.
-6 Conclusions and Future Work
-This work presented a noise-centric formulation of DRKF in which uncertainty in the process and measurement noises is modeled directly via Wasserstein ambiguity sets.
-In contrast to prior DRKF formulations, the proposed approach preserves dynamic consistency and yields tractable spectral bounds on the least-favorable covariances.
-The resulting recursion retains the analytical structure of the classical KF while providing robustness to distributional errors.
-We established the existence, uniqueness, and convergence of the steady-state DRKF and proved that it is asymptotically minimax optimal with respect to the worst-case mean-square error.
-Notably, the steady-state solution is obtained from a single stationary SDP, yielding a constant-gain filter with Kalman-level online complexity and strong robustness guarantees.
-Empirical studies corroborate the theory, demonstrating improved estimation accuracy, tighter and more reliable uncertainty quantification, and superior closed-loop performance in control tasks.
-Future directions include extending the framework to nonlinear systems and developing adaptive mechanisms for online calibration of the ambiguity set radius.
-It is also of interest to relax some of the structural assumptions underlying our convergence analysis (e.g., independence and range conditions on the process noise), to incorporate simultaneous model uncertainty in the system matrices, and to explore temporally coupled ambiguity sets that better capture long-range noise correlations while preserving computational tractability.
+By Theorem˜2, $\mathcal{E}(\underline{\Sigma}_{x,t})\subseteq\mathcal{E}(\Sigma_{x,t})\subseteq\mathcal{E}(\overline{\Sigma}_{x,t})$.
+As shown in Fig.˜7, the ellipses at representative time steps confirm the predicted nesting, providing a clear and interpretable robustness envelope.
+5.2.2 Convergence Property
+Finally, we examine convergence of the time-varying DRKF to its steady-state counterpart using the example from [34]: $A=\begin{bmatrix}0.1&1\\ 1&-1\end{bmatrix},\quad C=\begin{bmatrix}1&-1\end{bmatrix}$.
+[^15] We set the nominal covariances to $\hat{\Sigma}_{w}=I_{2}$ and $\hat{\Sigma}_{v}={1}$.
+and the ambiguity set radii $\theta_{x_{0}}=\theta_{w}=\theta_{v}=0.1$.
+Fig.˜9 shows that the relative error between $\operatorname{Tr}[\Sigma_{x,t}]$ and $\operatorname{Tr}[\Sigma_{x,\infty}^{*}]$ decays approximately linearly on a logarithmic scale, confirming exponential convergence, consistent with the theoretical analysis.
+Figure 9: Relative difference between $\operatorname{Tr}[\Sigma_{x,t}]$ and $\operatorname{Tr}[\Sigma_{x,\infty}^{*}]$.
+6 Conclusions
+We developed a DRKF based on a noise-centric Wasserstein ambiguity model.
+The proposed formulation preserves the structure of the classical KF, admits explicit spectral bounds, and ensures dynamic consistency under distributional uncertainty.
+We established existence, uniqueness, and convergence of the steady-state solution, proved asymptotic minimax optimality, and showed that the steady-state filter is obtained from a single stationary SDP with Kalman-level online complexity.
+Future directions include extending the framework to nonlinear systems, developing adaptive mechanisms for online calibration of the ambiguity set radius, incorporating model uncertainty in the system matrices, and exploring temporally coupled ambiguity sets.
 Appendix A Eigenvalue Lower Bounds for DRSE Maximizers
 Lemma 5.
-For $t=0$, the DRSE formulation (4) admits a maximizer pair $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ such that
-$\Sigma_{x,0}^{-,*}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\Sigma_{v,0}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}.$ (33)
-For $t>0$, the DRSE formulation (5) admits a maximizer pair $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ such that
-$\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}.$ (34)
+Consider the problems (4) and (5) with the explicit eigenvalue lower-bound constraints removed. Then:
+For $t=0$, (4) admits a maximizer $(\Sigma_{x,0}^{-,*},\Sigma_{v,0}^{*})$ such that
+$\Sigma_{x,0}^{-,*}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},\quad\Sigma_{v,0}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,0})I_{n_{y}}.$ (31)
+For $t>0$, (5) admits a maximizer $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ such that
+$\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\quad\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}.$ (32)
 Proof.
-For $t=0$, (33) is shown in [17, Theorem 3.1]. We now prove (34) by reformulating (5) in a form to which [17, Lemma A.3] applies. The objective of (5) is the trace of the posterior covariance matrix:
-$\begin{split}\operatorname{Tr}[\Sigma_{x,t}]&=\operatorname{Tr}[\Sigma_{x,t}^{-}-\Sigma_{x,t}^{-}C^{\top}(C\Sigma_{x,t}^{-}C^{\top}+\Sigma_{v,t})^{-1}C\Sigma_{x,t}^{-}]\\ &=\inf_{K_{t}}\operatorname{Tr}[(I_{n_{x}}-K_{t}C_{t})\Sigma_{x,t}^{-}(I_{n_{x}}-K_{t}C_{t})^{\top}+K_{t}\Sigma_{v,t}K_{t}^{\top}]\\ &=\inf_{K_{t}}(\langle(I_{n_{x}}-K_{t}C_{t})^{\top}(I_{n_{x}}-K_{t}C_{t}),\Sigma_{x,t}^{-}\rangle+\langle K_{t}^{\top}K_{t},\Sigma_{v,t}\rangle)\end{split}$
-where the last equality uses the inner product $\langle A,B\rangle:=\operatorname{Tr}[A^{\top}B]$.
-Given $\Sigma_{x,t-1}$, the deterministic part $A_{t}\Sigma_{x,t-1}A_{t}^{\top}$ is fixed, so $\Sigma_{x,t}^{-}=\bar{\Sigma}_{t}+\Sigma_{w,t-1}$, $\bar{\Sigma}_{t}\coloneqq A_{t}\Sigma_{x,t-1}A_{t}^{\top}\succeq 0$.
-Substituting this into the objective yields
-$\begin{split}\operatorname{Tr}[\Sigma_{x,t}]&=\inf_{K_{t}}\bigg(\langle(I_{n_{x}}-K_{t}C_{t})^{\top}(I_{n_{x}}-K_{t}C_{t}),\Sigma_{w,t-1}\rangle\\ &\qquad+\langle(I_{n_{x}}-K_{t}C_{t})^{\top}(I_{n_{x}}-K_{t}C_{t}),\bar{\Sigma}_{t}\rangle+\langle K_{t}^{\top}K_{t},\Sigma_{v,t}\rangle\bigg)\end{split}$ (35)
-First, we fix $\Sigma_{v,t}$ and view the objective as a function of $\Sigma_{w,t-1}$. From (35) this dependence can be written as
-$\inf_{K_{t}}\Big(\langle(I_{n_{x}}-K_{t}C_{t})^{\top}(I_{n_{x}}-K_{t}C_{t}),\Sigma_{w,t-1}\rangle+f_{1,t}(K_{t})\Big),$ (36)
-where $f_{1,t}(K_{t})$ collects the remaining terms, and is convex and continuous in $K_{t}$.
-Define the convex set $\mathcal{C}\coloneqq\{I_{n_{x}}-K_{t}C_{t}:K_{t}\in\mathbb{R}^{n_{x}\times n_{y}}\}$.
-For each $L\in\mathcal{C}$, define
-$f(L)\coloneqq\inf_{K_{t}:I_{n_{x}}-K_{t}C_{t}=L}f_{1,t}(K_{t}).$
-The function $f$ is convex and continuous since $f_{1,t}$ is quadratic in $K_{t}$ and we take the infimum over an affine subspace. Then (36) can be rewritten as
-$\inf_{L\in\mathcal{C}}\big(\langle L^{\top}L,\Sigma_{w,t-1}\rangle+f(L)\big).$
-Together with the Bures-Wasserstein constraint on $\Sigma_{w,t-1}$ in (5), this matches the setting of [17, Lemma A.3] with variable $\Sigma=\Sigma_{w,t-1}$.
-Hence there exists a maximizer $\Sigma_{w,t-1}^{*}$ such that $\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}}$.
-Next, we fix $\Sigma_{w,t-1}$ (and thus $\Sigma_{x,t}^{-}$) and view the objective as a function of $\Sigma_{v,t}$. From (35) the dependence on $\Sigma_{v,t}$ is
-$\inf_{K_{t}}\Big(\langle K_{t}^{\top}K_{t},\Sigma_{v,t}\rangle+f_{2,t}(K_{t})\Big),$
-where $f_{2,t}(K_{t})$ collects the remaining terms and is convex and continuous in $K_{t}$.
-This is exactly of the form in [17, Lemma A.3] with variable $\Sigma=\Sigma_{v,t}$.
-Applying [17, Lemma A.3] again yields a maximizer $\Sigma_{v,t}^{*}$ such that $\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}$.
-By choosing such partial maximizers when forming a joint maximizer of (5), we obtain (34), completing the proof.
+For $t=0$, (31) is shown in [23, Thm.
+3.1].
+We prove (32) for $t>0$.
+Fix $\Sigma_{x,t-1}$, so that $\Sigma_{x,t}^{-}=A_{t-1}\Sigma_{x,t-1}A_{t-1}^{\top}+\Sigma_{w,t-1}$.
+Using the standard identity $\operatorname{Tr}[\Sigma_{x,t}]=\inf_{K_{t}}\operatorname{Tr}\left[(I-K_{t}C_{t})\Sigma_{x,t}^{-}(I-K_{t}C_{t})^{\top}+K_{t}\Sigma_{v,t}K_{t}^{\top}\right]$, the objective can be written as
+$\operatorname{Tr}[\Sigma_{x,t}]=\inf_{K_{t}}\Big(\langle(I-K_{t}C_{t})^{\top}(I-K_{t}C_{t}),\,\Sigma_{w,t-1}\rangle+f_{t}(K_{t},\Sigma_{v,t})\Big),$
+where $\langle A,B\rangle\coloneqq\operatorname{Tr}[A^{\top}B]$ and $f_{t}$ collects the remaining terms (independent of $\Sigma_{w,t-1}$) and is convex and continuous in $K_{t}$.
+Let $(\bar{\Sigma}_{w,t-1},\bar{\Sigma}_{v,t})$ be a maximizer of (5) with the explicit eigenvalue bounds removed, which exists since the feasible set (Bures–Wasserstein balls intersected with $\mathbb{S}^{\cdot}_{+}$) is compact and the objective is continuous.
+Fixing $\Sigma_{v,t}=\bar{\Sigma}_{v,t}$, the maximization over $\Sigma_{w,t-1}$ matches the setting of [23, Lemma A.3], hence it admits a maximizer $\Sigma_{w,t-1}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}}$.
+Moreover, because $(\bar{\Sigma}_{w,t-1},\bar{\Sigma}_{v,t})$ is globally optimal, the fixed-$\Sigma_{v,t}$ subproblem achieves the same optimal value, so $(\Sigma_{w,t-1}^{*},\bar{\Sigma}_{v,t})$ is also a global maximizer.
+Now fix $\Sigma_{w,t-1}=\Sigma_{w,t-1}^{*}$, which in turn fixes $\Sigma_{x,t}^{-}=A_{t-1}\Sigma_{x,t-1}A_{t-1}^{\top}+\Sigma_{w,t-1}^{*}$.
+Then, the dependence on $\Sigma_{v,t}$ is of the form $\inf_{K_{t}}(\langle K_{t}^{\top}K_{t},\Sigma_{v,t}\rangle+\tilde{f}_{t}(K_{t}))$, so applying [23, Lemma A.3] again yields a maximizer $\Sigma_{v,t}^{*}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}}$, and $(\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*})$ remains globally optimal.
+This proves (32).
 Appendix B Proofs for Spectral Boundedness
 B.1 Proof of Proposition˜1
-Proof.
 By Lemma˜2, there exists $U\in O(n)$ such that $\|X^{1/2}-\hat{X}^{1/2}U\|_{F}\leq\theta.$ Since $\|\cdot\|_{2}\leq\|\cdot\|_{F}$, we also have
 $\|X^{1/2}-\hat{X}^{1/2}U\|_{2}\leq\theta.$ (33)
-For any matrices $A,B\in\real{n\times n}$, the singular-value perturbation bounds [41, Corollary 7.3.5] state that
-$\begin{split}&|\sigma_{\max}(A)-\sigma_{\max}(B)|\leq\|A-B\|_{2},\\ &|\sigma_{\min}(B)-\sigma_{\min}(A)|\leq\|A-B\|_{2}.\end{split}$ (38)
-Apply (LABEL:eqn:sig_per) with $A=X^{1/2}$ and $B=\hat{X}^{1/2}U$, and combine with (37) to obtain
-$\begin{split}&|\sigma_{\max}(X^{1/2})-\sigma_{\max}(\hat{X}^{1/2}U)|\leq\theta,\\ &|\sigma_{\min}(\hat{X}^{1/2}U)-\sigma_{\min}(X^{1/2})|\leq\theta\end{split}$
-Because $U$ is orthogonal, $X\succeq 0$ and $\hat{X}\succeq 0$, we have
-[…]
+For any matrices $A,B\in\real{n\times n}$, the singular-value perturbation bounds [9, Cor. 7.3.5] state that
+$\begin{split}&|\sigma_{\max}(A)-\sigma_{\max}(B)|\leq\|A-B\|_{2}\\ &|\sigma_{\min}(A)-\sigma_{\min}(B)|\leq\|A-B\|_{2}.\end{split}$ (34)
+Apply (LABEL:eqn:sig_per) with $A=X^{1/2}$ and $B=\hat{X}^{1/2}U$, and combine with (33) to obtain $|\sigma_{\max}(X^{1/2})-\sigma_{\max}(\hat{X}^{1/2}U)|\leq\theta$ and $|\sigma_{\min}(\hat{X}^{1/2}U)-\sigma_{\min}(X^{1/2})|\leq\theta$.
+Because $U$ is orthogonal, $X\succeq 0$ and $\hat{X}\succeq 0$, we have $\sigma_{\max}(\hat{X}^{1/2}U)=\sigma_{\max}(\hat{X}^{1/2})=\sqrt{\lambda_{\max}(\hat{X})}$ and $\sigma_{\min}(\hat{X}^{1/2}U)=\sigma_{\min}(\hat{X}^{1/2})=\sqrt{\lambda_{\min}(\hat{X})}$.
 Substituting these into the inequalities above yields
-$\sqrt{\lambda_{\max}(X)}=\sigma_{\max}(X^{1/2})\leq\sqrt{\lambda_{\max}(\hat{X})}+\theta,$
-$\sqrt{\lambda_{\min}(X)}=\sigma_{\min}(X^{1/2})\geq\sqrt{\lambda_{\min}(\hat{X})}-\theta$
-Squaring both sides and clipping the lower bound at zero yields the claim. ∎
+$\begin{split}\sqrt{\lambda_{\max}(X)}&=\sigma_{\max}(X^{1/2})\leq\sqrt{\lambda_{\max}(\hat{X})}+\theta,\\ \sqrt{\lambda_{\min}(X)}&=\sigma_{\min}(X^{1/2})\geq\sqrt{\lambda_{\min}(\hat{X})}-\theta.\end{split}$
+Squaring both sides and clipping the lower bound at zero yields the claim.
 B.2 Proof of Corollary˜2
-Proof.
 By Lemma˜1, the matrices $\Sigma_{v,t}^{*},\Sigma_{w,t}^{*}$ and $\Sigma_{x,0}^{-,*}$ belong to the Bures–Wasserstein balls centered at $\hat{\Sigma}_{v,t},\hat{\Sigma}_{w,t}$ and $\hat{\Sigma}_{x,0}^{-}$ with radii $\theta_{v},\theta_{w}$ and $\theta_{x_{0}}$, respectively.
 Hence, Proposition˜1 gives the corresponding upper bounds.
-The lower bounds follow directly from the constraints
-$\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}},\quad\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}},\quad\Sigma_{x,0}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}},$
-which are imposed in (4) and (5). Thus, any maximizers $\Sigma_{v,t}^{*}$, $\Sigma_{w,t-1}^{*}$, and $\Sigma_{x,0}^{-,*}$ necessarily satisfy these inequalities, giving the claimed lower bounds. ∎
+The lower bounds follow directly from the constraints $\Sigma_{v,t}\succeq\lambda_{\min}(\hat{\Sigma}_{v,t})I_{n_{y}},\Sigma_{w,t-1}\succeq\lambda_{\min}(\hat{\Sigma}_{w,t-1})I_{n_{x}}$, and $\Sigma_{x,0}^{-}\succeq\lambda_{\min}(\hat{\Sigma}_{x,0}^{-})I_{n_{x}}$, which are imposed in (4) and (5).
+Thus, any maximizers $\Sigma_{v,t}^{*}$, $\Sigma_{w,t-1}^{*}$, and $\Sigma_{x,0}^{-,*}$ necessarily satisfy these inequalities, giving the claimed lower bounds.
 B.3 Proof of Theorem˜2
-Proof.
-Define the posterior and prediction maps
-$\begin{split}&\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t})\coloneqq\Sigma_{x,t}^{-}-\Sigma_{x}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t})^{-1}C_{t}\Sigma_{x,t}^{-}\\ &\Phi(\Sigma_{x,t},\Sigma_{w,t})\coloneqq A_{t}\Sigma_{x,t}A_{t}^{\top}+\Sigma_{w,t}.\end{split}$
-The DRKF recursion is obtained by evaluating these maps at the least-favorable noise covariances:
-$\Sigma_{x,t}=\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t}^{*})\quad\text{and}\quad\Sigma_{x,t+1}^{-}=\Phi(\Sigma_{x,t},\Sigma_{w,t}^{*}).$
-Using the identity $\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t})=((\Sigma_{x,t}^{-})^{-1}+C_{t}^{\top}\Sigma_{v,t}^{-1}C_{t})^{-1}$ and the fact that $P_{1}\succeq P_{2}$ implies $P_{1}^{-1}\preceq P_{2}^{-1}$, the map $\Psi$ is monotone increasing in each argument.
+Define the posterior and prediction maps as
+$\begin{split}\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t})&\coloneqq\Sigma_{x,t}^{-}-\Sigma_{x,t}^{-}C_{t}^{\top}(C_{t}\Sigma_{x,t}^{-}C_{t}^{\top}+\Sigma_{v,t})^{-1}C_{t}\Sigma_{x,t}^{-}\\ \Phi(\Sigma_{x,t},\Sigma_{w,t})&\coloneqq A_{t}\Sigma_{x,t}A_{t}^{\top}+\Sigma_{w,t}.\end{split}$
+The DRKF recursion is obtained by evaluating these maps at the least-favorable noise covariances: $\Sigma_{x,t}=\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t}^{*})$ and $\Sigma_{x,t+1}^{-}=\Phi(\Sigma_{x,t},\Sigma_{w,t}^{*})$.
+For $\Sigma_{v,t}\succ 0$, the posterior covariance admits the variational representation
+$\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t})=\inf_{K_{t}}\Big[(I-K_{t}C_{t})\Sigma_{x,t}^{-}(I-K_{t}C_{t})^{\top}+K_{t}\Sigma_{v,t}K_{t}^{\top}\Big].$
+Hence, $\Psi$ is monotone increasing in each argument: if ${\Sigma_{x,t}^{-}}^{(1)}\succeq{\Sigma_{x,t}^{-}}^{(2)}$ and $\Sigma_{v,t}^{(1)}\succeq\Sigma_{v,t}^{(2)}$, then for any $K_{t}$ the corresponding quadratic form is larger, and taking the infimum preserves the Loewner order.
 The map $\Phi$ is also monotone since congruence with $A_{t}$ and addition of a PSD matrix preserve the Loewner order.
-Corollary˜2, together with the monotonicity of the posterior and prediction maps $\Psi$ and $\Phi$, implies by induction that the KF driven by $(\underline{\lambda}_{w,t}I_{n_{x}},\underline{\lambda}_{v,t}I_{n_{y}})$ produces covariances no larger than those of the DRKF, and the KF driven by $(\overline{\lambda}_{w,t}I_{n_{x}},\overline{\lambda}_{v,t}I_{n_{y}})$ produces covariances no smaller.
-∎
+Corollary˜2, together with the monotonicity of the posterior and prediction maps $\Psi$ and $\Phi$, implies by induction that, for all $t\geq 0$, the KF driven by $(\underline{\lambda}_{w,t}I_{n_{x}},\underline{\lambda}_{v,t}I_{n_{y}})$ produces covariances no larger than those of the DRKF, and the KF driven by $(\overline{\lambda}_{w,t}I_{n_{x}},\overline{\lambda}_{v,t}I_{n_{y}})$ produces covariances no smaller.
 Appendix C Proofs for Convergence
-C.1 Details of the Downsampled System
-For the convergence analysis, we define a downsampled state as $x_{t}^{d}=x_{tN}$ for some $N\in\mathbb{N}$. Then, the downsampled dynamics take the form
-$\begin{split}x_{t+1}^{d}&=A^{N}x_{t}^{d}+\mathcal{R}_{N}\bm{w}_{t}^{N}\\ y_{t}^{N}&=\mathcal{O}_{N}x_{t}^{d}+\mathcal{D}_{N}\bm{v}_{t}^{N}+\mathcal{H}_{N}\bm{w}_{t}^{N}\\ \bm{0}&=\mathcal{O}_{N}^{R}x_{t}^{d}+\bm{u}_{t}^{N}+\mathcal{L}_{N}\bm{w}_{t}^{N},\end{split}$
-where the block matrices collecting process and measurement noise are
-$\begin{split}\mathcal{R}_{N}&\coloneqq\left[\hat{\Sigma}_{w}^{\frac{1}{2}},\,A\hat{\Sigma}_{w}^{\frac{1}{2}},\,\ldots,\,A^{N-1}\hat{\Sigma}_{w}^{\frac{1}{2}}\right]\\ \mathcal{O}_{N}&\coloneqq\left[(CA^{N-1})^{\top},\,\ldots,\,(CA)^{\top},\,C^{\top}\right]^{\top}\\ \mathcal{O}_{N}^{R}&\coloneqq\left[(A^{N-1})^{\top},\,\ldots,\,(A)^{\top},\,I_{n_{x}}\right]^{\top}\\ \mathcal{D}_{N}&\coloneqq I_{N}\otimes\hat{\Sigma}_{v}^{\frac{1}{2}}.\end{split}$
-Here, $\mathcal{R}_{N}$ and $\mathcal{O}_{N}$ represent the $N$-block controllability and observability matrices.
-The block Hankel matrices $\mathcal{L}_{N}$ and $\mathcal{H}_{N}\coloneqq(I_{N}\otimes C)\mathcal{L}_{N}$ encode cross-correlations between observations and noise terms, with entries of $\mathcal{L}_{N}$ defined as $\mathcal{L}_{N}]_{ij}=A^{j-i-1}\hat{\Sigma}_{w}^{\frac{1}{2}}$ for $j>i$ and zero otherwise.
-Adopting the procedure in [36] and denoting the downsampled prior covariance matrix by $\Sigma_{x,t}^{-,d}:=\Sigma_{x,tN}^{-}$, the DR Riccati recursion associated with the downsampled model is given by (23), where
-$r_{\mathbb{D},t}^{d}(\Sigma_{x,t}^{-,d})\coloneqq\alpha_{N,t}\left[(\Sigma_{x,t}^{-,d})^{-1}+\Omega_{\bar{\Phi}_{N,t}}\right]^{-1}\alpha_{N,t}+W_{\bar{\Phi}_{N,t}},$
-with auxiliary matrices given by
-$\alpha_{N,t}$ $\coloneqq A^{N}-\mathcal{R}_{N}(\mathcal{H}_{N}^{\top}\mathcal{K}_{{\bar{\Phi}}_{N,t}}^{-1}\mathcal{O}_{N}+\mathcal{L}_{N}^{\top}\mathcal{K}_{{\bar{\Phi}}_{N,t}}^{-1}\mathcal{O}_{N}^{R})$
-$W_{\bar{\Phi}_{N,t}}$ $\coloneqq\mathcal{R}_{N}\mathcal{Q}_{\bar{\Phi}_{N,t}}\mathcal{R}_{N}^{\top}$
-$\mathcal{Q}_{\bar{\Phi}_{N,t}}$ $\coloneqq\left[I_{Nn_{x}}+\mathcal{H}_{N}^{\top}(\mathcal{D}_{N}\mathcal{D}_{N}^{\top})^{-1}\mathcal{H}_{N}-\mathcal{L}_{N}^{\top}\bar{\Phi}_{N,t}\mathcal{L}_{N}\right]^{-1}$
-$\Omega_{\bar{\Phi}_{N,t}}$ $\coloneqq\Omega_{N}+\mathcal{J}_{N}^{\top}S_{\bar{\Phi}_{N,t}}^{-1}\mathcal{J}_{N}$
-$\mathcal{J}_{N}$ $\coloneqq\mathcal{O}_{N}^{R}-\mathcal{L}_{N}\mathcal{H}_{N}^{\top}\left[\mathcal{D}_{N}\mathcal{D}_{N}^{\top}+\mathcal{H}_{N}\mathcal{H}_{N}^{\top}\right]^{-1}\mathcal{O}_{N}$
-$\Omega_{N}$ $\coloneqq\mathcal{O}_{N}^{\top}(\mathcal{D}_{N}\mathcal{D}_{N}^{\top}+\mathcal{H}_{N}\mathcal{H}_{N}^{\top})^{-1}\mathcal{O}_{N}$
-$S_{\bar{\Phi}_{N,t}}$ $\coloneqq\mathcal{L}_{N}(I_{Nn_{x}}+\mathcal{H}_{N}^{\top}(\mathcal{D}_{N}\mathcal{D}_{N}^{\top})^{-1}\mathcal{H}_{N})^{-1}\mathcal{L}_{N}^{\top}-\bar{\Phi}_{N,t}^{-1}$
-$\mathcal{K}_{\bar{\Phi}_{N,t}}$ $\coloneqq\begin{bmatrix}\mathcal{D}_{N}\mathcal{D}_{N}^{\top}+\mathcal{H}_{N}\mathcal{H}_{N}^{\top}&\mathcal{H}_{N}\mathcal{L}_{N}^{\top}\\ \mathcal{L}_{N}\mathcal{H}_{N}^{\top}&\mathcal{L}_{N}\mathcal{L}_{N}^{\top}-\bar{\Phi}_{N,t}^{-1}\end{bmatrix}$.
+C.1 Proof of Lemma˜3
+By Corollary˜2 and Assumption˜2, there exist constants $\underline{\lambda}_{w},\underline{\lambda}_{v}>0$ such that $\Sigma_{w,t}^{*}\succeq\underline{\lambda}_{w}I_{n_{x}}\succ 0$ and $\Sigma_{v,t}^{*}\succeq\underline{\lambda}_{v}I_{n_{y}}\succ 0$.
+This implies $\Sigma_{w,t}^{N,*}\succ 0$ and $\Sigma_{v,t}^{N,*}\succ 0$, and consequently,
+$\Sigma_{u,t}^{N}=\mathcal{H}_{N}\Sigma_{w,t}^{N,*}\mathcal{H}_{N}^{\top}+\Sigma_{v,t}^{N,*}\succ 0.$
+Using the matrix inversion lemma, the conditional covariance $\mathcal{Q}_{N,t}$ admits the equivalent representation
+$\mathcal{Q}_{N,t}=\big((\Sigma_{w,t}^{N,*})^{-1}+\mathcal{H}_{N}^{\top}(\Sigma_{v,t}^{N,*})^{-1}\mathcal{H}_{N}\big)^{-1}.$
+Since $(\Sigma_{w,t}^{N,*})^{-1}\succ 0$ and $\mathcal{H}_{N}^{\top}(\Sigma_{v,t}^{N,*})^{-1}\mathcal{H}_{N}\succeq 0$, it follows that $\mathcal{Q}_{N,t}\succ 0$.
+Moreover, $\mathcal{R}_{N}$ has full row rank for all $N\geq 1$.
+Hence, for any nonzero $x\in\real{n_{x}}$,
+$x^{\top}W_{N,t}x=(\mathcal{R}_{N}^{\top}x)^{\top}\mathcal{Q}_{N,t}(\mathcal{R}_{N}^{\top}x)>0,$
+which proves that $W_{N,t}\succ 0$.
+Finally, since $\Sigma_{u,t}^{N}\succ 0$, we have $(\Sigma_{u,t}^{N})^{-1}\succ 0$.
+By Assumption˜3, the observability of $(A,C)$ implies that $\mathcal{O}_{N}$ has full column rank for any $N\geq n_{x}$.
+Therefore, for any nonzero $x\in\real{n_{x}}$,
+$x^{\top}\Omega_{N,t}x=(\mathcal{O}_{N}x)^{\top}(\Sigma_{u,t}^{N})^{-1}(\mathcal{O}_{N}x)>0,$
+which establishes $\Omega_{N,t}\succ 0$ for all $N\geq n_{x}$.
 C.2 Proof of Proposition˜2
-Proof.
-The proof follows the argument in [36, Proposition 3.1].
-The downsampled DR Riccati mapping in (23) has the same structure as the robust Riccati mapping analyzed in [42, Theorem 5.3].
-According to this result, if the matrices $\Sigma_{x,t}^{-,d},\Omega_{\bar{\Phi}_{N,t}}$, and $W_{\bar{\Phi}_{N,t}}$ are positive definite, then $r_{\mathbb{D},t}^{d}$ is a contraction mapping.
-By Assumption˜1, $\hat{\Sigma}_{w}\succ 0$, which ensures that $\Sigma_{x,t}^{-,d}\succ 0$.
-Moreover, $\mathcal{Q}_{N,t}\succ 0$ whenever $0\preceq\bar{\Phi}_{N,t}\prec\tilde{\phi}_{N}I_{Nn_{x}}$, which, under the controllability assumption, is sufficient for the positive definiteness of $W_{\bar{\Phi}_{N,t}}$ for $N\geq n_{x}$.
-Similarly, $S_{\bar{\Phi}_{N,t}}\prec 0$ and the map $\bar{\Phi}_{N,t}\mapsto\Omega_{\bar{\Phi}_{N,t}}$ is monotonically non-increasing for $0\prec\bar{\Phi}_{N,t}\prec\tilde{\phi}_{N}I_{Nn_{x}}$.
-Under the observability assumption, the matrix $\Omega_{0}=\Omega_{N}$ is positive definite for $N\geq n_{x}$.
-Hence, by continuity, there exists $\phi_{N}\in(0,\tilde{\phi}_{N})$ such that $\Omega_{\bar{\Phi}_{N,t}}\succ 0$ for all $\bar{\Phi}_{N,t}$ satisfying (24).
-Thus all three matrices $\Sigma_{x,t}^{-,d},W_{\bar{\Phi}_{N,t}}$, and $\Omega_{\bar{\Phi}_{N,t}}$ are positive definite under (24), and therefore $r_{\mathbb{D},t}^{d}$ is a contraction for $t\geq q$.
-Since $r_{\mathbb{D}}$ is the corresponding time-unrolled operator, it inherits the same contraction property.
-∎
-C.3 Operator-norm bounds in a Bures–Wasserstein ball
-We bound the operator-norm deviation between a feasible covariance matrix and its nominal counterpart when both lie in a Bures–Wasserstein ball.
-Lemma 4.
-Fix $\hat{X}\succ 0$. For any $X\succ 0$ with $\mathcal{B}(X,\hat{X})\leq\theta$,
-$\|X-\hat{X}\|_{2}\;\leq\;\theta^{2}+2\theta\sqrt{\lambda_{\max}(\hat{X})}$.
-Proof.
-By Lemma˜2, there exists $U\in O(n)$ such that
-$\|X^{1/2}-\hat{X}^{1/2}U\|_{2}\leq\|X^{1/2}-\hat{X}^{1/2}U\|_{F}\leq\theta.$
-Let $E:=X^{1/2}-\hat{X}^{1/2}U$, so that $X^{1/2}=E+\hat{X}^{1/2}U$. Then,
-$X=\hat{X}+EE^{\top}+\hat{X}^{1/2}UE^{\top}+EU^{\top}\hat{X}^{1/2},$
-since $U$ is orthogonal. Taking operator norms and applying the triangle inequality and submultiplicativity of $|\cdot|_{2}$, we obtain
-$\|X-\hat{X}\|_{2}\leq\|E\|_{2}^{2}+2\|\hat{X}^{1/2}\|_{2}\|E\|_{2}.$
-Since $\|E\|_{2}\leq\theta$ and $\|\hat{X}^{1/2}\|_{2}=\sqrt{\lambda_{\max}(\hat{X})}$, the result follows. ∎
-C.4 Proof of Theorem˜3
-Proof.
-From (21) and the matrix inversion lemma,
-$\begin{split}\Phi_{t}&=C^{\top}\bigl(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\bigr)C+X_{t}-(I+X_{t}\Delta_{t})^{-1}X_{t}\\ &=C^{\top}(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1})C+X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}.\end{split}$
-We bound the two terms separately.
-$(i)$ Measurement uncertainty bound. Let $\delta_{v}\coloneq\|\Sigma_{v,t}^{*}-\hat{\Sigma}_{v}\|_{2}$. Since both matrices are invertible,
-$\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}=\hat{\Sigma}_{v}^{-1}(\Sigma_{v,t}^{*}-\hat{\Sigma}_{v})(\Sigma_{v,t}^{*})^{-1}.$
-By submultiplicativity, we have that
-$\|\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\|_{2}\leq\delta_{v}\|\hat{\Sigma}_{v}^{-1}\|_{2}\|(\Sigma_{v,t}^{*})^{-1}\|_{2}.$ (39)
-On the other hand, Weyl’s inequality gives
-$\lambda_{\min}(\Sigma_{v,t}^{*})\geq\lambda_{\min}(\hat{\Sigma}_{v})-\|\Sigma_{v,t}^{*}-\hat{\Sigma}_{v}\|_{2}=\lambda_{\min}(\hat{\Sigma}_{v})-\delta_{v},$
-so for $\delta_{v}<\lambda_{\min}(\hat{\Sigma}_{v})$,
-$\|\hat{\Sigma}_{v}^{-1}\|_{2}=\frac{1}{\lambda_{\min}(\hat{\Sigma}_{v})},$ (40)
+Fix $\theta_{w}$ and $\theta_{v}$ and let $\underline{\lambda}_{w}\coloneq\lambda_{\min}(\hat{\Sigma}_{w})$, $\underline{\lambda}_{v}\coloneq\lambda_{\min}(\hat{\Sigma}_{v})$, $\overline{\lambda}_{w}\coloneq\overline{\lambda}(\hat{\Sigma}_{w},\theta_{w})$, and $\overline{\lambda}_{v}\coloneq\overline{\lambda}(\hat{\Sigma}_{v},\theta_{v})$ as in Corollary˜2.
+Then, the least-favorable covariances satisfy $\underline{\lambda}_{w}I\preceq\Sigma_{w,t}^{N,*}\preceq\overline{\lambda}_{w}I$ and $\underline{\lambda}_{v}I\preceq\Sigma_{v,t}^{N,*}\preceq\overline{\lambda}_{v}I$.
+Hence,
+$\Sigma_{u,t}^{N}=\mathcal{H}_{N}\Sigma_{w,t}^{N,*}\mathcal{H}_{N}^{\top}+\Sigma_{v,t}^{N,*}\preceq(\overline{\lambda}_{w}\|\mathcal{H}_{N}\|_{2}^{2}+\overline{\lambda}_{v})I,$
+and $\Sigma_{u,t}^{N}\succeq\underline{\lambda}_{v}I$. Therefore,
+$\lambda_{\min}(\Omega_{N,t})\geq\frac{\sigma_{\min}(\mathcal{O}_{N})^{2}}{\lambda_{\max}(\Sigma_{u,t}^{N})}\geq\frac{\sigma_{\min}(\mathcal{O}_{N})^{2}}{\overline{\lambda}_{v}+\overline{\lambda}_{w}\|\mathcal{H}_{N}\|_{2}^{2}},$
+which implies
+$\|\Omega_{N,t}^{-1}\|_{2}\leq\frac{\overline{\lambda}_{v}+\overline{\lambda}_{w}\|\mathcal{H}_{N}\|_{2}^{2}}{\sigma_{\min}(\mathcal{O}_{N})^{2}}.$
+Next, using the definition of $\mathcal{G}_{N,t}$, we obtain
+$\|\mathcal{G}_{N,t}\|_{2}\leq\|\Sigma_{w,t}^{N,*}\|_{2}\|\mathcal{H}_{N}\|_{2}\,\|(\Sigma_{u,t}^{N})^{-1}\|_{2}\leq\frac{\overline{\lambda}_{w}}{\underline{\lambda}_{v}}\|\mathcal{H}_{N}\|_{2}.$
+Consequently,
+$\begin{split}\|\alpha_{N,t}\|_{2}&\leq\|A^{N}\|_{2}+\|\mathcal{R}_{N}\|_{2}\|\mathcal{G}_{N,t}\|_{2}\|\mathcal{O}_{N}\|_{2}\\ &\leq\|A^{N}\|_{2}+\frac{\overline{\lambda}_{w}}{\underline{\lambda}_{v}}\|\mathcal{R}_{N}\|_{2}\|\mathcal{H}_{N}\|_{2}\|\mathcal{O}_{N}\|_{2}.\end{split}$
+Finally, we have that
+$\begin{split}\lambda_{\min}(\mathcal{Q}_{N,t})&\geq\frac{1}{\lambda_{\max}((\Sigma_{w,t}^{N,*})^{-1})+\lambda_{\max}(\mathcal{H}_{N}^{\top}(\Sigma_{v,t}^{N,*})^{-1}\mathcal{H}_{N})}\\ &\geq\frac{1}{1/\underline{\lambda}_{w}+\|\mathcal{H}_{N}\|_{2}^{2}/\underline{\lambda}_{v}}.\end{split}$
+Since $\mathcal{R}_{N}\mathcal{R}_{N}^{\top}\succeq I$, we have $\lambda_{\min}(W_{N,t})\geq\lambda_{\min}(\mathcal{Q}_{N,t})$, and therefore
+$\|W_{N,t}^{-1}\|_{2}\leq 1/\underline{\lambda}_{w}+\|\mathcal{H}_{N}\|_{2}^{2}/\underline{\lambda}_{v}.$
+Combining the above bounds yields
+$\|M_{N,t}\|_{2}\leq\|\Omega_{N,t}^{-1}\|_{2}\|\alpha_{N,t}\|_{2}^{2}\|W_{N,t}^{-1}\|_{2}\leq\kappa_{N}(\theta_{v},\theta_{w}).$
+Substituting into (24) and using the monotonicity of $s\mapsto(\frac{\sqrt{s}}{1+\sqrt{1+s}})^{2}$ on $s\geq 0$ gives $\xi_{N,t}\leq\bar{\xi}_{N}(\theta_{v},\theta_{w})$; inequality (26) then ensures $\xi_{N,t}\leq\rho$ for all $t\geq 0$.
+C.3 Proof of Theorem˜3
+Fix $t\geq 0$ and $N\geq n_{x}$.
+By Lemma˜3, we have $W_{N,t}\succ 0$ and $\Omega_{N,t}\succ 0$.
+Hence, the downsampled DR Riccati mapping $r_{\mathbb{D},t}^{d}$ in (22) is well-defined on $\mathbb{S}^{n_{x}}_{++}$ and has the robust Riccati form considered in [14, Thm.
+5.3].
+Therefore, $r_{\mathbb{D},t}^{d}$ is a strict contraction with respect to Thompson’s part metric on $\mathbb{S}^{n_{x}}_{++}$, with contraction factor $\xi_{N,t}\in(0,1)$ satisfying the bound (24).
+By Proposition˜2, we have $\xi_{N,t}\leq\bar{\xi}_{N}(\theta_{v},\theta_{w})<1$ for all $t\geq 0$, and thus $r_{\mathbb{D},t}^{d}$ is uniformly contractive.
+Under Assumption˜2, the one-step DR Riccati update is stationary, so the downsampled map coincides with the NN-fold iterate of a single time-invariant map.
+The contraction argument for downsampled iterations in [36] then implies that the one-step DR Riccati recursion (19) converges to a unique fixed point $\Sigma_{x,\infty}^{-}\succ 0$.
+Finally, although the stage-wise SDP may admit multiple optimizers in $\Sigma_{v,t}^{*}$, the corresponding DR Kalman gain is unique.
+By Lemma˜1, for a fixed $\Sigma_{x,t}^{-}$ the stage-wise DRSE problem amounts to minimizing over $K$ the worst-case posterior MSE, with the adversary maximizing over $\Sigma_{v,t}$ in the SDP’s feasible set, denoted by $\mathcal{A}_{v,t}$.
+Using the standard KF covariance identity, the resulting worst-case one-step MSE is
+$\sup_{\Sigma_{v,t}\in\mathcal{A}_{v,t}}\operatorname{Tr}\big[(I-KC)\Sigma_{x,t}^{-}(I-KC)^{\top}+K\Sigma_{v,t}K^{\top}\big].$
+For each feasible $\Sigma_{v,t}$, the associated quadratic map
+$K\mapsto\operatorname{Tr}\big[(I-KC)\Sigma_{x,t}^{-}(I-KC)^{\top}+K\Sigma_{v,t}K^{\top}\big]$
+is $\underline{\lambda}_{v}$-strongly convex in $K$ (with respect to $\|\cdot\|_{F}$).
+Taking the supremum over the compact feasible set $\mathcal{A}_{v,t}$ preserves strong convexity, and hence the minimizer $K_{t}$ is unique.
+Moreover, this worst-case objective is continuous in $(\Sigma_{x,t}^{-},K)$, so the unique minimizer $K_{t}$ depends continuously on $\Sigma_{x,t}^{-}$.
+Therefore, $\Sigma_{x,t}^{-}\to\Sigma_{x,\infty}^{-}$ implies that the sequence $\{K_{t}\}$ converges.
+C.4 Proof of Corollary˜3
+For each $t\geq 1$, let $z_{t}:=(\Sigma_{x,t}^{-},\Sigma_{x,t},\Sigma_{w,t-1}^{*},\Sigma_{v,t}^{*},Y_{t},Z_{t})$ denote a primal optimal solution of the stage-wise SDP (7).
+By Theorem˜3, $(\Sigma_{x,t}^{-},\Sigma_{x,t})\to(\Sigma_{x,\infty}^{-},\Sigma_{x,\infty})$.
+Using the LMIs $\begin{bmatrix}\hat{\Sigma}_{w}&Y_{t}\\ Y_{t}^{\top}&\Sigma_{w,t-1}^{*}\end{bmatrix}\succeq 0$ and $\begin{bmatrix}\hat{\Sigma}_{v}&Z_{t}\\ Z_{t}^{\top}&\Sigma_{v,t}^{*}\end{bmatrix}\succeq 0$, together with $\hat{\Sigma}_{w}\succ 0$, $\hat{\Sigma}_{v}\succ 0$, Schur complements yield $\Sigma_{w,t-1}^{*}\succeq Y_{t}^{\top}\hat{\Sigma}_{w}^{-1}Y_{t}$ and $\Sigma_{v,t}^{*}\succeq Z_{t}^{\top}\hat{\Sigma}_{v}^{-1}Z_{t}$.
+Hence, it follows from Corollary˜2 that
+$\|Y_{t}\|_{2}^{2}\leq\|\hat{\Sigma}_{w}\|_{2}\ \|\Sigma_{w,t-1}^{*}\|_{2}\leq\|\hat{\Sigma}_{w}\|_{2}\ \bar{\lambda}_{w}$
 and
-$\|(\Sigma_{v,t}^{*})^{-1}\|_{2}\ =\frac{1}{\lambda_{\min}(\Sigma_{v,t}^{*})}\leq\frac{1}{\lambda_{\min}(\hat{\Sigma}_{v})-\delta_{v}}.$ (41)
-Substituting into (39) yields
-$\|\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\|_{2}\leq\frac{\delta_{v}}{\lambda_{\min}(\hat{\Sigma}_{v})\,(\lambda_{\min}(\hat{\Sigma}_{v})-\delta_{v})}.$
-Congruence with $C$ then gives
-$\bigl\|C^{\top}(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1})C\bigr\|_{2}\leq\|C\|_{2}^{2}\,\|\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1}\|_{2}\leq\frac{\|C\|_{2}^{2}\,\delta_{v}}{\lambda_{\min}(\hat{\Sigma}_{v})\,(\lambda_{\min}(\hat{\Sigma}_{v})-\delta_{v})}.$ (42)
-To guarantee $C^{\top}(\hat{\Sigma}_{v}^{-1}-(\Sigma_{v,t}^{*})^{-1})C\preceq\phi_{v}I$, it is sufficient that the right-hand side of (42) does not exceed $\phi_{v}$.
-This requirement is equivalent to $\delta_{v}\leq\delta_{v}^{\max}(\phi_{v})$ with $\delta_{v}^{\max}(\phi_{v})$ defined in (25).
-Moreover, any $\delta_{v}$ satisfying this bound automatically satisfies the condition $\delta_{v}<\lambda_{\min}(\hat{\Sigma}_{v})$ used in (40) and (41), ensuring consistency of the argument.
-$\delta_{v}\leq\theta_{v}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{v})}\theta_{v}.$
-Thus, it is sufficient to require that this upper bound does not exceed $\delta_{v}^{\max}(\phi_{v})$, which guarantees (27).
-Solving the resulting quadratic inequality in $\theta_{v}$ yields the explicit condition in (26).
-$(ii)$ Process uncertainty bound. Define $Y_{t}:=X_{t}^{1/2}\Delta_{t}X_{t}^{1/2}\succeq 0$. Then,
-$X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}=X_{t}^{1/2}Y_{t}(I+Y_{t})^{-1}X_{t}^{1/2}.$
-and since $Y_{t}\succeq 0$,
-$\|Y_{t}(I+Y_{t})^{-1}\|_{2}=\lambda_{\max}(Y_{t})/(1+\lambda_{\max}(Y_{t}))=\|Y_{t}\|_{2}/(1+\|Y_{t}\|_{2}).$
-Therefore,
-$\|X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}\|_{2}\leq\|X_{t}^{1/2}\|_{2}^{2}\cdot\frac{\|Y_{t}\|_{2}}{1+\|Y_{t}\|_{2}}.$
-However,
-$\|Y_{t}\|_{2}=\|X_{t}^{1/2}\Delta_{t}X_{t}^{1/2}\|_{2}\leq\|X_{t}\|_{2}\|\Delta_{t}\|_{2},$
-resulting in the bound
-$\|X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}\|_{2}\leq\frac{\|X_{t}\|_{2}^{2}\,\|\Delta_{t}\|_{2}}{1+\|X_{t}\|_{2}\|\Delta_{t}\|_{2}}.$ (43)
-Now, for the minimum-norm construction of $\Delta_{t}$ in (22),
-$\|\Delta_{t}\|_{2}\leq\|A^{+}\|_{2}^{2}\,\|\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}\|_{2}.$ (44)
-By Lemma˜4 in Appendix C.3, $\|\Sigma_{w,t}^{*}-\hat{\Sigma}_{w}\|_{2}\leq\theta_{w}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}\theta_{w}$. On the other hand,
-$\begin{split}\|X_{t}\|_{2}&\leq\|(\Sigma_{x,t}^{-})^{-1}\|_{2}+\|C\|_{2}^{2}\|(\Sigma_{v,t}^{*})^{-1}\|_{2}\\ &=\frac{1}{\lambda_{\min}(\Sigma_{x,t}^{-})}+\frac{\|C\|_{2}^{2}}{\lambda_{\min}(\Sigma_{v,t}^{*})}.\end{split}$
-By Corollary˜2, $\lambda_{\min}(\Sigma_{v,t}^{*})\geq\lambda_{\min}(\hat{\Sigma}_{v})$. Moreover, since by definition
-$\Sigma_{x,t}^{-}=A\Sigma_{x,t-1}A^{\top}+\Sigma_{w,t-1}^{*}\succeq\Sigma_{w,t-1}^{*},$
-we have $\lambda_{\min}(\Sigma_{x,t}^{-})\geq\lambda_{\min}(\Sigma_{w,t-1}^{*})\geq\lambda_{\min}(\hat{\Sigma}_{w})$. Thus
-$\|X_{t}\|_{2}\leq\delta_{x}^{\max}$ (45)
-with $\delta_{x}^{\max}$ as defined in (28).
-By the monotonicity of $f(x,y):=x^{2}y/(1+xy)$ on $\real{2}_{+}$, replacing $\|X_{t}\|_{2}$ and $\|\Delta_{t}\|_{2}$ in (43) by their respective upper bounds from (44) and (45) preserves the inequality.
-We thus obtain
-$\|X_{t}(I+\Delta_{t}X_{t})^{-1}\Delta_{t}X_{t}\|_{2}\leq\frac{(\delta_{x}^{\max})^{2}\|A^{+}\|_{2}^{2}\,\bigl(\theta_{w}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}\theta_{w}\bigr)}{1+\delta_{x}^{\max}\|A^{+}\|_{2}^{2}\,\bigl(\theta_{w}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}\,\theta_{w}\bigr)}.$
-Requiring the right-hand side to be $\leq\phi_{x}$ is equivalent to
-$\theta_{w}^{2}+2\sqrt{\lambda_{\max}(\hat{\Sigma}_{w})}\theta_{w}\leq\frac{\phi_{x}}{\|A^{+}\|_{2}^{2}\,\delta_{x}^{\max}\,(\delta_{x}^{\max}-\phi_{x})},$
-which ensures (30). Solving the quadratic inequality yields the explicit bound (29).
-Combining $(i)$ and $(ii)$ gives $\Phi_{t}\preceq\phi_{v}I+\phi_{x}I\preceq\phi_{N}I,t\geq q$.
-Since $\phi_{v}+\phi_{x}\leq\phi_{N}$Proposition˜2 applies: the downsampled DR Riccati map is a contraction in the Thompson part metric.
-By Banach’s fixed-point theorem, the recursion admits a unique fixed point, and the DR Kalman gain converges.
+$\|Z_{t}\|_{2}^{2}\leq\|\hat{\Sigma}_{v}\|_{2}\ \|\Sigma_{v,t}^{*}\|_{2}\leq\|\hat{\Sigma}_{v}\|_{2}\ \bar{\lambda}_{v},$
+so $\{Y_{t}\}$ and $\{Z_{t}\}$ are uniformly bounded.
+Therefore, $\{z_{t}\}$ is bounded, and by the Bolzano–Weierstrass theorem, there exists a subsequence $t_{k}$ such that $z_{t_{k}}\to z_{\infty}$.
+Writing
+$z_{\infty}:=(\Sigma_{x,\infty}^{-},\Sigma_{x,\infty},\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*},Y_{\infty},Z_{\infty}),$
+we have $(\Sigma_{w,t_{k}-1}^{*},\Sigma_{v,t_{k}}^{*})\to(\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*})$ along this subsequence.
+Passing to the limit along this subsequence in the constraints of (7), and using $\Sigma_{x,t_{k}-1}\to\Sigma_{x,\infty}$ and the closedness of the PSD cone, shows that $z_{\infty}$ is feasible for the stationary SDP (13).
+We now show that $z_{\infty}$ is in fact optimal for (13).
+Assume $\theta_{w}>0$ and $\theta_{v}>0$.
+[^16] The degenerate case $\theta_{w}=\theta_{v}=0$ reduces to the nominal KF, for which the statement is immediate.
+Then (13) satisfies Slater’s condition.
+For instance, choose $\Sigma_{w,\infty}=\hat{\Sigma}_{w}+\varepsilon I$, $\Sigma_{v,\infty}=\hat{\Sigma}_{v}+\varepsilon I$, $Y=\hat{\Sigma}_{w}$, $Z=\hat{\Sigma}_{v}$, $\Sigma_{x,\infty}=0$, and $\Sigma_{x,\infty}^{-}=\Sigma_{w,\infty}$ for sufficiently small $\varepsilon>0$; this makes all LMIs $\succ 0$ and the trace inequalities strict.
+Therefore, strong duality holds and the KKT conditions are necessary and sufficient for optimality.
+For each $t\geq 1$, let $\lambda_{t}$ be a dual optimal solution associated with the primal optimizer $z_{t}$ of (7) (existence follows from Slater’s theorem).
+Moreover, since the primal feasible sets are uniformly bounded and Slater’s condition holds, the set of primal–dual optimal pairs is bounded.
+Thus, after passing to a further subsequence if necessary, we have $\lambda_{t_{k}}\to\lambda_{\infty}$.
+Taking limits in the KKT conditions of (7) along $t_{k}$ (all primal/dual feasibility constraints are closed, and complementarity is preserved under limits) yields that $(z_{\infty},\lambda_{\infty})$ satisfies the KKT conditions of the stationary SDP (13).
+Consequently, $z_{\infty}$ is optimal for (13), and $(\Sigma_{x,\infty}^{-},\Sigma_{x,\infty})=(\Sigma_{x,\infty}^{-,*},\Sigma_{x,\infty}^{*})$.
+Finally, by Theorem˜3, the gain sequence $\{K_{t}\}$ converges.
+Along the subsequence $t_{k}$, $(\Sigma_{x,t_{k}}^{-},\Sigma_{v,t_{k}}^{*})\to(\Sigma_{x,\infty}^{-,*},\Sigma_{v,\infty}^{*})$, and continuity of $K(\Sigma_{x,t}^{-},\Sigma_{v})=\Sigma_{x,t}^{-}C^{\top}(C\Sigma_{x,t}^{-}C^{\top}+\Sigma_{v})^{-1}$ on $\mathbb{S}^{n_{x}}_{++}\times\mathbb{S}^{n_{y}}_{++}$ yields $\lim_{k\to\infty}K_{t_{k}}=K_{\infty}^{*}$.
+Since $\{K_{t}\}$ has a unique limit, we conclude that $K_{\infty}^{*}=\lim_{t\to\infty}K_{t}$.
 Appendix D Proofs for Stability and Optimality
 D.1 Proof of Theorem˜4
-Proof.
 Let $\tilde{e}_{t}:=x_{t}-\bar{x}^{-}_{t}$ denote the prediction error. From the DRKF update,
 […]
-The prediction dynamics satisfy
-$\tilde{e}_{t}=Ae_{t-1}+w_{t-1}-\hat{w}$.
-Eliminating $\tilde{e}_{t}$ yields the posterior error recursion
+The prediction error evolves as $\tilde{e}_{t}=Ae_{t-1}+w_{t-1}-\hat{w}$. Eliminating $\tilde{e}_{t}$ yields the posterior error recursion
 […]
 […]
-At steady state, the DRKF covariances satisfy
+At steady-state, the DRKF covariances satisfy
 $\Sigma_{x,\infty}^{-,*}=A\Sigma_{x,\infty}^{*}A^{\top}+\Sigma_{w,\infty}^{*}$
-with
-$K_{\infty}=\Sigma_{x,\infty}^{-,*}C^{\top}\!\bigl(C\,\Sigma_{x,\infty}^{-,*}C^{\top}+\Sigma_{v,\infty}^{*}\bigr)^{-1}.$
-Using the Joseph form,
-$\Sigma_{x,\infty}^{*}\;=\;(I-K_{\infty}C)\,\Sigma_{x,\infty}^{-,*}\,(I-K_{\infty}C)^{\top}+K_{\infty}\Sigma_{v,\infty}^{*}K_{\infty}^{\top}.$
+with $K_{\infty}^{*}=\Sigma_{x,\infty}^{-,*}C^{\top}\!\bigl(C\,\Sigma_{x,\infty}^{-,*}C^{\top}+\Sigma_{v,\infty}^{*}\bigr)^{-1}$. Using the Joseph form,
+$\Sigma_{x,\infty}^{*}\;=\;(I-K_{\infty}^{*}C)\,\Sigma_{x,\infty}^{-,*}\,(I-K_{\infty}^{*}C)^{\top}+K_{\infty}^{*}\Sigma_{v,\infty}^{*}(K_{\infty}^{*})^{\top}.$
 Substituting $\Sigma_{x,\infty}^{-,*}=A\Sigma_{x,\infty}^{*}A^{\top}+\Sigma_{w,\infty}^{*}$ gives
-$\Sigma_{x,\infty}^{*}$ $=F_{\infty}\,\Sigma_{x,\infty}^{*}\,F_{\infty}^{\top}+(I-K_{\infty}C)\,\Sigma_{w,\infty}^{*}\,(I-K_{\infty}C)^{\top}+K_{\infty}\Sigma_{v,\infty}^{*}K_{\infty}^{\top}.$
-Rearranging,
-$\Sigma_{x,\infty}^{*}-F_{\infty}\Sigma_{x,\infty}^{*}F_{\infty}^{\top}=(I-K_{\infty}C)\,\Sigma_{w,\infty}^{*}\,(I-K_{\infty}C)^{\top}+K_{\infty}\Sigma_{v,\infty}^{*}K_{\infty}^{\top}\succeq 0.$
-Since $\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*}\succ 0$, the right-hand side is in fact $\succ 0$, implying the strict Lyapunov inequality $\Sigma_{x,\infty}^{*}\succ F_{\infty}\Sigma_{x,\infty}^{*}F_{\infty}^{\top}.$ In fact, it implies that $F_{\infty}$ is Schur stable.
-Indeed, if $F_{\infty}$ had an eigenvalue on or outside the unit circle, no $\Sigma_{x,\infty}^{*}\succ 0$ could satisfy $\Sigma_{x,\infty}^{*}\succ F_{\infty}\Sigma_{x,\infty}^{*}F_{\infty}^{\top}$.
-[^9] This is the standard equivalence: $F$ is Schur $\Leftrightarrow$ there exists $P\succ 0$ with $P-FPF^{\top}\succ 0$.
-With $F_{\infty}$ Schur, the homogeneous error dynamics $e_{t}^{0}=F_{\infty}e_{t-1}^{0}$ thus converge to 0.
-With noises, the covariance of $e_{t}$ under the least-favorable distributions obeys the steady-state equation derived above and thus converges to $\Sigma_{x,\infty}^{*}$.
-Taking traces gives the claimed convergence and boundedness.
+$\Sigma_{x,\infty}^{*}=F_{\infty}\,\Sigma_{x,\infty}^{*}\,F_{\infty}^{\top}+(I-K_{\infty}^{*}C)\,\Sigma_{w,\infty}^{*}\,(I-K_{\infty}^{*}C)^{\top}+K_{\infty}^{*}\Sigma_{v,\infty}^{*}(K_{\infty}^{*})^{\top}.$
+Hence,
+$\Sigma_{x,\infty}^{*}-F_{\infty}\Sigma_{x,\infty}^{*}F_{\infty}^{\top}=Q_{\infty},$
+where $Q_{\infty}:=(I-K_{\infty}^{*}C)\Sigma_{w,\infty}^{*}(I-K_{\infty}^{*}C)^{\top}+K_{\infty}^{*}\Sigma_{v,\infty}^{*}(K_{\infty}^{*})^{\top}$.
+Since $\Sigma_{w,\infty}^{*},\Sigma_{v,\infty}^{*}\succ 0$, for any $x\neq 0$,
+$x^{\top}Q_{\infty}x=\|\Sigma_{w,\infty}^{*,1/2}(I-K_{\infty}^{*}C)^{\top}x\|^{2}+\|\Sigma_{v,\infty}^{*,1/2}(K_{\infty}^{*})^{\top}x\|^{2}>0,$
+which implies $Q_{\infty}\succ 0$.
+Thus, $\Sigma_{x,\infty}^{*}\succ F_{\infty}\Sigma_{x,\infty}^{*}F_{\infty}^{\top}$.
+By the standard discrete-time Lyapunov characterization, this strict inequality implies that $F_{\infty}$ is Schur stable.
+Using the temporal independence of $\{w_{t}\}$ and $\{v_{t}\}$ and their independence from $e_{t-1}$, the error covariance $P_{t}:=\mathrm{cov}(e_{t})$ satisfies
+$P_{t}=F_{\infty}P_{t-1}F_{\infty}^{\top}+Q_{\infty},$
+whose unique fixed point is $\Sigma_{x,\infty}^{*}$. Since $F_{\infty}$ is Schur, $P_{t}\to\Sigma_{x,\infty}^{*}$ and $\sup_{t\geq 0}\mathbb{E}[\|e_{t}\|^{2}]<\infty$, completing the proof.
 D.2 Proof of Corollary˜5
-Proof.
-Taking expectations in (46) yields
 […]
-Since $F_{\infty}$ is Schur by Theorem˜4, the affine recursion converges, yielding
 […]
-If $\hat{w}=\mu_{w}$ and $\hat{v}=\mu_{v}$, the constant input term vanishes, and thus $m_{t}\to 0$. ∎
+If $\hat{w}=\mu_{w}$ and $\hat{v}=\mu_{v}$, the constant input term vanishes, and thus $m_{t}\to 0$.
 D.3 Proof of Theorem˜5
-Proof.
-By the strong duality of Wasserstein DRO and the orthogonality principle for quadratic losses, the one-step minimax problem admits a saddle point $(\psi_{t}^{*},\mathbb{P}_{e,t}^{*})$ such that
-$\inf_{\psi_{t}\in\mathcal{F}_{t}}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})=\operatorname{Tr}[\Sigma_{x,t}^{*}],$
-where $\Sigma_{x,t}^{*}=\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t}^{*})$ is the DRKF posterior covariance.
-Thus, for any admissible $\psi_{t}$,
-$\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})\geq\operatorname{Tr}[\Sigma_{x,t}^{*}],$
-almost surely.
-Since $\Sigma_{x,t}^{*}\to\Sigma_{x,\infty}^{*}$ by Theorem˜3, both lower bounds (31) and (32) follow.
-For the steady-state DRKF $\psi_{\infty}$ with gain $K_{\infty}$, the least-favorable posterior covariance is exactly $\Sigma_{x,\infty}^{*}$.
-Hence, the worst-case one-step MSE equals $\operatorname{Tr}[\Sigma_{x,\infty}^{*}]$ for all sufficiently large $t$, establishing equality in the pointwise limit and in the long-run average.
-∎
+For each $t\geq 0$, define the stage-wise minimax value
+$V_{t}\coloneqq\inf_{\psi_{t}\in\mathcal{F}_{t}}\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t}),$
+where $J_{t}(\psi_{t},\mathbb{P}_{e,t})$ is understood almost surely with respect to $\sigma(\mathcal{Y}_{t-1})$.
+Under Assumption˜1, Lemma˜1 and Theorem˜1 imply that the minimax problem admits a Gaussian least-favorable distribution and that the corresponding optimal estimator is the DRKF $\psi_{t}^{*}$.
+Moreover, the minimax value is given by $V_{t}=\operatorname{Tr}[\Sigma_{x,t}]$, where $\Sigma_{x,t}=\Psi(\Sigma_{x,t}^{-},\Sigma_{v,t}^{*})$ is the DRKF posterior covariance at time $t$.
+Consequently, for any admissible estimator $\psi_{t}\in\mathcal{F}_{t}$,
+$\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t},\mathbb{P}_{e,t})\geq V_{t}=\operatorname{Tr}[\Sigma_{x,t}].$
+Taking $\liminf_{t\to\infty}$ and using Theorem˜3, which guarantees $\Sigma_{x,t}\to\Sigma_{x,\infty}^{*}$, yields (27).
+The corresponding long-run average lower bound (28) follows by Cesàro summability.
+We now prove the asymptotic optimality of the steady-state DRKF $\psi_{\infty}$ with gain $K_{\infty}^{*}$.
+Let $\bar{x}_{t}^{\infty}:=\psi_{\infty}(\mathcal{Y}_{t})$ and $\bar{x}_{t}^{*}:=\psi_{t}^{*}(\mathcal{Y}_{t})$ denote the posterior estimates produced by the steady-state and time-varying DRKFs, respectively.
+Define the difference $d_{t}:=\bar{x}_{t}^{\infty}-\bar{x}_{t}^{*}$.
+A direct comparison of the measurement updates yields
+$d_{t}=F_{\infty}d_{t-1}+\Delta_{t}\eta_{t}^{*},\quad t\geq 1,$ (36)
+where $F_{\infty}:=(I-K_{\infty}^{*}C)A$, $\Delta_{t}:=K_{\infty}^{*}-K_{t}$, and $\eta_{t}^{*}:=y_{t}-C{\bar{x}^{-}_{t}}^{*}-\hat{v}$ is the innovation of the time-varying DRKF.
+By Theorem˜4, $F_{\infty}$ is Schur, and by Theorem˜3 we have $\|\Delta_{t}\|_{2}\to 0$.
+[…]
+$\begin{split}\mathbb{E}\left[\|\eta_{t}^{*}\|^{2}\mid\mathcal{Y}_{t-1}\right]\leq 2\|C\|_{2}^{2}\,\mathbb{E}\left[\|x_{t}-{\bar{x}^{-}_{t}}^{*}\|^{2}\mid\mathcal{Y}_{t-1}\right]+2\,\mathbb{E}\left[\|v_{t}-\hat{v}\|^{2}\mid\mathcal{Y}_{t-1}\right].\end{split}$
+Under Assumption˜2, the ambiguity sets have finite Wasserstein radii, so all $\mathbb{P}_{v,t}\in\mathbb{D}_{v,t}$ have uniformly bounded second moments about $\hat{v}$.
+Moreover, Theorem˜3 implies $\Sigma_{x,t}\to\Sigma_{x,\infty}^{*}$, and hence $\sup_{t\geq 0}\operatorname{Tr}[\Sigma_{x,t}]<\infty$.
+Since $\Sigma_{x,t}^{-}=A\Sigma_{t-1}A^{\top}+\Sigma_{w,t-1}^{*}$ and $\Sigma_{w,t-1}^{*}\preceq\overline{\lambda}_{w}I$ by Corollary˜2, it follows that $\sup_{t\geq 0}\operatorname{Tr}[\Sigma_{x,t}^{-}]<\infty$, and therefore
+$\sup_{t\geq 0}\mathbb{E}[\|x_{t}-{\bar{x}^{-}_{t}}^{*}\|^{2}\mid\mathcal{Y}_{t-1}]<\infty.$
+Consequently, there exists $M_{\eta}<\infty$ such that
+$\sup_{t\geq 0}\ \sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}\mathbb{E}\left[\|\eta_{t}^{*}\|^{2}\mid\mathcal{Y}_{t-1}\right]\leq M_{\eta}.$
+Therefore,
+$\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}\mathbb{E}\left[\|\Delta_{t}\eta_{t}^{*}\|^{2}\mid\mathcal{Y}_{t-1}\right]\leq\|\Delta_{t}\|_{2}^{2}\,M_{\eta}\xrightarrow[t\to\infty]{}0.$
+Since $F_{\infty}$ is Schur, iterating (36) and using the above uniform bound yields that $d_{t}\to 0$ in mean square; in particular,
+$D_{t}:=\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}\mathbb{E}\left[\|d_{t}\|^{2}\mid\mathcal{Y}_{t-1}\right]\xrightarrow[t\to\infty]{}0.$
+Let $e_{t}^{*}:=x_{t}-\bar{x}_{t}^{*}$ denote the time-varying DRKF posterior error. Since $x_{t}-\bar{x}_{t}^{\infty}=e_{t}^{*}-d_{t}$, we have
+$\|x_{t}-\bar{x}_{t}^{\infty}\|^{2}\leq\|e_{t}^{*}\|^{2}+\|d_{t}\|^{2}+2\|e_{t}^{*}\|\|d_{t}\|.$
+Taking conditional expectations and then the supremum over $\mathbb{D}_{e,t}$ yields
+$\begin{split}\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})\leq\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t}^{*},\mathbb{P}_{e,t})+D_{t}+2\sqrt{D_{t}}\sqrt{\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t}^{*},\mathbb{P}_{e,t})}.\end{split}$
+By stage-wise minimax optimality of the time-varying DRKF, $\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{t}^{*},\mathbb{P}_{e,t})=V_{t}=\operatorname{Tr}[\Sigma_{x,t}]$. Thus,
+$\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})\leq V_{t}+D_{t}+2\sqrt{V_{t}}\sqrt{D_{t}}.$ (37)
+Taking $\limsup_{t\to\infty}$ and using $D_{t}\to 0$ and $V_{t}\to\operatorname{Tr}[\Sigma_{x,\infty}^{*}]$ gives
+$\limsup_{t\to\infty}\sup_{\mathbb{P}_{e,t}\in\mathbb{D}_{e,t}}J_{t}(\psi_{\infty},\mathbb{P}_{e,t})\leq\operatorname{Tr}[\Sigma_{x,\infty}^{*}].$
+Combined with the lower bound (27) applied to $\psi_{\infty}$, this proves (29).
+The Cesàro-average optimality statement (30) follows similarly since $\sup_{t}V_{t}<\infty$ and $D_{t}\to 0$ imply $\frac{1}{T}\sum_{t=0}^{T-1}D_{t}\to 0$, and hence all error terms in (37) have vanishing Cesàro mean.